### v1

# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

Abstract
In this article, we improve on two explicit bounds of order $\log t$ for $\sigma$ close to $1$: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$. Our method provides marked improvements in the constants, especially with regards to the latter. These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s.
1 Introduction
In the study of the Riemann zeta function $\zeta(s)$ and related functions, one often requires upper bounds for the logarithmic derivative of $\zeta(s)$ and the reciprocal of $\zeta(s)$. For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function. The reader is referred to [25], [3], [8], [24, (3.14)] for more information on the topic.
Trudgian in [25] gives an explicit version of results in [24, §3] originally due to Landau, obtaining bounds of order $O(\log t)$. This was then extended in [13] and [6] to bounds of order $O(\log t/\log\log t)$. However, such estimates tend to have large constants, often rendering them unsuitable for practical use. It is a non-trivial task to reduce these constants. This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
The method in [24, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$. It relies on two applications of the Borel–Carathédory theorem. The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$. As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [12, Lemma 3.2].
While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [25], we take the opportunity to point out an error in that paper. In the quantity $A$ of [25, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$. This leads to the overall constants being better than they should be. For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,473)$ and $(12,54.3)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively. In this article, we have recovered the erroneous bounds and even slightly improved upon them.
Furthermore, we directly derive a bound on $|1/\zeta(s)|$ from ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ rather than $|\zeta^{\prime}(s)/\zeta(s)|$, eliminating the need for the Borel–Carathédory theorem in this case, leading to a substantial improvement over [25]. In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [2, Proposition A.2].
2 Main results
Here we address matters of notation before listing some of the main results we prove.
2.1 Properties of the Riemann zeta function
[…]
The classical zero-free region for $\zeta(s)$ states that for $t\geq 2$,
$\zeta(s)\neq 0,\qquad\text{ for }\qquad\sigma\geq 1-\frac{1}{W_{0}\log t}.$
Currently, the best known result is due to [18], where above $H_{0}$, the value $W_{0}=5.558691$ is permissible. We take $W_{0}$ to be this value.
We denote the Riemann hypothesis up to height $T$ by $RH(T)$, by which we mean: for all zeros $\rho=\beta+i\gamma$ such that $\zeta(\rho)=0$ and $\gamma\leq T$, we have $\beta=1/2$. Similarly we denote the Riemann hypothesis by $RH$, by which we mean all zeros $\rho$ have $\beta=1/2$.
2.2 Bounds on $|\zeta^{\prime}(s)/\zeta(s)|$
Here we present our main results, which are corollaries of the theorems in Sections 4 and 5.
Corollary 1.
Let $W\geq W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 5$,
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q\log t,$
where $(W,Q)=(10,49.359)$ is admissible, and other values of $Q$ with corresponding $W$ are presented in Table 3.
Corollary 2.
For $\sigma\geq 1$ and $t\geq 5$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq 18.062\log t.$
Corollary 3.
Assume $RH(H_{0})$. Then in the region
$\sigma\geq\sigma_{0}>1/2\qquad\text{ and }\qquad 5\leq t\leq H,$
we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t,$
where $(\sigma_{0},Q_{RH(H_{0})})=(0.8,33.249)$ is admissible, and other values for $Q_{RH(H_{0})}$ with corresponding values of $\sigma_{0}$ are given in Table 2. Furthermore, by assuming $RH$ instead of $RH(H_{0})$, the upper bound restriction on $t$ can be removed.
2.3 Bounds on $|1/\zeta(s)|$
Corollary 4.
Let $W\geq W_{0}$. In a region
$\sigma\geq 1-\frac{1}{W\log t}$
where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y(W,t_{0})\log t,$
where $Y(W_{0},10^{2})=4.339$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
In addition, we have for $\sigma=1$ and $t\geq 2$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.751\log t.$
3 Some preliminaries
In this section, we present the results used in obtaining our estimates. For convenience, we begin by stating a few function theoretic results, which lie at the heart of our method.
Theorem 1.
[13, Theorem 4][Borel–Carathéodory] Let $s_{0}$ be a complex number, and let $R$ be a positive number possibly dependent on $s_{0}$. Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$. Let $M$ denote the maximum of $|\textup{Re}\,f(s)|$ on the boundary $|s-s_{0}|=R$. Then, for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
$|f(s)|\leq\frac{2r}{R-r}M+\frac{R+r}{R-r}|f(s_{0})|.$
Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [12, Lemma 3.2]. The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [23, §3.71]).
Lemma 1.
[12, Lemma 3.2] Let $f(z)$ be holomorphic for $|z-a|\leq R$, and non-vanishing both at $z=a$ and on the circle $|z-a|=R$. Let $\rho_{j}=a+r_{j}\exp(i\theta_{j})$ be the zeros of $f(z)$ in the disc, where $\rho_{j}$ has multiplicity $n_{j}$. Then
${\textup{Re}}\frac{f^{\prime}}{f}(a)=-\sum_{\rho_{j}}n_{j}\left(\frac{1}{r_{j}}-\frac{r_{j}}{R^{2}}\right)\cos\theta_{j}+\frac{1}{\pi R}\int_{0}^{2\pi}(\cos\theta)\log|f(a+Re^{i\theta})|\text{d}\theta.$
Lemma 2.
Let $d>0$ be some real constant and suppose that $\zeta(s)\neq 0$ in the region $\sigma\geq 1-1/W_{0}\log t$. Let $R^{\prime}>0$ be a parameter, chosen such that
$\frac{d}{\log t}\leq\frac{1}{W_{0}\log(t+R^{\prime})}.$ (1)
Then for $t\geq 3$, we have
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\leq\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+it+R^{\prime}e^{i\theta})||,$
which holds uniformly for
$\sigma\geq 1-\frac{d}{\log t}.$ (2)
Proof.
For $\sigma_{0}$ in the region stated in (2), we take $R=R^{\prime}$, $f(s)=\zeta(s)$, $a=s_{0}=\sigma_{0}+it_{0}$, and $\rho_{k}=s_{0}+r_{k}\exp(i\theta_{k})$ to be a zero of $\zeta(s)$ in Lemma 1 (where we will make the substitution $s_{0}\to s$ at the end of the proof). This gives us
${\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(s_{0})=-\sum_{\rho_{k}}n_{k}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}+\frac{1}{\pi R^{\prime}}\int_{0}^{2\pi}(\cos\theta)\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta.$
Note that if a zero of $\zeta(s)$ lies on the boundary $|s-s_{0}|=R^{\prime}$, then since the zeros are separable, we can always find an arbitrarily smaller $R^{\prime}$ where the boundary is free of zeros.
For convenience, we define
$J:=\int_{0}^{2\pi}(\cos\theta)\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta,$
for which we seek a lower bound.
For $20\leq\theta\leq\pi/$ 2 and $3\pi/2\leq\theta\leq 2\pi$, i.e. when $0\leq\cos\theta\leq 1$, we have
$\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\geq-\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||,$
so in the ranges $20\leq\theta\leq\pi/$ 2 and $3\pi/2\leq\theta\leq 2\pi$ combined,
$J\geq-2\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||.$
For the remaining range $\pi/2\leq\theta\leq 3\pi/2$, noting that $-1\leq\cos\theta\leq 0$, we have
$\int_{\pi/2}^{3\pi/2}(\cos\theta)$ $\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta\geq-2\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||,$
and hence for the whole range $\theta\in[0,2\pi]$,
$J\geq-4\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||.$
Now Lemma 1 yields
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(s_{0})=\sum_{|s_{0}-\rho|\leq R^{\prime}}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}-\frac{1}{\pi R^{\prime}}J,$ (3)
and since $r_{k}\cos\theta_{k}=-{\textup{Re}}(s_{0}-\rho_{k})$ and
$\frac{\cos\theta_{k}}{r_{k}}=\frac{r_{k}\cos\theta_{k}}{r_{k}^{2}}=\frac{-{\textup{Re}}(s_{0}-\rho_{k})}{|s_{0}-\rho_{k}|^{2}}=-{\textup{Re}}\left(\frac{1}{s_{0}-\rho_{k}}\right),$
the sum over all zeros in $|s_{0}-\rho|\leq R^{\prime}$ from (3) is
$\sum_{|s_{0}-\rho|\leq R^{\prime}}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}$ $=\sum_{|s_{0}-\rho|\leq R^{\prime}}-{\textup{Re}}\left(\frac{1}{s_{0}-\rho}-\frac{s_{0}-\rho}{R^{\prime 2}}\right)$
$=-\sum_{|s_{0}-\rho|\leq R^{\prime}}(\sigma_{0}-\beta)\left(\frac{1}{|s_{0}-\rho|^{2}}-\frac{1}{R^{\prime 2}}\right),$ (4)
where the second factor in (3) is non-negative.
Notice that if
$1-\frac{1}{W_{0}\log(t_{0}+R^{\prime})}\leq\sigma_{0},$
then by our non-vanishing condition on $\zeta(s)$, we always have that $\sigma_{0}-\beta\geq 0$ in the disc $|s_{0}-\rho|\leq R^{\prime}$. This condition is fulfilled by our choice of $R^{\prime}$ in (1). Thus the entire right-hand side of (3) is negative, and so we can discard any contribution from this.
Combining this with the bound on $J$ and (3) proves the theorem. ∎
The following corollary is useful for proving results in a right-half plane, under $RH(T)$ or $RH$.
Corollary 5.
Let $\epsilon>0$ be a constant and suppose that $\zeta(s)\neq 0$ in the region $\sigma>1/2$ and $t\leq t+\epsilon$. Then for a fixed $R^{\prime}<\epsilon$ and $t\geq 3$, we have
$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\leq\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+it+R^{\prime}e^{i\theta})||,$
which holds uniformly for
$\sigma\geq\frac{1}{2}+\epsilon.$ (5)
Proof.
The proof is analogous to that of Lemma 2 with the additional assumption that the disc $|s_{0}-\rho|\leq R^{\prime}$ is completely free of zeros. ∎
Next we require some uniform bounds for $\zeta(s)$. We slightly improve [26, Corollary 1], which is proven using the Phragmén–Lindelöf principle. Note that we also use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$. This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing in $\sigma$. Additionally, we found that further splitting the interval complicates (8) and raises the threshold of $t$ for the lemma to be valid.
Lemma 3.
Let there be a real number $h(s)=(s-1)\zeta(s)$, and let $\delta_{r}$ be a positive real number. Furthermore, let $Q_{0}\geq 0$ be a number for which
$|h(\tfrac{1}{2}+it)|$ $\leq k_{1}|Q_{0}+\tfrac{1}{2}+it|^{k_{2}+1}(\log|Q_{0}+\tfrac{1}{2}+it|)^{k_{3}}$ (6)
$|h(1+\delta_{r}+it)|$ $\leq\zeta(1+\delta_{r})|Q_{0}+1+\delta_{r}+it|,$
for all $t$. Then for $\sigma\in[\tfrac{1}{2},1+\delta_{r}]$ and $t\geq t_{0}\geq e$, we have
$|\zeta$ $(s)|\leq\left(\zeta(1+\delta_{r})(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))\right)^{\frac{\sigma-0.5}{0.5+\delta_{r}}}$ (7)
×$\bm{\times}\left(k_{1}(1+a_{0}(1+\delta_{r},Q_{0},t_{0}))^{k_{3}}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}+1}t^{k_{2}}\log^{k_{3}}t\right)^{\frac{1+\delta_{r}-\sigma}{0.5+\delta_{r}}},$
where
$a_{0}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{2t^{2}\log t}+\frac{\pi}{2\log t}+\frac{\pi(\sigma+Q_{0})^{2}}{4t\log^{2}t},\qquad a_{1}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{t},$
and the right-hand side of (7) is decreasing in $\sigma$, provided that $k_{i}\geq 0$, and the following hold:
$t^{k_{2}}\log^{k_{3}}t\geq\frac{k_{1}(1+a_{0})^{k_{3}}(1+a_{1})^{k_{2}}}{\zeta(1+\delta_{r})}.$ (8)
Proof.
The proof is identical to that of [26, Lemma 2.1], except that we are more precise in estimating (7) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$. ∎
Corollary 6.
For $\sigma\in[\tfrac{1}{2},1+\delta_{r}]$ and $t\geq t_{0}\geq e$, we have
$|\zeta(s)|\leq 0.618(1+a_{0}(1+\delta_{r},1.31,t_{0}))(1+a_{1}(1+\delta_{r},1.31,t_{0}))^{7/6}t^{1/6}\log t,$
provided that
$t^{1/6}\log t\geq\frac{0.618(1+a_{0})(1+a_{1})^{1/6}}{\zeta(1+\delta_{r})}.$ (9)
Proof.
Recall that Hiary, Patel, and Yang [14] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$.
We now refer to Lemma 3, taking the maximum of (7) at $\sigma=1/2$. It has been shown in [27, p. 15] that $Q_{0}=1.31$ is permissible in (6) together with
$(k_{1},k_{2},k_{3},Q_{0})=\left(0.618,\tfrac{1}{6},1,1.31\right),$
hence we are done. ∎
In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones. We would like to point out that Lemma 5 and Corollary 8, which can both be viewed as an extension of Corollary 7 inside the critical strip, are not actually required for the purposes of this paper. However, we include them here as the proofs are similar to that of Corollary 7, and at the same time we provide a brief discussion on their usefulness. Lemma 5 is an improvement on [15, Proposition 5.3.2], which proves a similar result.
Lemma 4.
[22, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\frac{4}{\sigma-(1/2)}.$
Lemma 5.
Let $W>0$ be a real constant. Then in the region
$1-\frac{1}{W\log t}\leq\sigma\leq\sigma_{1},$
we have for $t\geq t_{0}\geq 3$ and any $\eta\in[2/t_{0},1-1/t_{0}]$ that
$|\zeta(\sigma+it)|\leq e^{1/W}\log t+C_{0}(W,\sigma_{1},t_{0},\eta)$
where
$C_{0}(W,\sigma_{1},t_{0},\eta)=$ $e^{1/W}\Bigg{(}\log\left(\eta+\frac{1}{t_{0}}\right)+\gamma+\frac{1}{12\eta^{2}}$
$+\frac{1}{t_{0}}\left(\frac{1}{6\eta^{2}}+\frac{1}{\eta}+1\right)+\frac{16(\sigma_{1}^{2}+2\sigma_{1}-1)+3}{192\eta^{2}t_{0}^{2}}\Bigg{)}.$
Proof.
We will use the following representation for $\zeta(s)$ for $N\geq 2$,
$\zeta(s)=\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}+\frac{N^{1-s}}{s-1}+\frac{s}{12N^{s+1}}-\frac{s(s+1)}{2}\int_{N}^{\infty}\frac{\varphi^{*}(u)}{u^{s+2}}\text{d}u,$ (10)
where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$. Note that (10) holds for $\sigma>0$ by analytic continuation (see [24, Theorem 3.5]).
In the region of $\sigma$ considered, we have
$|N^{-s}|=N^{-\sigma}\leq\exp\left(-\Big{(}1-\frac{1}{W\log t}\Big{)}\log N\right)=N^{-1}\exp\left(\frac{\log N}{W\log t}\right).$ (11)
This bound can be used throughout (10). The first term in (10) can also be bounded using the estimate in [10, Lemma 2.1], so for $N\geq 2$,
$\left|\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}\right|$ $=\left|\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{1}{2N^{s}}\right|$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\sum_{n=1}^{N}\frac{1}{n}+\frac{1}{2N}\right)$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}-\frac{1}{12N^{2}}+\frac{1}{64N^{4}}\right).$
The term $N^{1-s}/(s-1)$ will be bounded by $t^{-1}\exp\left(\frac{\log N}{W\log t}\right)$. The integral in (10) can be evaluated using the bound $|\varphi^{*}(u)|\leq 1/6$, so
$\int_{N}^{\infty}\frac{|\varphi^{*}(u)|}{u^{\sigma+2}}\text{d}u\leq\frac{1}{6(\sigma+1)}\frac{1}{N^{\sigma+1}}\leq\frac{1}{6}\exp\left(\frac{\log N}{W\log t}\right)\frac{1}{N^{2}}.$
This implies
$\left|\frac{s}{12N^{s+1}}-\frac{s^{2}+s}{2}\int_{N}^{\infty}\frac{\varphi^{*}(u)}{u^{s+2}}\text{d}u\right|$ $\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\frac{2|s|}{12N^{2}}+\frac{|s^{2}|}{12N^{2}}\right)$
$\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\frac{t+\sigma_{1}}{6N^{2}}+\frac{t^{2}+\sigma_{1}^{2}}{12N^{2}}\right).$
We thus have
$|\zeta(s)|$ $\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}+\frac{1}{64N^{4}}+\frac{1}{t}+\frac{t^{2}+\sigma_{1}^{2}+2(t+\sigma_{1})-1}{12N^{2}}\right).$
It remains to choose $N$. We can minimise the expression by taking $N=\lfloor\eta t\rfloor+1$ for some constant $\eta\in[2/t_{0},1-1/t_{0}]$. This will make
$\exp\left(\frac{\log N}{W\log t}\right)\leq\exp\left(\frac{1}{W}\right)$
for all $t\geq t_{0}$. The bound then becomes
$|\zeta(s)|$ $\leq e^{1/W}\left(\log t+\log(\eta+t^{-1})+\gamma+\frac{1}{t}\left(\frac{1}{\eta}+1\right)+\frac{1}{64\eta^{2}t^{2}}\right.$
$\hskip 142.26378pt+\left.\frac{1}{12\eta^{2}}+\frac{1}{6\eta^{2}t}+\frac{\sigma_{1}^{2}+2\sigma_{1}-1}{12\eta^{2}t^{2}}\right)$
$\leq e^{1/W}\log t+C_{0}(W,\sigma_{1},t_{0},\eta)$
for $t\geq t_{0}$. ∎
We also provide a corollary, specialised for $\sigma\geq 1$, which sharpens the above estimates.
Corollary 7.
Let $s=\sigma+ikt$ with $1\leq\sigma\leq\sigma_{1}\leq 2$, $t\geq t_{0}\geq 3$, and a fixed $k\geq 1$. For $0<\eta/k<t_{0}$, we have
$|\zeta(\sigma+ikt)|<\log kt+C(\sigma_{1},t_{0},k,\eta),$
where
$C(\sigma_{1},t_{0},k,\eta)=$ $\log\left(\frac{1}{\eta}+\frac{1}{kt_{0}}\right)+\gamma+\frac{\eta^{4}}{64(kt_{0})^{4}}$ (12)
$+\frac{12+(c_{1}(\sigma_{1},t_{0},k)-(kt_{0})^{-1})\eta^{2}}{12kt_{0}}+\frac{c_{2}(\sigma_{1},t_{0},k)\eta^{2}}{24},$
with $\gamma$ denoting Euler’s constant, and
$c_{1}(\sigma_{1},t_{0},k)$ $=\left(\frac{\sigma_{1}^{2}}{(kt_{0})^{2}}+1\right)^{1/2},$
$c_{2}(\sigma_{1},t_{0},k)$ $=c_{1}(\sigma_{1},t_{0},k)\left(\frac{(\sigma_{1}+1)^{2}}{(kt_{0})^{2}}+1\right)^{1/2}.$
Proof.
The proof follows the argument of [2, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. We omit the proof since it is almost identical to that of Lemma 5. ∎
To use Lemma 5, we want to choose a $t_{0}$ and optimise over $\eta$. For the smallest $t_{0}=3$, we can only use $\eta=2/3$. However, as $t_{0}\rightarrow\infty$, the optimal $\eta$ quickly approaches $1/\sqrt{6}\approx 0.4082$. We also have the following corollary, which is not required for our purposes, but is useful to have for other applications.
Corollary 8.
In the region
$1-\frac{1}{W_{0}\log t}\leq\sigma\leq 1,$ (13)
we have for $t\geq t_{0}\geq 3$,
$|\zeta(s)|\leq 1.1971\log t+B_{1},$ (14)
where Table 1 gives values of $B_{1}$ for select $t_{0}$.
Proof.
For a fixed $t_{0}$ and $\sigma_{1}=1$, in Lemma 5 optimise over $\eta$ to find the minimal $B_{1}$. ∎
      $t_{0}$       $B_{1}$       $\eta$
      $3$       $2.1173$       $2/3$
      $10^{3}$       $0.2254$       $0.41$
      $10^{12}$       $0.2171$       $0.41$
Table 1: Values of $B_{1}$ for Corollary 8 with the corresponding optimised $\eta$ from Lemma 5, where $\sigma_{1}=1$.
Corollary 8 can be used in conjunction with other estimates for $\zeta(s)$ in the critical strip, such as Ford’s Theorem 1 in [9] or Trudgian’s Lemma 2.1 [26]. Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$. For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 8 is an improvement to [9, Theorem 1]. This makes it useful for results over smaller $t$.
It is not uncommon to degrade asymptotically sharper results for $\zeta(s)$ to $|\zeta(s)|\leq C\log t$ for the sake of a smaller main-term constant $C$. Corollary 8 now directly offers a better constant. For instance, over $\sigma\in[1-1/(W_{0}\log t),1]$ and $t\geq 3$ our result implies
$|\zeta(s)|\leq 3.13\log t,$
whereas degrading a Richert-type bound like Ford’s [9, Theorem 1] would give
$|\zeta(s)|\leq 102\log t,$
and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [26, Lem. 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [14]) and estimates on the $1$-line (see Backlund [1]) would yield
$|\zeta(s)|\leq 364\log t.$
Next we have some bounds on $\zeta(\sigma)$ for real $\sigma>1$. These are useful in their own right, but we show they can be used to obtain good bounds on $-\zeta^{\prime}(\sigma)/\zeta(\sigma)$ in a similar region.
Lemma 6.
[20, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
$|\zeta(\sigma+it)|\leq\zeta(\sigma)\leq\frac{e^{\gamma(\sigma-1)}}{\sigma-1}.$
Lemma 7.
Let $\gamma_{n}$ be the Laurent–Stieltjes constants (with $\gamma:=\gamma_{0}$). For integers $k$ such that $1\leq k\leq 1500$, and real $\sigma$ satisfying $31<\sigma\leq\sigma_{k}<$ 3, we have
$\zeta(\sigma)\geq\frac{1}{\sigma-1}+\gamma+\phi_{0}(\sigma,k),$
where
$\phi_{0}(\sigma,k)=\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}}{n!}\gamma_{n}+\frac{1}{(\sigma-3)}\left(\frac{\sigma-1}{2}\right)^{k+1}\geq 0.$ (15)
In addition, we note some admissible pairs:
$(\sigma_{k},k)=(1.45,1),\,(2,3),\,(2.51,10),\,(2.97,500),\,(2.98,1000),\,(2.99,1500).$
Proof.
Recall from the Laurent series expansion of the zeta function
$\zeta(\sigma)=\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}+\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}.$ (16)
To prove the non-negativity of (15) we require the last two terms of (16) together to be positive. We have the bounds on $\gamma_{n}$ due to Lavrik [16]:
$|\gamma_{n}|\leq\frac{n!}{2^{n+1}},\quad\quad n=1,2,3,\ldots,$ (17)
hence
$\left|\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}\right|\leq\sum_{n=k+1}^{\infty}\frac{(\sigma-1)^{n}}{2^{n+1}}=-\frac{1}{(\sigma-3)}\left(\frac{\sigma-1}{2}\right)^{k+1}$
by the geometric series with $|1-\sigma|<2$. Taking the lower bound from this, one can then verify that $\phi_{0}(\sigma,k)\geq 0$ for $1<s\leq\sigma_{k}$ by calculating the first $k$ terms. ∎
It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [20, Lemma 5.3]. Our method in Lemmas 7 and 8, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes. It is worth noting that bounds sharper than (17) are available, but are more complicated to work with.
Lemma 8.
Let the hypotheses of Lemma 7 be fulfilled. Then for $\sigma$ real, we have
$-\zeta(\sigma)\leq\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}},\qquad 1<\sigma<3,$
and
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\frac{-\zeta^{\prime}(\sigma)}{\zeta(\sigma)}\leq\frac{\phi_{1}(\sigma,k)}{\sigma-1},\qquad 1<\sigma\leq\sigma_{k},$
where
$\phi_{1}(\sigma,k)=\frac{(\sigma-3)^{2}+(\sigma-1)^{2}}{(\sigma-3)^{2}(1+(\sigma-1)(\phi_{0}(\sigma,k)+\gamma))}.$ (18)
Proof.
The first inequality of the second assertion is trivial. The subsequent inequality is proved by differentiating (16) and repeating a process similar to the proof of Lemma 7 (not splitting the sum for brevity), to obtain
$-\zeta^{\prime}(\sigma)$ $\leq\frac{1}{(\sigma-1)^{2}}+\sum_{n=1}^{\infty}\frac{n(\sigma-1)^{n-1}}{2^{n+1}}$
$=\frac{1}{(\sigma-1)^{2}}+\frac{1}{2(\sigma-1)}\sum_{n=1}^{\infty}n\left(\frac{\sigma-1}{2}\right)^{n}=\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}}$
for $|1-\sigma|<2$ by identification with the polylogarithm $\textup{Li}_{-1}(z)=z/(1-z)^{2}$. Dividing by the bound in the Lemma 7 proves the result. ∎
These bounds are good for $\sigma$ close to $1$. For instance, we have that
$\phi_{1}(\sigma,3)<1\qquad\text{for}\qquad\sigma\leq 1.83,$ (19)
which is an improvement on Delange’s result [7] for a fixed $\sigma\leq 1.65$. Although the bounds get sharper as $k$ increases, in practice, taking $k\leq 10$ is more than sufficient for small $\sigma$ and any savings gained by subsequent increases in $k$ would be after the fifth decimal place.
Corollary 9.
Let $\phi_{1}(\sigma,k)$ be defined as in (18). For $1<\sigma_{0}\leq\sigma\leq 1.5$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\frac{-\zeta^{\prime}(\sigma)}{\zeta(\sigma)}\leq\frac{\phi_{2}(\sigma_{0})}{\sigma-1},$
[…]
Proof.
For $k=10$, we have from Lemma 16 that $\phi_{0}(\sigma,10)\geq 0$ in our stated range of $\sigma$. It can also be verified that $\phi_{1}(\sigma,10)$ is decreasing in the region $1<\sigma\leq 1.48$, and therefore $\phi_{1}(\sigma,10)\leq\phi_{1}(\sigma_{0},10)$.
To complete the proof, verify that $\phi_{1}(\sigma,10)\leq 0.851$ in the region $1.48\leq\sigma\leq 1.5$. ∎
Finally, we have a lower bound on $|\zeta(s)|$, for $\sigma>1$, which is better than the trivial bound $|\zeta(s)|\geq\zeta(2\sigma)/\zeta(\sigma)$ over certain ranges of $\sigma$ (for more information, see [17]). This is the $\log t$ analogue of [6, Lemma 8].
Lemma 9.
Let $\kappa,\,\sigma_{1}$ be positive parameters, and $\gamma$ be Euler’s constant. Then for $t\geq t_{0}\geq 3$ and $1+\kappa\leq\sigma_{1}\leq 2$, we have
$\left|\frac{1}{\zeta(1+\kappa+it)}\right|\leq V(\kappa,\sigma_{1},t_{0},\eta)\log t$
with
$V(\kappa,\sigma_{1},t_{0},\eta):=\left(\frac{\exp\left(\gamma\kappa\right)}{\kappa\log t_{0}}\right)^{3/4}\left(1+\frac{\log 2+C(\sigma_{1},t_{0},2,\eta)}{\log t_{0}}\right)^{1/4},$
where $\eta$ and $C(\sigma_{1},t_{0},k,\eta)$ are defined in Corollary 7.
Proof.
Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [24, § 3.3] and [17]). That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq|\zeta(\sigma)|^{3/4}|\zeta(\sigma+2it)|^{1/4}.$ (20)
Taking $\sigma=1+\kappa$ in (20) gives
$\left|\frac{1}{\zeta(1+\kappa+it)}\right|\leq|\zeta(1+\kappa)|^{3/4}|\zeta(1+\kappa+2it)|^{1/4}.$ (21)
The two factors on the right-hand side can be bounded with Lemmas 6 and 7, respectively, with the aid of the bound
$\log 2t$ $\leq\left(1+\frac{\log 2}{\log t_{0}}\right)\log t.$
This proves the lemma. ∎
4 Estimates for $|\zeta^{\prime}(s)/\zeta(s)|$
4.1 Assuming a partial Riemann hypothesis
In this section, we assume $RH(T)$, and estimate $|\zeta^{\prime}(s)/\zeta(s)|$ in a right-half plane: $\sigma\geq 1/2+\epsilon$, $t_{0}\leq t\leq T$.
We will first prove the result for $T_{0}$ and then make the substitution to $t$ at the end. The idea is to construct discs $|s-s_{0}|\leq 1/2$ which have origins slightly to the right of the $1$-line, centred at $s_{0}=\sigma_{0}+iT_{0}$, where
$\sigma_{0}=1+\epsilon,\qquad e+1/2\leq t_{0}\leq T_{0}\leq T-1/2,$
and $\epsilon\leq 1/2$ is a parameter to be determined later. We then extend their radii into the critical strip, so that Theorem 1 applied to $f(s):=-\zeta^{\prime}(s)/\zeta(s)$ gives us the desired estimates. Due to $RH(T)$, $\zeta(s)$ is analytic and non-zero in the discs.
Choosing a fixed $R^{\prime}<\epsilon$, we have by Corollary 5 that for $1/2+\epsilon\leq\sigma\leq 3/2+\epsilon$,
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT_{0})\right|\leq M:=\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+iT_{0}+R^{\prime}e^{i\theta})||,$ (22)
which we estimate, by Lemma 6 with $\delta_{r}=1/2+\epsilon$ and $T_{0}\geq t_{0}\geq e+R^{\prime}$. In the region $1/2\leq\sigma\leq 3/2+2\epsilon$, we have
$|\zeta(s)|\leq A_{0}(T_{0}+R^{\prime})^{1/6}\log(T_{0}+R^{\prime})\leq A_{1}T_{0}^{1/6}\log T_{0},$ (23)
where
$A_{1}$ $=A_{0}\left(1+\frac{R^{\prime}}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{t_{0}}\right)}{\log t_{0}}\right),$ (24)
with
$A_{0}$ $=0.618\left(1+a_{1}\left(\frac{3}{2}+2\epsilon,1.31,t_{0}-R^{\prime}\right)\right)^{7/6}$ (25)
×$\qquad\qquad\bm{\times}\left(1+a_{0}\left(\frac{3}{2}+2\epsilon,1.31,t_{0}-R^{\prime}\right)\right).$
Combining (23), (25), and (24) with (22) thus gives for $1/2+\epsilon\leq\sigma\leq 3/2+\epsilon$,
$M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{1}\right).$ (26)
Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we can use the trivial bound for $\sigma>1$ [11, Lemma 70.1], which gives us
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{1}{\sigma_{0}-1}.$ (27)
However, when bounding $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$ in an interval away from the $1$-line, that is to say, where $\sigma_{0}\leq\sigma\leq 1.5$ with $\sigma_{0}$ not tending to $1$, we can do better than the trivial bound by using Corollary 9. We have for $\sigma_{0}\leq\sigma\leq 1.5$,
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{\phi_{2}(\sigma_{0})}{\sigma_{0}-1}.$ (28)
Up to this point, the above with Theorem 1, where $M$ is as in (26), and $R=1/2$, will give a result that holds for
$1+\epsilon-\frac{\alpha}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha}{2},$ (29)
and for a result on any larger $\sigma$, we simply use the trivial bound in (27).
Finally, putting everything together and making the substitution $T_{0}\mapsto t$, we prove our result.
Theorem 2.
Assume $RH(T)$. Let $\alpha$, $\epsilon$, and $R^{\prime}$ be parameters, such that $20<R^{\prime}<\epsilon\leq 1/$ 2 and $10<\alpha<$ 1. In the region
$\sigma\geq 1+\epsilon-\frac{\alpha}{2}\quad\text{and}\quad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(T)}\log t,$
with
$Q_{RH(T)}$ $=\max\Bigg{\{}\frac{2\alpha}{(1-\alpha)\pi R^{\prime}}\Bigg{(}\frac{1}{6}+B_{t_{0}}+\frac{\log A_{1}}{\log t_{0}}\Bigg{)}$ (30)
$\qquad\qquad+\left(\frac{1+\alpha}{1-\alpha}\right)\frac{\phi_{2}(1+\epsilon)}{\epsilon\log t_{0}},\,\frac{1}{(\epsilon+(\alpha/2))\log t_{0}}\Bigg{\}},$
where $A_{1}$ is defined in (24), $\phi_{2}(\sigma)$ is defined in Corollary 9, and
$B_{t_{0}}=\begin{cases}\frac{1}{e}&\text{ if }t_{0}\leq e^{e},\\ \frac{\log\log t_{0}}{\log t_{0}}&\text{ if }t_{0}>e^{e}.\end{cases}$
Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
Assuming $RH(H_{0})$ allows us to prove Corollary 3, which holds for $t\leq H$.
Proof of Corollary 3.
For $t\geq t_{0}$, in Theorem 2 set $\sigma_{0}=1+\epsilon-\frac{\alpha}{2}$. Then fix $t_{0}$ and optimise (30) over $\epsilon,\,R^{\prime},\,\alpha$, while ensuring that the restrictions set in Theorem 2 and (9) hold. In practice $R^{\prime}$ is taken to be arbitrarily smaller than $\epsilon$, differing only after the sixth decimal place. The final value of $t_{0}=5$ is taken to be the one where (9) holds for all $\sigma_{0}$ in Table 2. This gives the result. ∎
$\sigma_{0}$ $Q_{RH(H_{0})}$ $\epsilon$ $\alpha$
$0.51$ $38889$ $0.004948$ $0.989897$
$0.55$ $1508$ $0.024334$ $0.948668$
$0.6$ $361.726$ $0.047605$ $0.89521$
$0.65$ $153.957$ $0.070599$ $0.841198$
$0.7$ $82.739$ $0.092279$ $0.784558$
$0.75$ $50.441$ $0.113238$ $0.726476$
$0.8$ $33.249$ $0.129407$ $0.658814$
$0.85$ $23.086$ $0.144234$ $0.588468$
$0.9$ $16.601$ $0.160246$ $0.520492$
$0.95$ $12.235$ $0.174136$ $0.448272$
$1$ $9.161$ $0.176533$ $0.353067$
Table 2: Values for $Q_{RH(H_{0})}$ and $\sigma\geq\sigma_{0}$ in Corollary 3, with corresponding $\epsilon,\,\alpha$ (all rounded to six decimal places). Each entry is valid for $5\leq t\leq H$.
4.2 Above the Riemann height
Suppose $t_{0}\geq H$ and let there be a positive real number $d$. As before, we construct discs which have origins slightly to the right of the $1$-line, at $s_{0}=\sigma_{0}+iT_{0}$, where $T_{0}\geq t_{0}$,
$\sigma_{0}=1+\delta_{T_{0}}=1+\frac{d}{\log T_{0}},$ (31)
and extend their radii slightly into the critical strip, and then apply Theorem 1. This time however, we do not have the assumption of $RH(T)$ to help us with potential zeros. We will instead utilise the zero-free region mentioned in Section 2.1 and keep our discs within said region. This requires discs with radii decreasing to 00 at the rate of $1/\log T_{0}$, similar to the zero-free region.
Let
$R_{T_{0}}=2\delta_{T_{0}}=\frac{2d}{\log T_{0}}\qquad\text{and}\qquad r=\alpha R_{T_{0}},\quad 0<\alpha<1.$ (32)
In addition, let $d$, and hence $R_{T_{0}}$, be chosen such that
$\sigma_{0}-R_{T_{0}}=\sigma_{0}-2\delta_{T_{0}}\geq 1-\frac{1}{W_{0}\log(T_{0}+R_{T_{0}})},$
which is fulfilled by
$d\leq\frac{1}{W_{0}}\left(1+\frac{\log\big{(}1+\frac{R_{H}}{H}\big{)}}{\log H}\right)^{-1}.$ (33)
This ensures that the disc $|s-s_{0}|\leq R_{T_{0}}$ lies entirely in the zero-free region, so $f(s)$ is analytic there.
Next, we want to obtain a bound on $M$. This is done by choosing
$R^{\prime}=\frac{1}{2}-\delta_{H}=\frac{1}{2}-\frac{d}{\log H}$
in Lemma 2. Then as long as $d$ is chosen to satisfy (1), i.e.,
$d\leq\frac{1}{W_{0}}\left(1+\frac{\log\big{(}1+\frac{R^{\prime}}{H}\big{)}}{\log H}\right)^{-1},$ (34)
we have for $\sigma\geq 1-\delta_{H}$,
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT_{0})\right|\leq M:=\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+iT_{0}+R^{\prime}e^{i\theta})||.$ (35)
Estimating $M$ thus requires a uniform bound of $\zeta(s)$ in the region
$\frac{1}{2}\leq 1-\delta_{T_{0}}-R^{\prime}\leq\sigma\leq 1+3\delta_{T_{0}}+R^{\prime}\leq 1+\frac{1}{2}+2\delta_{H}.$
By Lemma 6 with $\delta_{r}=1/2+2\delta_{H}$, we have for $T_{0}\geq H$, and the region $1/2\leq\sigma\leq 3/2+2\delta_{H}$ that
$|\zeta(s)|\leq A_{2}(T_{0}+R^{\prime})^{1/6}\log(T_{0}+R^{\prime})\leq A_{3}T_{0}^{1/6}\log T_{0},$ (36)
where
$A_{3}$ $=A_{2}\left(1+\frac{R^{\prime}}{H}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{H}\right)}{\log H}\right),$ (37)
with
$A_{2}$ $=0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{2d}{\log H},1.31,H-R^{\prime}\right)\right)^{7/6}$ (38)
×$\qquad\qquad\bm{\times}\left(1+a_{0}\left(\frac{3}{2}+\frac{2d}{\log H},1.31,H-R^{\prime}\right)\right).$
Combining (36), (38), and (37) with (35) thus gives for $1-\delta_{T_{0}}\leq\sigma\leq 1+3\delta_{T_{0}}$,
$M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{3}\right).$ (39)
Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we use the trivial bound for $\sigma>1$, given in (27). This yields
$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq\frac{1}{\sigma_{0}-1}\leq\frac{\log T_{0}}{d}.$ (40)
Up to this point, Theorem 1 will give a result that holds for
$1-\delta_{T_{0}}(2\alpha-1)\leq\sigma\leq 1+\delta_{T_{0}}(2\alpha+1),$ (41)
and for a result on any larger $\sigma$, we simply use (27). At the same time, if we want a result that holds for $\sigma\geq 1-1/W\log T_{0}$, then by the left inequality of (41), we must have
$W=\frac{1}{d(2\alpha-1)}\qquad\text{and}\qquad\alpha=\frac{1}{2}+\frac{1}{2dW}.$ (42)
Finally, putting everything together and making the substitution $T_{0}\mapsto t$, we prove our result.
Theorem 3.
Let $d$ and $W\geq W_{0}$ be positive parameters, subject to (33), (34), and (42). Let also $10<\alpha<$ 1 be a parameter defined in (42), and depending on $d$ and $W$. In a region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H$
where $\zeta(\sigma+it)\neq 0$, we have
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{H}\log t,$
where
$Q_{H}=\max\left\{\frac{\lambda}{6}+\frac{1+\alpha}{d(1-\alpha)}+\frac{\lambda(\log\log H+\log A_{3})}{\log H},\,\frac{1}{d(1+2\alpha)}\right\},$
with $A_{3}$ defined in (37),
$\lambda=\frac{8\alpha}{(1-\alpha)\pi R^{\prime}},\qquad\text{and}\qquad R^{\prime}=\frac{1}{2}-\frac{d}{\log H}.$
4.3 Combining both cases $t\leq H$ and $t\geq H$
Proof of Corollary 1.
For the case of $t\geq H$, with a fixed choice of $W$ as defined in (42), it remains to optimise $Q_{H}$ in Theorem 3 over $\alpha$ and $d$, subject to the conditions mentioned in said theorem, while also checking that (9) holds. We found that the optimal value for $d$ in each case was approximately $d=0.179898\ldots$, and any differences in the final result caused by varying $d$ further occurred only after the sixth decimal place. We note also that the optimal $d$ is one that is as close to $1/W_{0}$ as possible while still satisfying (33) and (34).
On the other hand, a simple calculation shows that
$0.821-\frac{1}{W_{0}\log t}>$ 0.82
for $t\geq t_{0}=e$. Now, for $\sigma\geq 0.8$ and $5\leq t\leq H$, Corollary 3 and Table 2 asserts that
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=33.249\log t,$
which is smaller than any value of optimised $Q_{H}$ found. Thus this same value of $Q_{H}$ holds for all $t\geq 5$. ∎
We also prove Corollary 2, especially relevant for $\sigma=1$.
Proof of Corollary 2.
First note that Corollary 3 and Table 2 has for $\sigma\geq 1$ and $5\leq t\leq H$,
$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=9.161\log t.$
Then for $t\geq H$, the proof is identical to that of Theorem 3 and Corollary 1, except we can restrict the size of $\alpha$ even further, since we now require $\sigma_{0}-\alpha R_{T_{0}}\geq 1$, which is equivalent to $\alpha\leq 1/2$. In order for the result to hold for all $\sigma\geq 1$, we choose $\alpha=1/2$, so in this case,
$Q_{H}=\max\left\{\frac{8}{6\pi R^{\prime}}+\frac{3}{d}+\frac{8(\log\log H+\log A_{3})}{\pi R^{\prime}\log H},\,\frac{1}{2d}\right\}.$
As before, we then optimise over $d$, while ensuring that we fulfil the conditions in Theorem 3 and (9). The optimal value found was $d=0.179898$. ∎
$W$ $Q$ $\alpha$ $W$ $Q$ $\alpha$
$5.559$ $450041$ $0.999972$ $5.59$ $4458$ $0.997200$
$5.56$ $106223$ $0.999882$ $5.6$ $3383$ $0.996312$
$5.5601$ $98674$ $0.999873$ $5.7$ $1002$ $0.987604$
$5.5602$ $92137$ $0.999864$ $5.8$ $594.087$ $0.979198$
$5.5603$ $86414$ $0.999855$ $5.9$ $425.318$ $0.971076$
$5.5604$ $81354$ $0.999846$ $6$ $333.034$ $0.963224$
$5.5605$ $76859$ $0.999837$ $6.1$ $274.847$ $0.955630$
$5.561$ $60223$ $0.999792$ $6.2$ $234.806$ $0.948282$
$5.562$ $42026$ $0.999703$ $6.3$ $205.568$ $0.941166$
$5.563$ $32278$ $0.999613$ $6.4$ $183.281$ $0.934273$
$5.564$ $26201$ $0.999523$ $6.5$ $165.729$ $0.927592$
$5.565$ $22051$ $0.999433$ $6.75$ $134.741$ $0.911755$
$5.566$ $19036$ $0.999343$ $7$ $114.502$ $0.897049$
$5.567$ $16748$ $0.999254$ $8$ $74.999$ $0.847418$
$5.568$ $14951$ $0.999164$ $9$ $58.454$ $0.808816$
$5.569$ $13502$ $0.999074$ $10$ $49.359$ $0.777935$
$5.57$ $12310$ $0.998985$ $11$ $43.607$ $0.752668$
$5.575$ $8541$ $0.998537$ $12$ $39.642$ $0.731612$
$5.58$ $6542$ $0.998091$ $13$ $36.742$ $0.713796$
Table 3: Values for $Q$ in Corollary 1, with corresponding $W$ and $\alpha$ (rounded to six decimal places). Each entry is valid for $t\geq 5$.
5 Bounds on $|1/\zeta(s)|$
Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [24, Theorem 3.11]). It revolves around estimating the right-hand side of
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|=-\textup{Re}\log\zeta\left(1+\delta_{1}+it\right)+\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x$ (43)
for some $\delta_{1}>0$.
The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [2, Proposition A.2], [17]), while in the past the above integral has been estimated by bounding the real part of the logarithmic derivative by a uniform bound $|\zeta^{\prime}(s)/\zeta(s)|$. Here some savings can be made by splitting up the interval $[\sigma,1+\delta_{1}]$ and using the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval (see [6, Lemma 10]).
However, we show that by using directly a bound on the real part instead, through Lemma 2, we are able to do much better. As before, the proof will be split into two cases, $t\leq H$ and $t\geq H$.
5.1 Assuming a partial Riemann hypothesis
Theorem 4.
Assume $RH(T)$. Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log t_{0}\leq\sigma_{1}\leq 2$. Suppose also $R^{\prime}=1/2-d/\log t_{0}$. Then in the region
$\sigma\geq 1-\frac{1}{W\log t}\qquad\text{and}\qquad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{RH(T)}(d,d_{1},\sigma_{1},t_{0},\eta)\log t,$
where
$Y_{RH(T)}$ $(d,d_{1},\sigma_{1},t_{0},\eta)=\max\Bigg{\{}\frac{1}{d_{1}}+\frac{1}{\log t_{0}},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\Bigg{\}},$
with
$A_{4}$ $=\left(1+\frac{R^{\prime}}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{t_{0}}\right)}{\log t_{0}}\right)\left(1+a_{0}\left(\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},1.31,t_{0}-R^{\prime}\right)\right)$
×$\qquad\qquad\bm{\times}0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},1.31,t_{0}-R^{\prime}\right)\right)^{7/6},$
$B_{t_{0}}$ defined in Theorem 2, and $V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta),\,\eta$ defined in Lemma 9.
Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
Proof.
Let $\delta_{t}:=d/\log t$ and $\delta_{1,t}:=d_{1}/\log t$, and first suppose $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. Since by $RH(T)$ there are no zeros for $\sigma\geq 1/2$, we do not have to check the condition in (1). The same argument in Section 4.1 with Lemma 2 leading to (26), requires a uniform bound on $|\zeta(s)|$ in the region
$\frac{1}{2}\leq 1-\delta_{t}-R^{\prime}\leq\sigma\leq 1+\delta_{1,t}+R^{\prime}\leq\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},$
and in return gives us
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\log t$
uniformly for $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. Thus we have from (43) that
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\log\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|+\int_{1-\delta_{t}}^{1+\delta_{1,t}}\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(x+it)\right|\text{d}x,$ (44)
where the integral is bounded from above by
$\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right).$
To bound the first term on the right-hand side of (44), we use Lemma 9 with $\kappa=\delta_{1,t}=d_{1}/\log t$ to see that
$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|\leq V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$
Hence (44) becomes
$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$
for $t_{0}\leq t\leq T-1/2$ and $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
Now we turn to the case $\sigma\geq 1+\delta_{1,t}$. Here we use the trivial bound $|\zeta(s)|\leq\zeta(\sigma)\leq\sigma/(\sigma-1)$, which is decreasing for all $\sigma>1$. Together with the trigonometric inequality from (21), we see that
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\zeta(1+\delta_{1,t})\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t.$
We reiterate that although there are sharper bounds available to use here, this is sufficient for our purposes as the case where $\sigma\leq 1+\delta_{1,t}$ is dominant when computing constants.
Note that the results proved hold also if one assumes $RH$, and allow us to remove the upper bound restriction on $t$, as explained at the end of the proof of Theorem 2. ∎
5.2 Above the Riemann height
Now we consider $t\geq H$ and prove the following.
Theorem 5.
Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log H\leq\sigma_{1}\leq 2$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W\geq W_{0}$ is chosen such that $d$ satisfies (34). Then in the region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{H}(d,d_{1},\sigma_{1},\eta)\log t,$
where
$Y_{H}$ $(d,d_{1},\sigma_{1},\eta)=\max\Bigg{\{}\frac{1}{d_{1}}+\frac{1}{\log H},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}V\left(\frac{d_{1}}{\log H},\sigma_{1},H,\eta\right)\Bigg{\}},$
with
$A_{5}$ $=\left(1+\frac{R^{\prime}}{H}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{H}\right)}{\log H}\right)\left(1+a_{0}\left(\frac{3}{2}+\frac{d_{1}-d}{\log H},1.31,H-R^{\prime}\right)\right)$
×$\qquad\qquad\bm{\times}0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{d_{1}-d}{\log H},1.31,H-R^{\prime}\right)\right)^{7/6},$
and $V(d_{1}/\log H,\sigma_{1},H,\eta),\,\eta$ are defined in Lemma 9. Admissible values of $Y_{H}$ with corresponding $W$ are presented in Table 4.
Proof.
The proof is analogous to that of Theorem 4. Let $\delta_{t}:=d/\log t$ and $\delta_{1,t}:=d_{1}/\log t$, and first suppose $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. As we do not assume $RH(T)$, we have to check that (1) holds, and we do this by our hypothesis of $d$ satisfying (34). The same argument in Section 4.2 with Lemma 2 leading to (39), requires a uniform bound on $|\zeta(s)|$ in the region
$\frac{1}{2}\leq 1-\delta_{t}-R^{\prime}\leq\sigma\leq 1+\delta_{1,t}+R^{\prime}\leq\frac{3}{2}+\frac{d_{1}-d}{\log H},$
and in return gives us
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\log t$
uniformly for $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
The rest of the proof then follows that of Theorem 4, and it remains to optimise $Y_{H}$ for a fixed choice of $W=1/d$, over $d_{1}$, $\sigma_{1}$ and $\eta$, subject to the conditions mentioned in the statement of the theorem, while also checking that (9) holds. ∎
$W$ $Y_{H}$ $d_{1}$ $\sigma_{1}$ $\eta$
$W_{0}$ $2.307$ $1.059275$ $1.608256$ $3.775265$
$6$ $2.284$ $1.060503$ $1.608124$ $3.775135$
$7$ $2.244$ $1.062065$ $1.604395$ $3.772048$
$8$ $2.214$ $1.063547$ $1.895304$ $3.621918$
$9$ $2.191$ $1.064570$ $1.399356$ $3.464227$
$10$ $2.173$ $1.065554$ $1.135564$ $3.65928$
$11$ $2.158$ $1.065999$ $1.43905$ $3.446264$
$12$ $2.146$ $1.066605$ $1.345521$ $3.769427$
$13$ $2.136$ $1.067103$ $1.345557$ $3.77319$
Table 4: Values for $Y_{H}$ in Theorem 5, with corresponding $W$, and parameters (rounded to six decimal places) $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq H$.
5.3 Combining both cases $t\leq H$ and $t\geq H$
Proof of Corollary 4.
For the first assertion of the corollary, we begin by assuming $RH(H_{0})$ and dealing with the case $t_{0}\leq t\leq H$.
To obtain $Y(W,5)$, fix $t_{0}$ and $W=1/d$ in Theorem 4 and optimise $Y_{RH(H_{0})}$ over $d_{1}$, $\sigma_{1}$, and $\eta$. At the same time ensure that the restrictions set in Theorem 4 and Corollary 7 hold, and raise $t_{0}$ if (9) is not satisfied. We find that the smallest value of $t_{0}$ such that (9) is fulfilled for all entries in Tables 6 and 5 is $t_{0}=5$. Then we compare the resulting values with the corresponding $Y_{H}$ (for $t\geq H$) in Theorem 5 and Table 4, taking the larger of the two.
[…]
For the second assertion, i.e., when $\sigma=1$, we set $d=0$ and repeat the above process using Theorems 4 (under $RH(H_{0})$) and 5 (unconditionally). This allows us to obtain a result for $\sigma\geq 1$. We have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 2.018\log t\quad\text{for}\quad t\geq H,$ (45)
[…]
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 2.751\log t\quad\text{for}\quad 2\exp(e^{2})\leq t\leq H,$ (46)
[…]
Finally, combine (45) and (46) with the following two bounds. First, we utilise Lemma 4, which was proven using numerical methods. Taking $\sigma=1$ in said lemma, we derive
$\left|\frac{1}{\zeta(1+it)}\right|\leq\frac{8}{\log(500)}\log t\leq 1.288\log t\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}).$
Second, a computation from the proof of [2, Proposition A.2], implemented in interval arithmetic, states
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.079\log t\quad\text{for}\quad 2\leq t\leq 500.$
This completes the proof. ∎
When $\sigma$ is fixed (in this case $\sigma=1$), the limiting factor for the constant in the second assertion of Corollary 4 comes from the case $2\exp(e^{2})\leq t\leq H$. In particular, the lower threshold of $2\exp(e^{2})$ is what determines the constant $2.751$. Using numerical methods, it is possible to push this threshold higher and hence improve (46). To illustrate this idea, following the proof of Corollary 4, we can prove unconditionally that for $t\geq 10^{6}$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.391\log t,$ (47)
which is an improvement over what one gets from [4, Theorem 1.1].
We make the remark that we include our computed values of $\sigma_{1}$ and $\eta$ in Tables 4, 5, and 6, for completeness sake. In practice, as we limit the overall constant to three decimal places, $d_{1}$ is the main parameter we optimise over. For $\sigma_{1}$ and $\eta$, any sufficiently large values will do, and further optimisations here will only improve the final constant from the fourth decimal place onwards.
$t_{0}$ $Y(W_{0},t_{0})$ $d_{1}$ $\sigma_{1}$ $\eta$
$10$ $8.101$ $0.307574$ $1.133578$ $3.071916$
$10^{2}$ $4.339$ $0.542834$ $1.388735$ $3.417735$
$10^{3}$ $3.632$ $0.648794$ $1.470345$ $3.459386$
$10^{4}$ $3.264$ $0.725027$ $1.834773$ $3.463610$
$10^{5}$ $3.021$ $0.787051$ $1.447903$ $3.464020$
$10^{7}$ $2.711$ $0.885013$ $1.30488$ $3.464071$
$10^{9}$ $2.518$ $0.960532$ $1.322282$ $3.464052$
$10^{11}$ $2.383$ $1.021393$ $1.511713$ $3.464182$
Table 5: Values for $Y(W_{0},t_{0})$ in Corollary 4, with corresponding $t_{0}$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq t_{0}$ and $W=W_{0}$.
$W$ $Y(W,5)$ $d_{1}$ $\sigma_{1}$ $\eta$
$W_{0}$ $18.060$ $0.170011$ $1.105634$ $2.789282$
$6$ $16.722$ $0.173040$ $1.107519$ $2.781151$
$7$ $14.654$ $0.178370$ $1.110828$ $2.783844$
$8$ $13.339$ $0.182334$ $1.113292$ $2.786533$
$9$ $12.433$ $0.185403$ $1.115198$ $2.782321$
$10$ $11.771$ $0.187839$ $1.116712$ $2.782458$
$11$ $11.268$ $0.189804$ $1.117933$ $2.786745$
$12$ $10.872$ $0.191448$ $1.118954$ $2.786457$
$13$ $10.553$ $0.192838$ $1.119819$ $2.786611$
Table 6: Values for $Y(W,5)$ in Corollary 4, with corresponding $W$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq t_{0}=5$.
6 Concluding Remarks
Here we state some possible ways to improve our estimates. The most impactful would be having a larger zero-free region or a higher Riemann height. Using an asymptotically stronger zero-free region is also helpful. See for instance [6], where a Littlewood-style zero-free region is used to obtain bounds of order $\log t/\log\log t$. Minor improvements can be made by a more careful treatment of the sum over zeros and the integral arising from the application of Lemma 1. Other possibilities are: an improvement on the Borel–Carathéodory by specialising it more to this problem, proving a uniform (in $\sigma$) bound for $\arg\zeta(s)$, or using stronger estimates for $\zeta(s)$ under $RH$ or $RH(T)$.
It is evident from the results that the closer $s$ is to the boundary of the zero-free region, and hence a potential zero, the larger the constants. However, the large constants can also be attributed in part to the factor $(R+r)/(R-r)$ in Theorem 1, which increases rapidly as $r$ approaches $R$. This contributes significantly to the overall estimate, even if $|f(s_{0})|$ is small, and seems to suggest that Theorem 1 in its current form might not be optimal for this problem.
The main advantage of our method using Lemma 1 is that it eliminates the need for Theorem 1 in estimating $|1/\zeta(s)|$, and also reduces the number of applications of the theorem when estimating $|\zeta^{\prime}(s)/\zeta(s)|$. It might be possible to remove this entirely, by directly bounding the real and imaginary part of $|\zeta^{\prime}(s)/\zeta(s)|$, as one can easily prove an analogue of Lemma 1 which relates $\textup{Im}(f^{\prime}/f)$ to $\arg f$. To use such a result in the case of $f(s)=\zeta(s)$, besides dealing with the corresponding sum over zeros, a uniform bound (over a range of $\sigma$) on $S_{\sigma}(t):=(1/\pi)\arg\zeta(\sigma+it)$ is needed.
To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results. As an example, for $W=W_{0}$, the condition (34) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$. The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us. In fact we only need a ‘weak’ version of $RH$, for instance in (22), that up to height $T$, no zeros lie to the right of $1/2-R^{\prime}$. The downside to this is that we have not used better available bounds that require the full strength of $RH$. For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [24, §14], [21], [22], and [5] for more details).
Acknowledgements
We thank Tim Trudgian for his guidance throughout the writing of this paper. We are also very grateful to Bryce Kerr, Olivier Ramaré, and Aleksander Simonič for the helpful discussions.

### v2

@@
 These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s.
+Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
 1 Introduction
@@
 For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function.
-The reader is referred to [25], [3], [8], [24, (3.14)] for more information on the topic.
-Trudgian in [25] gives an explicit version of results in [24, §3] originally due to Landau, obtaining bounds of order $O(\log t)$.
+The reader is referred to [26], [4], [9], [25, (3.14)] for more information on the topic.
+Trudgian in [26] gives an explicit version of results in [25, §3] originally due to Landau, obtaining bounds of order $O(\log t)$.
 This was then extended in [14] and [7] to bounds of order $O(\log t/\log\log t)$.
@@
 This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
-The method in [24, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$.
+The method in [25, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$.
 It relies on two applications of the Borel–Carathédory theorem.
 The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$.
-As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [12, Lemma 3.2].
+As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [13, Lemma 3.2].
 While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [26], we take the opportunity to point out an error in that paper.
-In the quantity $A$ of [25, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$.
+In the quantity $A$ of [26, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$.
 This leads to the overall constants being better than they should be.
@@
 Furthermore, we directly derive a bound on $|1/\zeta(s)|$ from ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ rather than $|\zeta^{\prime}(s)/\zeta(s)|$, eliminating the need for the Borel–Carathédory theorem in this case, leading to a substantial improvement over [26].
-In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [2, Proposition A.2].
+In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [3, Proposition A.2].
 2 Main results
@@
 $\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y(W,t_{0})\log t,$
-where $Y(W_{0},10^{2})=4.339$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
+where $Y(W_{0},10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
 In addition, we have for $\sigma=1$ and $t\geq 2$,
 $\left|\frac{1}{\zeta(1+it)}\right|\leq 2.751\log t.$
+Corollary 5.
+Let $W\geq W_{0}$. In a region
+$\sigma\geq 1-\frac{1}{W\log t}$
+where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(W,t_{0})(\log t)^{11/12},$
+where $Y^{\prime}(W_{0},10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
+In addition, we have for $\sigma=1$ and $t\geq 2$,
+$\left|\frac{1}{\zeta(1+it)}\right|\leq 7.475(\log t)^{11/12}.$
 3 Some preliminaries
@@
 Theorem 1.
-[13, Theorem 4][Borel–Carathéodory] Let $s_{0}$ be a complex number, and let $R$ be a positive number possibly dependent on $s_{0}$.
+[14, Theorem 4][Borel–Carathéodory] Let $s_{0}$ be a complex number, and let $R$ be a positive number possibly dependent on $s_{0}$.
 Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$.
@@
 $|f(s)|\leq\frac{2r}{R-r}M+\frac{R+r}{R-r}|f(s_{0})|.$
-Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [12, Lemma 3.2].
-The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [23, §3.71]).
+Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [13, Lemma 3.2].
+The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [24, §3.71]).
 Lemma 1.
-[12, Lemma 3.2] Let $f(z)$ be holomorphic for $|z-a|\leq R$, and non-vanishing both at $z=a$ and on the circle $|z-a|=R$.
+[13, Lemma 3.2] Let $f(z)$ be holomorphic for $|z-a|\leq R$, and non-vanishing both at $z=a$ and on the circle $|z-a|=R$.
 Let $\rho_{j}=a+r_{j}\exp(i\theta_{j})$ be the zeros of $f(z)$ in the disc, where $\rho_{j}$ has multiplicity $n_{j}$.
@@
 The proof is analogous to that of Lemma 2 with the additional assumption that the disc $|s_{0}-\rho|\leq R^{\prime}$ is completely free of zeros. ∎
-Next we require some uniform bounds for $\zeta(s)$.
-We slightly improve [26, Corollary 1], which is proven using the Phragmén–Lindelöf principle.
-Note that we also use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$.
-This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing in $\sigma$.
-Additionally, we found that further splitting the interval complicates (8) and raises the threshold of $t$ for the lemma to be valid.
+Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [27, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
+Lemma 3.
+[28, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
+$|f(s)|<C\exp(e^{k|t|})$
+for some $C>0$ and $0<k<\tfrac{\pi}{b-a}$. Suppose further that there are $A,B,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\geq 0$ such that $\alpha_{1}\geq\beta_{1}$ and
+$|f(s)|\leq\begin{cases}A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|\quad\text{for}\textup{Re}(s)=a;\\ B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|\quad\text{for }\textup{Re}(s)=b.\end{cases}$
+Then for $a\leq\textup{Re}(s)\leq b$, one has
+$|f(s)|\leq\Big{(}{A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|}\Big{)}^{\tfrac{b-\textup{Re}(s)}{b-a}}\Big{(}{B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|}\Big{)}^{\tfrac{\textup{Re}(s)-a}{b-a}}.$
 Lemma 4.
@@
 Proof.
-The proof is identical to that of [26, Lemma 2.1], except that we are more precise in estimating (7) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$. ∎
+The proof is analogous to that of [27, Lemma 2.1], except that we are more precise in estimating (7) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
+Note that different to the proof of [27, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$.
+This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing (by the condition (8)) in $\sigma$.
+Additionally, we found that further splitting the interval complicates (8) and raises the threshold of $t$ for the lemma to be valid.
+∎
 Corollary 7.
@@
 Recall that Hiary, Patel, and Yang [15] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$.
-We now refer to Lemma 3, taking the maximum of (7) at $\sigma=1/2$. It has been shown in [27, p. 15] that $Q_{0}=1.31$ is permissible in (6) together with
+We now refer to Lemma 4, taking the maximum of (7) at $\sigma=1/2$. It has been shown in [29, p. 15] that $Q_{0}=1.31$ is permissible in (6) together with
 $(k_{1},k_{2},k_{3},Q_{0})=\left(0.618,\tfrac{1}{6},1,1.31\right),$
 hence we are done. ∎
+Lemma 5.
+For every $|t|\geq 3$, we have
+$|\zeta(1+it)|\leq 58.096(\log|t|)^{2/3}.$
+Proof.
+We use same argument in the first paragraph of [28, p.
+2], which modifies Ford’s proof in [10].
+The only difference is that we now apply the argument to [2, §5], which itself is an improvement on Ford’s result.
+This proves the lemma.
+∎
+Lemma 6.
+For $t\geq t_{0}\geq 3$ and $\sigma\geq 1$, we have
+$|\zeta(\sigma+it)|\leq C_{3}(\log t)^{2/3},$
+with
+$C_{3}=58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right).$ (10)
+Proof.
+The proof is similar to that of [14, Lemma 7]. First, to show that the result holds for $1\leq\sigma\leq 2$, we use the Phragmén–Lindelöf Principle, i.e., Lemma 3, on the holomorphic function
+$f(s)=(s-1)\zeta(s).$
+To this end, on the $1$-line we have
+$|f(1+it)|=|t|\,|\zeta(1+it)|\leq 58.096\,|2+it|(\log|2+it|)^{2/3}.$
+This inequality is verified numerically for $|t|<3$, and is a consequence of Lemma 5 for $|t|\geq 3$.
+Also, on the line $s=2+it$,
+$|f(2+it)|\leq|1+it|\,\zeta(2)<58.096|3+it|(\log|3+it|)^{2/3},$
+for all real $t$. Now apply Lemma 3 with the following parameter values: $a=1$, $b=2$, $Q=1$, $\alpha_{1}=\beta_{1}=1$, $\alpha_{2}=\beta_{2}=2/3$, and $A=B=58.096$. We obtain
+$|f(s)|\leq 58.096|1+s|(\log|1+s|)^{2/3},$ (11)
+for $1\leq\sigma\leq 2$. Moreover, for the same range of $\sigma$, we have the following estimates:
+$\begin{split}\frac{|1+s|}{|s-1|}&\leq\frac{\sqrt{9+t^{2}}}{t}\leq\sqrt{1+9t_{0}^{-2}},\\ \log|1+s|&\leq\log\sqrt{9+t^{2}}\leq\log t+\log\sqrt{1+9t_{0}^{-2}}.\end{split}$
+Thus for $1\leq\sigma\leq 2$ and $t\geq t_{0}$, the inequality (11) implies that
+$|\zeta(\sigma+it)|$ $\leq 58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right)\log^{2/3}t<C_{3}\log^{2/3}t,$
+as desired, where $C_{3}$ is defined in (10).
+Finally, we address the case where $\sigma\geq 2$. Here since $\sigma>1$, we apply the trivial bound in (48). Then for $\sigma\geq 2$, it is easy to verify that
+$|\zeta(\sigma+it)|\leq 2\leq C_{3}(\log t)^{2/3}$
+for all $t\geq t_{0}$. ∎
+Lemma 7.
+[22, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\frac{4}{\sigma-(1/2)}.$
 In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones.
@@
 However, we include them here as the proofs are similar to that of Corollary 8, and at the same time we provide a brief discussion on their usefulness.
-Lemma 5 is an improvement on [15, Proposition 5.3.2], which proves a similar result.
-Lemma 4.
-[22, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\frac{4}{\sigma-(1/2)}.$
+Lemma 8 is an improvement on [16, Proposition 5.3.2], which proves a similar result.
 Lemma 8.
@@
 where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$.
-Note that (10) holds for $\sigma>0$ by analytic continuation (see [24, Theorem 3.5]).
+Note that (12) holds for $\sigma>0$ by analytic continuation (see [25, Theorem 3.5]).
 In the region of $\sigma$ considered, we have
 $|N^{-s}|=N^{-\sigma}\leq\exp\left(-\Big{(}1-\frac{1}{W\log t}\Big{)}\log N\right)=N^{-1}\exp\left(\frac{\log N}{W\log t}\right).$ (13)
-This bound can be used throughout (10). The first term in (10) can also be bounded using the estimate in [10, Lemma 2.1], so for $N\geq 2$,
+This bound can be used throughout (12). The first term in (12) can also be bounded using the estimate in [11, Lemma 2.1], so for $N\geq 2$,
 $\left|\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}\right|$ $=\left|\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{1}{2N^{s}}\right|$
@@
 Proof.
-The proof follows the argument of [2, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. We omit the proof since it is almost identical to that of Lemma 5. ∎
+The proof follows the argument of [3, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. We omit the proof since it is almost identical to that of Lemma 8. ∎
 To use Lemma 8, we want to choose a $t_{0}$ and optimise over $\eta$.
@@
 Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$.
-For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 8 is an improvement to [9, Theorem 1].
+For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 9 is an improvement to [10, Theorem 1].
 This makes it useful for results over smaller $t$.
@@
 $|\zeta(s)|\leq 3.13\log t,$
-whereas degrading a Richert-type bound like Ford’s [9, Theorem 1] would give
+whereas degrading a Richert-type bound like Ford’s [10, Theorem 1] would give
 $|\zeta(s)|\leq 102\log t,$
-and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [26, Lem.
+and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [27, Lem.
 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [15]) and estimates on the $1$-line (see Backlund [1]) would yield
 $|\zeta(s)|\leq 364\log t.$
+We also remark that although Lemma 6 is asymptotically superior to Corollary 8, one would expect that the former beats the latter only for a sufficiently large $t$; approximately $t\geq\exp(58^{3})\geq 10^{84736}$, say.
+This can be improved by lowering the constant in Lemma 5, and correspondingly $C_{3}$.
 Next we have some bounds on $\zeta(\sigma)$ for real $\sigma>1$.
@@
 Lemma 9.
-[20, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
+[21, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
 $|\zeta(\sigma+it)|\leq\zeta(\sigma)\leq\frac{e^{\gamma(\sigma-1)}}{\sigma-1}.$
@@
 by the geometric series with $|1-\sigma|<2$. Taking the lower bound from this, one can then verify that $\phi_{0}(\sigma,k)\geq 0$ for $1<s\leq\sigma_{k}$ by calculating the first $k$ terms. ∎
-It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [20, Lemma 5.3].
-Our method in Lemmas 7 and 8, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes.
+It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [21, Lemma 5.3].
+Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes.
 It is worth noting that bounds sharper than (19) are available, but are more complicated to work with.
@@
 Finally, we have a lower bound on $|\zeta(s)|$, for $\sigma>1$, which is better than the trivial bound $|\zeta(s)|\geq\zeta(2\sigma)/\zeta(\sigma)$ over certain ranges of $\sigma$ (for more information, see [18]).
-This is the $\log t$ analogue of [6, Lemma 8].
+This is the $\log t$ analogue of [7, Lemma 8].
 Lemma 12.
@@
 Proof.
-Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [24, § 3.3] and [17]).
+Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [25, § 3.3] and [18]).
 That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
@@
 $\left|\frac{1}{\zeta(1+\kappa+it)}\right|\leq|\zeta(1+\kappa)|^{3/4}|\zeta(1+\kappa+2it)|^{1/4}.$ (23)
-The two factors on the right-hand side can be bounded with Lemmas 6 and 7, respectively, with the aid of the bound
+The two factors on the right-hand side can be bounded with Lemmas 9 and 8, respectively, with the aid of the bound
 $\log 2t$ $\leq\left(1+\frac{\log 2}{\log t_{0}}\right)\log t.$
@@
 $M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{1}\right).$ (28)
-Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we can use the trivial bound for $\sigma>1$ [11, Lemma 70.1], which gives us
+Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we can use the trivial bound for $\sigma>1$ [12, Lemma 70.1], which gives us
 $\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{1}{\sigma_{0}-1}.$ (29)
@@
 5 Bounds on $|1/\zeta(s)|$
-Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [24, Theorem 3.11]). It revolves around estimating the right-hand side of
+Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [25, Theorem 3.11]). It revolves around estimating the right-hand side of
 $\log\left|\frac{1}{\zeta(\sigma+it)}\right|=-\textup{Re}\log\zeta\left(1+\delta_{1}+it\right)+\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x$ (45)
 for some $\delta_{1}>0$.
-The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [2, Proposition A.2], [17]), while in the past the above integral has been estimated by bounding the real part of the logarithmic derivative by a uniform bound $|\zeta^{\prime}(s)/\zeta(s)|$.
-Here some savings can be made by splitting up the interval $[\sigma,1+\delta_{1}]$ and using the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval (see [6, Lemma 10]).
-However, we show that by using directly a bound on the real part instead, through Lemma 2, we are able to do much better. As before, the proof will be split into two cases, $t\leq H$ and $t\geq H$.
+The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [3, Proposition A.2], [18]), while in the past the above integral has been estimated by bounding the real part of the logarithmic derivative by a uniform bound $|\zeta^{\prime}(s)/\zeta(s)|$.
+Here some savings can be made by splitting up the interval $[\sigma,1+\delta_{1}]$ and using the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval (see [7, Lemma 10]).
+However, we show that by using directly a bound on the real part instead, through Lemma 2, we are able to do much better.
+Furthermore, through a trick utilising the trigonometric polynomial, we are able to get an slight asymptotic improvement on the bounds, at the cost of a slightly worse constant.
+As before, the proofs will be split into two cases, $t\leq H$ and $t\geq H$.
 5.1 Assuming a partial Riemann hypothesis
@@
 To bound the first term on the right-hand side of (46), we use Lemma 12 with $\kappa=\delta_{1,t}=d_{1}/\log t$ to see that
-$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|\leq V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$
+$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|\leq V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$ (47)
 Hence (46) becomes
@@
 for $t_{0}\leq t\leq T-1/2$ and $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
-Now we turn to the case $\sigma\geq 1+\delta_{1,t}$.
-Here we use the trivial bound $|\zeta(s)|\leq\zeta(\sigma)\leq\sigma/(\sigma-1)$, which is decreasing for all $\sigma>1$.
-Together with the trigonometric inequality from (21), we see that
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\zeta(1+\delta_{1,t})\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t.$
+Now we turn to the case $\sigma\geq 1+\delta_{1,t}$. Here we use the trivial bound
+$|\zeta(s)|\leq\zeta(\sigma)\leq\frac{\sigma}{\sigma-1},\qquad(\sigma>1),$ (48)
+where $\sigma/(\sigma-1)$ is decreasing in $\sigma$. Together with the trigonometric inequality from (23), we see that
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\zeta(1+\delta_{1,t})\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t.$ (49)
 We reiterate that although there are sharper bounds available to use here, this is sufficient for our purposes as the case where $\sigma\leq 1+\delta_{1,t}$ is dominant when computing constants.
@@
 $\left|\frac{1}{\zeta(1+it)}\right|\leq\frac{8}{\log(500)}\log t\leq 1.288\log t\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}).$
-Second, a computation from the proof of [2, Proposition A.2], implemented in interval arithmetic, states
+Second, a computation from the proof of [3, Proposition A.2], implemented in interval arithmetic, states
 $\left|\frac{1}{\zeta(1+it)}\right|\leq 2.079\log t\quad\text{for}\quad 2\leq t\leq 500.$
@@
 $\left|\frac{1}{\zeta(1+it)}\right|\leq 2.391\log t,$ (52)
-which is an improvement over what one gets from [4, Theorem 1.1].
+which is an improvement over what one gets from [5, Theorem 1.1].
 We make the remark that we include our computed values of $\sigma_{1}$ and $\eta$ in Tables 4, 5, and 6, for completeness sake.
@@
 Each entry is valid for $t\geq t_{0}=5$.
+5.4 An asymptotic improvement
+In this section we prove an asymptotic improvement over the bounds given earlier in Section 5, at the expense of a worse constant.
+For the sake of clarity, we first outline the general idea here, before proving the next two theorems.
+Assume we have the bound
+$\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
+for $t\geq t_{0}$ and $1-\delta\leq\sigma\leq 1+\delta_{1}$, where $\delta=d/\log t$ and $\delta_{1}=d_{1}/\log t$, as before. Then from (45), we have
+$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((\delta_{1}+\delta)\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\right)$
+$\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((d_{1}+d)C\right).$ (53)
+Notice that since the exponential factor is $O(1)$, the overall order is determined by the first factor.
+With the same argument involving the trigonometric inequality (22),
+$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq|\zeta(1+\delta_{1})|^{3/4}|\zeta(1+\delta_{1}+2it)|^{1/4},$ (54)
+and at the expense of a worse constant, we can estimate the second factor in (54) differently from the previous sections.
+We do this by using Lemma 6, which states that $|\zeta(s)|\leq C_{3}(\log t)^{2/3}$ for $\sigma\geq 1$.
+Along with Lemma 9, this leads to
+$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq(C_{4}\log t)^{3/4}(C_{3}(\log t)^{2/3})^{1/4},$
+which applied to (5.4) ultimately furnishes the bound
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq C_{5}(\log t)^{11/12}.$
+Theorem 6.
+Assume $RH(T)$.
+Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log t_{0}\leq 1$.
+Suppose also $R^{\prime}=1/2-d/\log t_{0}$.
+Then in the region
+$\sigma\geq 1-\frac{1}{W\log t}\qquad\text{and}\qquad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
+we have
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{RH(T)}(d,d_{1},t_{0})(\log t)^{11/12},$
+where
+$Y^{\prime}_{RH(T)}$ $(d,d_{1},t_{0})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)^{3/4},$
+$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log t_{0}}}{d_{1}}\right)^{3/4}\Bigg{\}},$
+with $A_{4}$ defined in Theorem 4, $B_{t_{0}}$ defined in Theorem 2, and $C_{3}$ defined in Lemma 6.
+Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
+Proof.
+The proof is identical to that of Theorem 4, except that for $\sigma\leq 1+\delta_{1,t}$, by Lemmas 6 and 9, we can replace (47) with
+$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|$ $\leq|\zeta(1+\delta_{1,t})|^{3/4}|\zeta(1+\delta_{1,t}+2it)|^{1/4}$
+$\leq\left(\frac{e^{\gamma(d_{1}/\log t_{0})}}{d_{1}}\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}.$
+Meanwhile, for $\sigma\geq 1+\delta_{1,t}$, the same argument with (22) and (48) leads to
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\left(\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}$
+in place of (49). ∎
+Theorem 7.
+Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log H\leq 1$.
+Suppose also $R^{\prime}=1/2-d/\log H$, and $W\geq W_{0}$ is chosen such that $d$ satisfies (36).
+Then in the region
+$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
+we have
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{H}(d,d_{1})(\log t)^{11/12},$
+where
+$Y^{\prime}_{H}$ $(d,d_{1})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log H}\right)^{3/4},$
+$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log H}}{d_{1}}\right)^{3/4}\Bigg{\}},$
+with $A_{5}$ defined in Theorem 5, and $C_{3}$ defined in Lemma 6.
+Proof.
+The proof is similar to that of Theorem 6, except we do not assume $RH(T)$. Then the process follows the proof of Theorem 5, making similar changes as with the earlier case $t\leq H$ in Theorem 6. ∎
+Combining Theorems 6 and 7, we prove Corollary 5, a result which holds for all $t\geq t_{0}\geq 5$.
+Proof of Corollary 5.
+The proof is analogous to that of Corollary 4, where the final result is obtained by optimising over the appropriate parameters.
+For the result on $\sigma=1$, the limiting region is again $2\exp(e^{2})\leq t\leq H$, where we have
+$\left|\frac{1}{\zeta(1+it)}\right|\leq 7.475(\log t)^{11/12},$
+with $d_{1}=0.72042$. ∎
+$t_{0}$ $Y^{\prime}(W_{0},t_{0})$ $d_{1}$ $W$ $Y^{\prime}(W,5)$ $d_{1}$
+$10$ $21.123$ $0.307587$ $W_{0}$ $47.054$ $0.170052$
+$10^{2}$ $11.643$ $0.542829$ $6$ $43.568$ $0.173052$
+$10^{3}$ $9.842$ $0.648796$ $7$ $38.179$ $0.178401$
+$10^{4}$ $8.885$ $0.725024$ $8$ $34.753$ $0.182368$
+$10^{5}$ $8.247$ $0.787048$ $9$ $32.391$ $0.185428$
+$10^{7}$ $7.425$ $0.885013$ $10$ $30.668$ $0.187861$
+$10^{9}$ $6.907$ $0.960526$ $11$ $29.356$ $0.18984$
+$10^{11}$ $6.545$ $1.021396$ $12$ $28.325$ $0.191481$
+$H$ $6.340$ $1.059541$ $13$ $27.494$ $0.192868$
+Table 7: Values for $Y^{\prime}(W,t_{0})$ in Corollary 5, with corresponding $W$ or $t_{0}$, and parameter $d_{1}$ for the case $t_{0}\leq t\leq H$ (rounded to six decimal places).
+Each entry is valid for $t\geq t_{0}$.
 6 Concluding Remarks
@@
 The downside to this is that we have not used better available bounds that require the full strength of $RH$.
-For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [24, §14], [21], [22], and [5] for more details).
+For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [25, §14], [23], [22], and [6] for more details).
+Even in the unconditional results used, there is room for improvement.
+As an illustration, Corollary 5, though asymptotically superior, beats Corollary 4 only when $t$ is large.
+In the case of $\sigma=1$, this happens for $t\geq\exp(161975)$.
+This is in part due to the relatively large constants in Corollary 5, and a direct way of lowering these constants is an improvement in the bound $|\zeta(1+it)|$, which correspondingly lowers $C_{3}$ in Lemma 5.
+Finally, we conclude with a discussion on the trigonometric inequality (22) used in many of our above proofs, in particular when bounding $|1/\zeta(s)|$ for $\sigma>1$.
+Under certain conditions, this method is superior to using the trivial bound.
+In general, when studying zero-free regions for $\zeta(s)$, increasing the degree of the non-negative trigonometric polynomial improves the results one gets.
+Surprisingly, the same cannot be said in our setting (with the exception of results in Section 5.4), and the classical polynomial we used is actually optimal.
+Take for instance Lemma 12, which is used in Theorems 4 and 5.
+Here the overall order of the bound is $O(\log t)$, and in this scenario the article [18] has an in-depth discussion on why the classical polynomial is best.
+On the other hand, in Theorems 6 and 7, the question is whether higher degree polynomials lead to a reduction in the exponent $11/12$, even if overall constants get worse.
+This is an interesting question to consider, since in this case the answer is less obvious, although preliminary calculations suggest that the classical polynomial is still the optimal one.
 Acknowledgements

### v3

@@
 These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s.
-Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
+Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement within the classical zero-free region: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
 1 Introduction
@@
 For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function.
-The reader is referred to [26], [4], [9], [25, (3.14)] for more information on the topic.
-Trudgian in [26] gives an explicit version of results in [25, §3] originally due to Landau, obtaining bounds of order $O(\log t)$.
+The reader is referred to [27], [4], [9], [26, (3.14)] for more information on the topic.
+Our main results will be presented in detail in Section 2, but here we highlight two of the more significant ones. In particular, Corollaries 4 and 5 of this article show respectively that for
+$\sigma\geq 1-\frac{1}{5.5587\log t}\qquad\text{and}\qquad t\geq 1000,$
+we have
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 3.632\log t,\qquad\text{and}\qquad\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 9.842(\log t)^{11/12}.$
+Now we give a brief history of results on $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$.
+Trudgian in [27] gives an explicit version of results in [26, §3] originally due to Landau, obtaining bounds of order $O(\log t)$.
 This was then extended in [14] and [7] to bounds of order $O(\log t/\log\log t)$.
@@
 This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
-The method in [25, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$.
+The method in [26, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$.
 It relies on two applications of the Borel–Carathédory theorem.
@@
 While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [27], we take the opportunity to point out an error in that paper.
-In the quantity $A$ of [26, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$.
+In the quantity $A$ of [27, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$.
 This leads to the overall constants being better than they should be.
@@
 In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [3, Proposition A.2].
+Finally, a novel argument utilising a combination of the classical non-negative trigonometric polynomial and a uniform bound of the form $\zeta(s)\ll(\log t)^{2/3}$ for $\sigma\geq 1$, allows us get an asymptotic improvement in the case of $1/\zeta(s)$.
+In particular, working within zero-free regions of the forms
+$\sigma\geq 1-\frac{1}{Z_{1}\log t},\qquad\sigma\geq 1-\frac{\log\log t}{Z_{2}\log t},\qquad\sigma\geq 1-\frac{1}{Z_{3}(\log t)^{2/3}(\log\log t)^{1/3}},$
+for some positive constants $Z_{1},\,Z_{2},\,Z_{3}$, one usually proves upper bounds for $1/\zeta(s)$ of the order
+$\log t,\qquad\frac{\log t}{\log\log t},\qquad(\log t)^{2/3}(\log\log t)^{1/3},$ (1)
+respectively. Without making any changes to the zero-free region used, our new method would allow one to refine the results in (1), to
+$(\log t)^{11/12},\qquad\frac{(\log t)^{11/12}}{(\log\log t)^{3/4}},\qquad(\log t)^{2/3}(\log\log t)^{1/4},$ (2)
+respectively.
+As far as we are aware, results of these types were not previously known.
+In this paper, we prove an explicit version of the first bound in (2).
+An explicit version of the third bound in (2) is given in [18].
 2 Main results
@@
 Currently, the best known result is due to [20], where above $H_{0}$, the value $W_{0}=5.558691$ is permissible. We take $W_{0}$ to be this value.
-We denote the Riemann hypothesis up to height $T$ by $RH(T)$, by which we mean: for all zeros $\rho=\beta+i\gamma$ such that $\zeta(\rho)=0$ and $\gamma\leq T$, we have $\beta=1/2$.
-Similarly we denote the Riemann hypothesis by $RH$, by which we mean all zeros $\rho$ have $\beta=1/2$.
+We denote the Riemann hypothesis up to height $T$ by $RH(T)$, by which we mean: for all non-trivial zeros $\rho=\beta+i\gamma$ such that $\zeta(\rho)=0$ and $\gamma\leq T$, we have $\beta=1/2$.
+Similarly we denote the Riemann hypothesis by $RH$, by which we mean all non-trivial zeros $\rho$ have $\beta=1/2$.
 2.2 Bounds on $|\zeta^{\prime}(s)/\zeta(s)|$
@@
 Corollary 1.
-Let $W\geq W_{0}$. In a region
+Let $W>W_{0}$. In a region
 $\sigma\geq 1-\frac{1}{W\log t}$
@@
 Corollary 4.
-Let $W\geq W_{0}$. In a region
+Let $W>W_{0}$. In a region
 $\sigma\geq 1-\frac{1}{W\log t}$
@@
 $\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y(W,t_{0})\log t,$
-where $Y(W_{0},10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
+where $Y(5.5587,10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
 In addition, we have for $\sigma=1$ and $t\geq 2$,
@@
 Corollary 5.
-Let $W\geq W_{0}$. In a region
+Let $W>W_{0}$. In a region
 $\sigma\geq 1-\frac{1}{W\log t}$
@@
 $\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(W,t_{0})(\log t)^{11/12},$
-where $Y^{\prime}(W_{0},10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
+where $Y^{\prime}(5.5587,10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
 In addition, we have for $\sigma=1$ and $t\geq 2$,
@@
 Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [13, Lemma 3.2].
-The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [24, §3.71]).
+The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [25, §3.71]).
 Lemma 1.
@@
 Thus the entire right-hand side of (3) is negative, and so we can discard any contribution from this.
-Combining this with the bound on $J$ and (3) proves the theorem. ∎
+Combining this with the bound on $J$ and (5) proves the lemma. ∎
 The following corollary is useful for proving results in a right-half plane, under $RH(T)$ or $RH$.
@@
 The proof is analogous to that of Lemma 2 with the additional assumption that the disc $|s_{0}-\rho|\leq R^{\prime}$ is completely free of zeros. ∎
-Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [27, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
+Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [28, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
 Lemma 3.
-[28, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
+[29, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
 $|f(s)|<C\exp(e^{k|t|})$
@@
 Proof.
-The proof is analogous to that of [27, Lemma 2.1], except that we are more precise in estimating (7) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
-Note that different to the proof of [27, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$.
+The proof is analogous to that of [28, Lemma 2.1], except that we are more precise in estimating (9) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
+Note that different to the proof of [28, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$.
 This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing (by the condition (10)) in $\sigma$.
@@
 Recall that Hiary, Patel, and Yang [15] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$.
-We now refer to Lemma 4, taking the maximum of (7) at $\sigma=1/2$. It has been shown in [29, p. 15] that $Q_{0}=1.31$ is permissible in (6) together with
+We now refer to Lemma 4, taking the maximum of (9) at $\sigma=1/2$. It has been shown in [30, p. 15] that $Q_{0}=1.31$ is permissible in (8) together with
 $(k_{1},k_{2},k_{3},Q_{0})=\left(0.618,\tfrac{1}{6},1,1.31\right),$
@@
 Proof.
-We use same argument in the first paragraph of [28, p.
+We use same argument in the first paragraph of [29, p.
 2], which modifies Ford’s proof in [10].
@@
 Lemma 7.
-[22, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
+[23, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
 $\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\frac{4}{\sigma-(1/2)}.$
@@
 where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$.
-Note that (12) holds for $\sigma>0$ by analytic continuation (see [25, Theorem 3.5]).
+Note that (14) holds for $\sigma>0$ by analytic continuation (see [26, Theorem 3.5]).
 In the region of $\sigma$ considered, we have
@@
 $\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\frac{t+\sigma_{1}}{6N^{2}}+\frac{t^{2}+\sigma_{1}^{2}}{12N^{2}}\right).$
-We thus have
-$|\zeta(s)|$ $\leq\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}+\frac{1}{64N^{4}}+\frac{1}{t}+\frac{t^{2}+\sigma_{1}^{2}+2(t+\sigma_{1})-1}{12N^{2}}\right).$
+We thus have that $|\zeta(s)|$ is upper bounded by
+$\exp\left(\frac{\log N}{W\log t}\right)\left(\log N+\gamma+\frac{1}{N}+\frac{1}{64N^{4}}+\frac{1}{t}+\frac{t^{2}+\sigma_{1}^{2}+2(t+\sigma_{1})-1}{12N^{2}}\right).$
 It remains to choose $N$. We can minimise the expression by taking $N=\lfloor\eta t\rfloor+1$ for some constant $\eta\in[2/t_{0},1-1/t_{0}]$. This will make
@@
 $|\zeta(s)|\leq 102\log t,$
-and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [27, Lem.
+and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [28, Lem.
 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [15]) and estimates on the $1$-line (see Backlund [1]) would yield
@@
 Lemma 9.
-[21, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
+[22, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
 $|\zeta(\sigma+it)|\leq\zeta(\sigma)\leq\frac{e^{\gamma(\sigma-1)}}{\sigma-1}.$
@@
 by the geometric series with $|1-\sigma|<2$. Taking the lower bound from this, one can then verify that $\phi_{0}(\sigma,k)\geq 0$ for $1<s\leq\sigma_{k}$ by calculating the first $k$ terms. ∎
-It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [21, Lemma 5.3].
+It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [22, Lemma 5.3].
 Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes.
@@
 Proof.
-Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [25, § 3.3] and [18]).
+Since $\sigma>1$, we use the classical non-negativity argument involving the trigonometric polynomial $2(1+\cos\theta)^{2}=3+4\cos\theta+\cos 2\theta$ (see [26, § 3.3] and [19]).
 That is, we use $\zeta^{3}(\sigma)|\zeta^{4}(\sigma+it)\zeta(\sigma+2it)|\geq 1$, which implies
@@
 Up to this point, the above with Theorem 1, where $M$ is as in (30), and $R=1/2$, will give a result that holds for
-$1+\epsilon-\frac{\alpha}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha}{2},$ (31)
+$1+\epsilon-\frac{\alpha}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha}{2},$
 and for a result on any larger $\sigma$, we simply use the trivial bound in (31).
@@
 Theorem 3.
-Let $d$ and $W\geq W_{0}$ be positive parameters, subject to (35), (36), and (44). Let also $10<\alpha<$ 1 be a parameter defined in (44), and depending on $d$ and $W$. In a region
+Let $d$ and $W>W_{0}$ be positive parameters, subject to (36), (37), and (45). Let also $10<\alpha<$ 1 be a parameter defined in (45), and depending on $d$ and $W$. In a region
 $\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H$
@@
 Table 3: Values for $Q$ in Corollary 1, with corresponding $W$ and $\alpha$ (rounded to six decimal places). Each entry is valid for $t\geq 5$.
-5 Bounds on $|1/\zeta(s)|$
-Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [25, Theorem 3.11]). It revolves around estimating the right-hand side of
+5 Estimates for $|1/\zeta(s)|$
+Moving from a bound on the logarithmic derivative of $\zeta$ to one on the reciprocal of $\zeta$ is done in the usual way (see [26, Theorem 3.11]). It revolves around estimating the right-hand side of
 $\log\left|\frac{1}{\zeta(\sigma+it)}\right|=-\textup{Re}\log\zeta\left(1+\delta_{1}+it\right)+\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x$ (46)
@@
 Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log H\leq\sigma_{1}\leq 2$.
-Suppose also $R^{\prime}=1/2-d/\log H$, and $W\geq W_{0}$ is chosen such that $d$ satisfies (36).
+Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37).
 Then in the region
@@
 $W$ $Y_{H}$ $d_{1}$ $\sigma_{1}$ $\eta$
-$W_{0}$ $2.307$ $1.059275$ $1.608256$ $3.775265$
+$5.5587$ $2.307$ $1.059544$ $1.412209$ $3.464102$
 $6$ $2.284$ $1.060503$ $1.608124$ $3.775135$
@@
 For $\sigma_{1}$ and $\eta$, any sufficiently large values will do, and further optimisations here will only improve the final constant from the fourth decimal place onwards.
-$t_{0}$ $Y(W_{0},t_{0})$ $d_{1}$ $\sigma_{1}$ $\eta$
-$10$ $8.101$ $0.307574$ $1.133578$ $3.071916$
-$10^{2}$ $4.339$ $0.542834$ $1.388735$ $3.417735$
-$10^{3}$ $3.632$ $0.648794$ $1.470345$ $3.459386$
-$10^{4}$ $3.264$ $0.725027$ $1.834773$ $3.463610$
-$10^{5}$ $3.021$ $0.787051$ $1.447903$ $3.464020$
-$10^{7}$ $2.711$ $0.885013$ $1.30488$ $3.464071$
-$10^{9}$ $2.518$ $0.960532$ $1.322282$ $3.464052$
-$10^{11}$ $2.383$ $1.021393$ $1.511713$ $3.464182$
-Table 5: Values for $Y(W_{0},t_{0})$ in Corollary 4, with corresponding $t_{0}$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$.
-Each entry is valid for $t\geq t_{0}$ and $W=W_{0}$.
+$t_{0}$ $Y(5.5587,t_{0})$ $d_{1}$ $\sigma_{1}$ $\eta$
+$10$ $8.101$ $0.307606$ $1.330692$ $3.068541$
+$10^{2}$ $4.339$ $0.542825$ $1.851563$ $3.417699$
+$10^{3}$ $3.632$ $0.648814$ $1.282661$ $3.465621$
+$10^{4}$ $3.264$ $0.725026$ $1.748467$ $3.463629$
+$10^{5}$ $3.021$ $0.787049$ $1.208003$ $3.464054$
+$10^{7}$ $2.711$ $0.885011$ $1.101137$ $3.464109$
+$10^{9}$ $2.518$ $0.960512$ $1.272668$ $3.445896$
+$10^{11}$ $2.383$ $1.021396$ $1.655004$ $3.464097$
+Table 5: Values for $Y(5.5587,t_{0})$ in Corollary 4, with corresponding $t_{0}$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$.
+Each entry is valid for $t\geq t_{0}$ and $W=5.5587$.
 $W$ $Y(W,5)$ $d_{1}$ $\sigma_{1}$ $\eta$
-$W_{0}$ $18.060$ $0.170011$ $1.105634$ $2.789282$
+$5.5587$ $18.060$ $0.170031$ $1.105647$ $2.784753$
 $6$ $16.722$ $0.173040$ $1.107519$ $2.781151$
@@
 Assume we have the bound
-$\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
+$\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
 for $t\geq t_{0}$ and $1-\delta\leq\sigma\leq 1+\delta_{1}$, where $\delta=d/\log t$ and $\delta_{1}=d_{1}/\log t$, as before. Then from (46), we have
-$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((\delta_{1}+\delta)\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\right)$
+$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((\delta_{1}+\delta)\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\right)$
 $\leq\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\exp\left((d_{1}+d)C\right).$ (54)
@@
 Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log H\leq 1$.
-Suppose also $R^{\prime}=1/2-d/\log H$, and $W\geq W_{0}$ is chosen such that $d$ satisfies (36).
+Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37).
 Then in the region
@@
 with $d_{1}=0.72042$. ∎
-$t_{0}$ $Y^{\prime}(W_{0},t_{0})$ $d_{1}$ $W$ $Y^{\prime}(W,5)$ $d_{1}$
-$10$ $21.123$ $0.307587$ $W_{0}$ $47.054$ $0.170052$
-$10^{2}$ $11.643$ $0.542829$ $6$ $43.568$ $0.173052$
-$10^{3}$ $9.842$ $0.648796$ $7$ $38.179$ $0.178401$
-$10^{4}$ $8.885$ $0.725024$ $8$ $34.753$ $0.182368$
+$t_{0}$ $Y^{\prime}(5.5587,t_{0})$ $d_{1}$ $W$ $Y^{\prime}(W,5)$ $d_{1}$
+$10$ $21.123$ $0.307588$ $5.5587$ $47.054$ $0.170052$
+$10^{2}$ $11.643$ $0.542830$ $6$ $43.568$ $0.173052$
+$10^{3}$ $9.842$ $0.648792$ $7$ $38.179$ $0.178401$
+$10^{4}$ $8.885$ $0.725028$ $8$ $34.753$ $0.182368$
 $10^{5}$ $8.247$ $0.787048$ $9$ $32.391$ $0.185428$
-$10^{7}$ $7.425$ $0.885013$ $10$ $30.668$ $0.187861$
-$10^{9}$ $6.907$ $0.960526$ $11$ $29.356$ $0.18984$
-$10^{11}$ $6.545$ $1.021396$ $12$ $28.325$ $0.191481$
-$H$ $6.340$ $1.059541$ $13$ $27.494$ $0.192868$
+$10^{7}$ $7.425$ $0.885011$ $10$ $30.668$ $0.187861$
+$10^{9}$ $6.907$ $0.960531$ $11$ $29.356$ $0.18984$
+$10^{11}$ $6.545$ $1.021397$ $12$ $28.325$ $0.191481$
+$H$ $6.340$ $1.059544$ $13$ $27.494$ $0.192568$
 Table 7: Values for $Y^{\prime}(W,t_{0})$ in Corollary 5, with corresponding $W$ or $t_{0}$, and parameter $d_{1}$ for the case $t_{0}\leq t\leq H$ (rounded to six decimal places).
@@
 To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results.
-As an example, for $W=W_{0}$, the condition (36) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$.
+As an example, for $W=5.5587$, the condition (37) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$.
 The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us.
@@
 The downside to this is that we have not used better available bounds that require the full strength of $RH$.
-For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [25, §14], [23], [22], and [6] for more details).
+For example, under $RH$ we could have used that $\zeta(1/2+it)=O(\exp(C\log t/\log\log t))$ for some $C>0$ to prove something stronger (see [26, §14], [24], [23], and [6] for more details).
 Even in the unconditional results used, there is room for improvement.

### v4

@@
-Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function
+Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta function
 Abstract
-In this article, we improve on two explicit bounds of order $\log t$ for $\sigma$ close to $1$: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$.
-Our method provides marked improvements in the constants, especially with regards to the latter.
-These savings come from replacing a lemma of Borel and Carathédory, with a Jensen-type one of Heath-Brown’s.
+In this article, we give explicit bounds of order $\log t$ for $\sigma$ close to $1$, for two quantities: $|\zeta^{\prime}(\sigma+it)/\zeta(\sigma+it)|$ and $|1/\zeta(\sigma+it)|$.
+We correct an error in the literature, and especially in the case of $|1/\zeta(\sigma+it)|$, also provide improvements in the constants.
 Using an argument involving the trigonometric polynomial, we additionally provide a slight asymptotic improvement within the classical zero-free region: $1/\zeta(\sigma+it)\ll(\log t)^{11/12}$.
+The same method applied to the Korobov–Vinogradov zero-free region gives a new record: the unconditional bound $1/\zeta(\sigma+it)\ll(\log t)^{2/3}(\log\log t)^{1/4}$.
 1 Introduction
@@
 For example, these come up in the Perron formula in the process of estimating the Chebyshev function $\psi(x)=\sum_{n\leq x}\Lambda(n)$, where $\Lambda(n)$ is the von Mangoldt function, or $M(x)=\sum_{n\leq x}\mu(n)$, where $\mu(n)$ is the Möbius function.
-The reader is referred to [27], [4], [9], [26, (3.14)] for more information on the topic.
-Our main results will be presented in detail in Section 2, but here we highlight two of the more significant ones. In particular, Corollaries 4 and 5 of this article show respectively that for
-$\sigma\geq 1-\frac{1}{5.5587\log t}\qquad\text{and}\qquad t\geq 1000,$
-we have
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 3.632\log t,\qquad\text{and}\qquad\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 9.842(\log t)^{11/12}.$
-Now we give a brief history of results on $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$.
-Trudgian in [27] gives an explicit version of results in [26, §3] originally due to Landau, obtaining bounds of order $O(\log t)$.
-This was then extended in [14] and [7] to bounds of order $O(\log t/\log\log t)$.
-However, such estimates tend to have large constants, often rendering them unsuitable for practical use.
+The reader is referred to [29], [4], [9], [26, (3.14)] for more information on this.
+The chief purpose of this article is to prove explicit upper bounds for $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$.
+Along the way, we also provide in §3 a collection of other improved bounds on quantities involving $\zeta(s)$, which might be useful for anyone seeking to do explicit computations in the area.
+Our main results will be presented in detail in §2, but let us first give a brief history of results on the two quantities $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$.
+Trudgian in [29] gives an explicit version of all the lemmas in [26, §3.9], which were originally due to Landau.
+This then allowed Trudgian to obtain explicit bounds on $\zeta^{\prime}(s)/\zeta(s)$ and $1/\zeta(s)$ of order $O(\log t)$.
+This was then extended in [15] and [7] to explicit bounds of order $O(\log t/\log\log t)$.
+However, estimates for the two quantities mentioned above tend to have large constants, often rendering them unsuitable for practical use.
 It is a non-trivial task to reduce these constants.
 This is due to the fact that the factor of $\zeta(s)$ in the denominator means that such estimates can only be valid in a zero-free region, and the constants rapidly increase in size as we approach the boundary of the region, i.e., $\zeta(s)\to 0$.
-The method in [26, §3] first estimates $|\zeta^{\prime}(s)/\zeta(s)|$, and uses the result to derive a bound for $|1/\zeta(s)|$.
-It relies on two applications of the Borel–Carathédory theorem.
-The first application serves to obtain a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$, which is then used in the second to arrive at the desired bound on $|\zeta^{\prime}(s)/\zeta(s)|$.
-As each application slightly worsens the final estimate, we were able to improve Trudgian’s results by cutting out the first application, and instead estimating ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ using a lemma of Heath-Brown [13, Lemma 3.2].
-While the improvements to bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem minor compared to [27], we take the opportunity to point out an error in that paper.
-In the quantity $A$ of [27, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$.
+While the bounds on $|\zeta^{\prime}(s)/\zeta(s)|$ given in this paper might seem worse compared to [29], we take the opportunity to point out an error in that paper.
+In the quantity $A$ of [29, Lemma 4], there should be a factor of $1/c$ instead of $1/2c$.
 This leads to the overall constants being better than they should be.
-For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,473)$ and $(12,54.3)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively.
-In this article, we have recovered the erroneous bounds and even slightly improved upon them.
-Furthermore, we directly derive a bound on $|1/\zeta(s)|$ from ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ rather than $|\zeta^{\prime}(s)/\zeta(s)|$, eliminating the need for the Borel–Carathédory theorem in this case, leading to a substantial improvement over [27].
-In addition, when $\sigma=1$, the second assertion of Corollary 4 is significantly better than [3, Proposition A.2].
-Finally, a novel argument utilising a combination of the classical non-negative trigonometric polynomial and a uniform bound of the form $\zeta(s)\ll(\log t)^{2/3}$ for $\sigma\geq 1$, allows us get an asymptotic improvement in the case of $1/\zeta(s)$.
+For instance, in Table $2$ of said paper, which shows bounds of the form $|\zeta^{\prime}(s)/\zeta(s)|\leq R_{1}\log t$ for $\sigma\geq 1-1/W\log t$, one should have $(W,R_{1})=(6,538.16)$ and $(12,59.49)$, instead of $(6,382.58)$ and $(12,40.14)$, respectively.
+This also impacts the subsequent bounds on $|1/\zeta(s)|$ ($R_{2}$ in [29]), which are derived from $R_{1}$.
+In this article, we have corrected the erroneous bounds and even slightly improved upon the (corrected) bounds.
+Furthermore, by adapting a method of [7, §5] and being more careful in our derivation on bounds of $|1/\zeta(s)|$, we obtain improvements in some (erroneously better than they should be) bounds in [29].
+In addition, when $\sigma=1$, the final assertion of Corollary 4 is significantly better than [3, Proposition A.2].
+Finally, a novel argument utilising a combination of the classical non-negative trigonometric polynomial and a uniform bound of the form $\zeta(s)\ll(\log t)^{2/3}$ for $\sigma\geq 1$, allows us to get an asymptotic improvement in the case of $1/\zeta(s)$.
 In particular, working within zero-free regions of the forms
@@
 In this paper, we prove an explicit version of the first bound in (2).
-An explicit version of the third bound in (2) is given in [18].
+The third bound in (2) is the best unconditional bound currently known, and an explicit version of it is given in [19].
 2 Main results
-Here we address matters of notation before listing some of the main results we prove.
+Here we address matters of notation before listing some of our main results. We note that the arguments used in this article are largely those of [7].
 2.1 Properties of the Riemann zeta function
-[…]
+In what follows, let $s=\sigma+it$, with $\sigma,\,t\in\mathbb{R}$, and define $H:=H_{0}-1/2$, where $H_{0}:=3.\,000\,175\,332\,800\cdot 10^{12}$ is the Riemann height, i.e., the height to which the Riemann hypothesis has been verified (see [22]).
+Our arguments will consider $t\leq H$ and $t\geq H$ as separate cases.
 The classical zero-free region for $\zeta(s)$ states that for $t\geq 2$,
 $\zeta(s)\neq 0,\qquad\text{ for }\qquad\sigma\geq 1-\frac{1}{W_{0}\log t}.$
-Currently, the best known result is due to [20], where above $H_{0}$, the value $W_{0}=5.558691$ is permissible. We take $W_{0}$ to be this value.
+Currently, the best known result is due to [21], where the value $W_{0}=5.558691$ is permissible. We take $W_{0}$ to be this value.
 We denote the Riemann hypothesis up to height $T$ by $RH(T)$, by which we mean: for all non-trivial zeros $\rho=\beta+i\gamma$ such that $\zeta(\rho)=0$ and $\gamma\leq T$, we have $\beta=1/2$.
@@
 2.2 Bounds on $|\zeta^{\prime}(s)/\zeta(s)|$
-Here we present our main results, which are corollaries of the theorems in Sections 4 and 5.
+Here we present our main results, which are corollaries of the theorems in §4 and §5.
 Corollary 1.
@@
 $\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q\log t,$
-where $(W,Q)=(10,49.359)$ is admissible, and other values of $Q$ with corresponding $W$ are presented in Table 3.
+where $(W,Q)=(10,71.220)$ is admissible, and other values of $Q$ with corresponding $W$ are presented in Table 3.
 Corollary 2.
@@
 Corollary 3.
-Assume $RH(H_{0})$. Then in the region
-$\sigma\geq\sigma_{0}>1/2\qquad\text{ and }\qquad 5\leq t\leq H,$
+In the region
+$\sigma\geq\sigma_{0}>1/2\qquad\text{ and }\qquad t_{0}\leq t\leq H,$
 we have
 $\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t,$
-where $(\sigma_{0},Q_{RH(H_{0})})=(0.8,33.249)$ is admissible, and other values for $Q_{RH(H_{0})}$ with corresponding values of $\sigma_{0}$ are given in Table 2.
-Furthermore, by assuming $RH$ instead of $RH(H_{0})$, the upper bound restriction on $t$ can be removed.
+where $(\sigma_{0},\,t_{0},\,Q_{RH(H_{0})})=(0.8,14,23.759)$ is admissible, and other values for $Q_{RH(H_{0})}$ with corresponding values of $\sigma_{0}$ and $t_{0}$ are given in Table 2.
+Furthermore, by assuming $RH$, the upper bound restriction on $t$ can be removed.
 2.3 Bounds on $|1/\zeta(s)|$
@@
 $\sigma\geq 1-\frac{1}{W\log t}$
-where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y(W,t_{0})\log t,$
-where $Y(5.5587,10^{3})=3.632$ is admissible, and other values of $Y(W_{0},t_{0})$ with corresponding $t_{0}$ are presented in Table 5, and values of $Y(W,5)$ with corresponding $W$ are presented in Table 6.
-In addition, we have for $\sigma=1$ and $t\geq 2$,
+where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 13$,
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y\log t,$
+where $(W,Y)=(13,1719)$ is admissible, and other values of $Y$ with corresponding $W$ are presented in Table 4.
+In addition, we have for $\sigma\geq 1$ and $t\geq 13$,
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 30.812\log t,$
+while for $\sigma=1$ and $t\geq 2$,
 $\left|\frac{1}{\zeta(1+it)}\right|\leq 29.388\log t.$
@@
 $\sigma\geq 1-\frac{1}{W\log t}$
-where $\zeta(\sigma+it)\neq 0$, we have for $t\geq t_{0}$,
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(W,t_{0})(\log t)^{11/12},$
-where $Y^{\prime}(5.5587,10^{3})=9.842$ is admissible, and other values of $Y^{\prime}(W,t_{0})$ with corresponding $W$ and $t_{0}$ are presented in Table 7.
-In addition, we have for $\sigma=1$ and $t\geq 2$,
+where $\zeta(\sigma+it)\neq 0$, we have for $t\geq 13$,
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}(\log t)^{11/12},$
+where $(W,Y)=(13,4904)$ is admissible, and other values of $Y^{\prime}$ with corresponding $W$ are presented in Table 5.
+In addition, we have for $\sigma\geq 1$ and $t\geq 13$,
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 87.725(\log t)^{11/12},$
+while for $\sigma=1$ and $t\geq 2$,
 $\left|\frac{1}{\zeta(1+it)}\right|\leq 81.900(\log t)^{11/12}.$
 3 Some preliminaries
-In this section, we present the results used in obtaining our estimates. For convenience, we begin by stating a few function theoretic results, which lie at the heart of our method.
-Theorem 1.
-[14, Theorem 4][Borel–Carathéodory] Let $s_{0}$ be a complex number, and let $R$ be a positive number possibly dependent on $s_{0}$.
+In this section, we present the results used in obtaining our estimates.
+The argument is largely that of [15] and [7], in which detailed proofs of the core lemmas can be found.
+At the heart of the method lies the Borel–Carathéodory Theorem (see [15, Theorem 4]), by which the following two lemmas are proven.
+We remark that Lemma 1 is used in the process of proving Lemma 2.
+Theorem 1 (Borel–Carathéodory).
+[15, Theorem 14] Let $s_{0}$ be a complex number.
+Let $R$ be a positive number, possibly depending on $s_{0}$.
 Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq R$.
-Let $M$ denote the maximum of $|\textup{Re}\,f(s)|$ on the boundary $|s-s_{0}|=R$.
+Let $M$ denote the maximum of $\textup{Re}\,f(s)$ on the boundary $|s-s_{0}|=R$.
 Then, for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
 $|f(s)|\leq\frac{2r}{R-r}M+\frac{R+r}{R-r}|f(s_{0})|.$
-Seeking to apply Theorem 1 with $f(s)=-\zeta(s)^{\prime}/\zeta(s)$, we obtain a bound on $M$ using the following lemma due to Heath-Brown [13, Lemma 3.2].
-The lemma is related to the well-known one of Jensen’s, and also to one of Carleman’s (see [25, §3.71]).
+If in addition $f(s_{0})=0$, then for any $r\in(0,R)$ and any $s$ such that $|s-s_{0}|\leq r$,
+$|f^{\prime}(s)|\leq\frac{2R}{(R-r)^{2}}M.$
 Lemma 1.
-[13, Lemma 3.2] Let $f(z)$ be holomorphic for $|z-a|\leq R$, and non-vanishing both at $z=a$ and on the circle $|z-a|=R$.
-Let $\rho_{j}=a+r_{j}\exp(i\theta_{j})$ be the zeros of $f(z)$ in the disc, where $\rho_{j}$ has multiplicity $n_{j}$.
-Then
-${\textup{Re}}\frac{f^{\prime}}{f}(a)=-\sum_{\rho_{j}}n_{j}\left(\frac{1}{r_{j}}-\frac{r_{j}}{R^{2}}\right)\cos\theta_{j}+\frac{1}{\pi R}\int_{0}^{2\pi}(\cos\theta)\log|f(a+Re^{i\theta})|\text{d}\theta.$
+[15, Lemma 15] Let $s_{0}$ be a complex number.
+Let $r$ and $\alpha$ be positive numbers (possibly depending on $s_{0}$) such that $\alpha<1/2$.
+Suppose that the function $f(s)$ is analytic in a region containing the disc $|s-s_{0}|\leq r$.
+Suppose further there is a number $A_{1}$ independent of $s$ such that,
+$\left|\frac{f(s)}{f(s_{0})}\right|\leq A_{1},\qquad|s-s_{0}|\leq r.$
+Then, for any $s$ in the disc $|s-s_{0}|\leq\alpha r$ we have
+$\left|\frac{f^{\prime}(s)}{f(s)}-\sum_{\rho}\frac{1}{s-\rho}\right|\leq\frac{4\log A_{1}}{r(1-2\alpha)^{2}},$
+where $\rho$ runs through the zeros of $f(s)$ in the disc $|s-s_{0}|\leq\tfrac{1}{2}r$, counted with multiplicity.
 Lemma 2.
-Let $d>0$ be some real constant and suppose that $\zeta(s)\neq 0$ in the region $\sigma\geq 1-1/W_{0}\log t$. Let $R^{\prime}>0$ be a parameter, chosen such that
-$\frac{d}{\log t}\leq\frac{1}{W_{0}\log(t+R^{\prime})}.$ (3)
-Then for $t\geq 3$, we have
-$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\leq\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+it+R^{\prime}e^{i\theta})||,$
-which holds uniformly for
-$\sigma\geq 1-\frac{d}{\log t}.$ (4)
-Proof.
-For $\sigma_{0}$ in the region stated in (4), we take $R=R^{\prime}$, $f(s)=\zeta(s)$, $a=s_{0}=\sigma_{0}+it_{0}$, and $\rho_{k}=s_{0}+r_{k}\exp(i\theta_{k})$ to be a zero of $\zeta(s)$ in Lemma 1 (where we will make the substitution $s_{0}\to s$ at the end of the proof).
-This gives us
-${\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(s_{0})=-\sum_{\rho_{k}}n_{k}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}+\frac{1}{\pi R^{\prime}}\int_{0}^{2\pi}(\cos\theta)\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta.$
-Note that if a zero of $\zeta(s)$ lies on the boundary $|s-s_{0}|=R^{\prime}$, then since the zeros are separable, we can always find an arbitrarily smaller $R^{\prime}$ where the boundary is free of zeros.
-For convenience, we define
-$J:=\int_{0}^{2\pi}(\cos\theta)\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta,$
-for which we seek a lower bound.
-For $20\leq\theta\leq\pi/$ 2 and $3\pi/2\leq\theta\leq 2\pi$, i.e. when $0\leq\cos\theta\leq 1$, we have
-$\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\geq-\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||,$
-so in the ranges $20\leq\theta\leq\pi/$ 2 and $3\pi/2\leq\theta\leq 2\pi$ combined,
-$J\geq-2\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||.$
-For the remaining range $\pi/2\leq\theta\leq 3\pi/2$, noting that $-1\leq\cos\theta\leq 0$, we have
-$\int_{\pi/2}^{3\pi/2}(\cos\theta)$ $\log|\zeta(s_{0}+R^{\prime}e^{i\theta})|\text{d}\theta\geq-2\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||,$
-and hence for the whole range $\theta\in[0,2\pi]$,
-$J\geq-4\max_{0\leq\theta\leq 2\pi}|\log|\zeta(s_{0}+R^{\prime}e^{i\theta})||.$
-Now Lemma 1 yields
-$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(s_{0})=\sum_{|s_{0}-\rho|\leq R^{\prime}}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}-\frac{1}{\pi R^{\prime}}J,$ (5)
-and since $r_{k}\cos\theta_{k}=-{\textup{Re}}(s_{0}-\rho_{k})$ and
-$\frac{\cos\theta_{k}}{r_{k}}=\frac{r_{k}\cos\theta_{k}}{r_{k}^{2}}=\frac{-{\textup{Re}}(s_{0}-\rho_{k})}{|s_{0}-\rho_{k}|^{2}}=-{\textup{Re}}\left(\frac{1}{s_{0}-\rho_{k}}\right),$
-the sum over all zeros in $|s_{0}-\rho|\leq R^{\prime}$ from (5) is
-$\sum_{|s_{0}-\rho|\leq R^{\prime}}\left(\frac{1}{r_{k}}-\frac{r_{k}}{R^{\prime 2}}\right)\cos\theta_{k}$ $=\sum_{|s_{0}-\rho|\leq R^{\prime}}-{\textup{Re}}\left(\frac{1}{s_{0}-\rho}-\frac{s_{0}-\rho}{R^{\prime 2}}\right)$
-$=-\sum_{|s_{0}-\rho|\leq R^{\prime}}(\sigma_{0}-\beta)\left(\frac{1}{|s_{0}-\rho|^{2}}-\frac{1}{R^{\prime 2}}\right),$ (6)
-where the second factor in (3) is non-negative.
-Notice that if
-$1-\frac{1}{W_{0}\log(t_{0}+R^{\prime})}\leq\sigma_{0},$
-then by our non-vanishing condition on $\zeta(s)$, we always have that $\sigma_{0}-\beta\geq 0$ in the disc $|s_{0}-\rho|\leq R^{\prime}$.
-This condition is fulfilled by our choice of $R^{\prime}$ in (3).
-Thus the entire right-hand side of (3) is negative, and so we can discard any contribution from this.
-Combining this with the bound on $J$ and (5) proves the lemma. ∎
+[15, Lemma 16] Let $s$ and $s_{0}$ be complex numbers with real parts $\sigma$ and $\sigma_{0}$, respectively.
+Let $r$, $\alpha$, $\beta$, $A_{1}$ and $A_{2}$ be positive numbers, possibly depending on $s_{0}$, such that $\alpha<1/2$ and $\beta<1$.
+Suppose that the function $f(s)$ satisfies the conditions of Lemma 1 with $r$, $\alpha$ and $A_{1}$, and that
+$\left|\frac{f^{\prime}(s_{0})}{f(s_{0})}\right|\leq A_{2}.$
+Suppose, in addition, that $f(s)\neq 0$ for any $s$ in the intersection of the disc $|s-s_{0}|\leq r$ and the right half-plane $\sigma\geq\sigma_{0}-\alpha r$.
+Then, for any $s$ in the disc $|s-s_{0}|\leq\alpha\beta r$,
+$\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{8\beta\log A_{1}}{r(1-\beta)(1-2\alpha)^{2}}+\frac{1+\beta}{1-\beta}A_{2}.$
 The following corollary is useful for proving results in a right-half plane, under $RH(T)$ or $RH$.
 Corollary 6.
-Let $\epsilon>0$ be a constant and suppose that $\zeta(s)\neq 0$ in the region $\sigma>1/2$ and $t\leq t+\epsilon$. Then for a fixed $R^{\prime}<\epsilon$ and $t\geq 3$, we have
-$-{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\leq\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+it+R^{\prime}e^{i\theta})||,$
-which holds uniformly for
-$\sigma\geq\frac{1}{2}+\epsilon.$ (7)
+If in the disc $|s-s_{0}|\leq r$, we have that $f(s)$ is analytic, non-zero, and satisfies $\left|\tfrac{f(s)}{f(s_{0})}\right|\leq A_{1}$, then for $0<\alpha_{0}<1$,
+$\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{2\log A_{1}}{r(1-\alpha_{0})^{2}},\qquad|s-s_{0}|\leq\alpha_{0}r.$
 Proof.
-The proof is analogous to that of Lemma 2 with the additional assumption that the disc $|s_{0}-\rho|\leq R^{\prime}$ is completely free of zeros. ∎
-Next we require some uniform bounds for $\zeta(s)$. Lemma 4 slightly improves [28, Corollary 1], which is proven using the Phragmén–Lindelöf Principle (Lemma 3).
+We follow a similar proof to that of Lemma 1 (also [15, Lemma 15]).
+Define $g(s)=\log(f(s)/f(s_{0}))$, where the logarithm branch is determined by $g(s_{0})=0$.
+Furthermore, since $f(s)\neq 0$ and is analytic in the disc $|s-s_{0}|\leq r$, hence $g(s)$ is analytic there.
+By hypothesis, we have $|f(s)/f(s_{0})|\leq A_{1}$ in $|s-s_{0}|\leq r$, so in the same region,
+$\textup{Re}(g(s))=\log\left|\frac{f(s)}{f(s_{0})}\right|\leq\log A_{1}.$
+To complete the proof, notice that $g^{\prime}(s)=f^{\prime}(s)/f(s)$, and apply Theorem 1 in the two discs $|s-s_{0}|\leq\alpha_{0}r<r$ to obtain
+$|g^{\prime}(s)|=\left|\frac{f^{\prime}(s)}{f(s)}\right|\leq\frac{2r\log A_{1}}{(r-\alpha_{0}r)^{2}}=\frac{2\log A_{1}}{r(1-\alpha_{0})^{2}},$
+for $|s-s_{0}|\leq\alpha_{0}r$. ∎
+Next we require some uniform bounds for $\zeta(s)$.
+Lemma 4 slightly improves [30, Corollary 1], which is proven using a generalisation of the Phragmén–Lindelöf Principle, due to Trudgian [27, Lemma 3].
+Unfortunately, Trudgian’s proof contains an error, the consequence of which is that his result is true only if one includes the restriction $\alpha_{2}=\beta_{2}$ (according to the notation of [27, Lemma 3]).
+In Lemma 3 below, we present a corrected version which includes this restriction.
+We refer the reader to [10] for an in-depth discussion on this error and a strategy to recovering the original result [^1] The author thanks Andrew Fiori for bringing this to our attention..
 Lemma 3.
-[29, Lemma 3] Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
+Let $a,b,Q$ be real numbers such that $b>a$ and $Q+a>1$. Let $f(s)$ be a holomorphic function on the strip $a\leq\textup{Re}(s)\leq b$ such that
 $|f(s)|<C\exp(e^{k|t|})$
-for some $C>0$ and $0<k<\tfrac{\pi}{b-a}$. Suppose further that there are $A,B,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\geq 0$ such that $\alpha_{1}\geq\beta_{1}$ and
-$|f(s)|\leq\begin{cases}A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|\quad\text{for}\textup{Re}(s)=a;\\ B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|\quad\text{for }\textup{Re}(s)=b.\end{cases}$
+for some $C>0$ and $0<k<\tfrac{\pi}{b-a}$. Suppose further that there are $A,B,\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\geq 0$ such that $\alpha_{1}\geq\beta_{1}$, $\alpha_{2}=\beta_{2}$, and
+$|f(s)|\leq\begin{cases}A|Q+s|^{\alpha_{1}}(\log|Q+s|)^{\alpha_{2}}\quad\text{for }\textup{Re}(s)=a;\\ B|Q+s|^{\beta_{1}}(\log|Q+s|)^{\beta_{2}}\quad\text{for }\textup{Re}(s)=b.\end{cases}$
 Then for $a\leq\textup{Re}(s)\leq b$, one has
-$|f(s)|\leq\Big{(}{A|Q+s|^{\alpha_{1}}\log^{\alpha_{2}}|Q+s|}\Big{)}^{\tfrac{b-\textup{Re}(s)}{b-a}}\Big{(}{B|Q+s|^{\beta_{1}}\log^{\beta_{2}}|Q+s|}\Big{)}^{\tfrac{\textup{Re}(s)-a}{b-a}}.$
+$|f(s)|\leq\Big{(}{A|Q+s|^{\alpha_{1}}|\log(Q+s)|^{\alpha_{2}}}\Big{)}^{\tfrac{b-\textup{Re}(s)}{b-a}}\Big{(}{B|Q+s|^{\beta_{1}}|\log(Q+s)|^{\beta_{2}}}\Big{)}^{\tfrac{\textup{Re}(s)-a}{b-a}}.$
 Lemma 4.
-Let there be a real number $h(s)=(s-1)\zeta(s)$, and let $\delta_{r}$ be a positive real number. Furthermore, let $Q_{0}\geq 0$ be a number for which
+Let $h(s)=(s-1)\zeta(s)$, and let $\delta_{r}$ be a positive real number. Furthermore, let $Q_{0}>1/2$ be a number for which
 $|h(\tfrac{1}{2}+it)|$ $\leq k_{1}|Q_{0}+\tfrac{1}{2}+it|^{k_{2}+1}(\log|Q_{0}+\tfrac{1}{2}+it|)^{k_{3}}$ (3)
-$|h(1+\delta_{r}+it)|$ $\leq\zeta(1+\delta_{r})|Q_{0}+1+\delta_{r}+it|,$
+$|h(1+\delta_{r}+it)|$ $\leq k_{4}|Q_{0}+1+\delta_{r}+it|(\log|Q_{0}+1+\delta_{r}+it|)^{k_{3}},$
 for all $t$. Then for $\sigma\in[\tfrac{1}{2},1+\delta_{r}]$ and $t\geq t_{0}\geq e$, we have
-$|\zeta$ $(s)|\leq\left(\zeta(1+\delta_{r})(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))\right)^{\frac{\sigma-0.5}{0.5+\delta_{r}}}$ (9)
-×$\bm{\times}\left(k_{1}(1+a_{0}(1+\delta_{r},Q_{0},t_{0}))^{k_{3}}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}+1}t^{k_{2}}\log^{k_{3}}t\right)^{\frac{1+\delta_{r}-\sigma}{0.5+\delta_{r}}},$
+$|\zeta(s)|\leq$ $\left(k_{4}\right)^{\frac{\sigma-0.5}{0.5+\delta_{r}}}(1+a_{0}(1+\delta_{r},Q_{0},t_{0}))^{k_{3}}\log^{k_{3}}t$ (4)
+$\bm{\times}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))\left(k_{1}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}}t^{k_{2}}\right)^{\frac{1+\delta_{r}-\sigma}{0.5+\delta_{r}}},$
 where
-$a_{0}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{2t^{2}\log t}+\frac{\pi}{2\log t}+\frac{\pi(\sigma+Q_{0})^{2}}{4t\log^{2}t},\qquad a_{1}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{t},$
-and the right-hand side of (9) is decreasing in $\sigma$, provided that $k_{i}\geq 0$, and the following hold:
-$t^{k_{2}}\log^{k_{3}}t\geq\frac{k_{1}(1+a_{0})^{k_{3}}(1+a_{1})^{k_{2}}}{\zeta(1+\delta_{r})}.$ (10)
+$a_{0}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{2t^{2}\log t}+\frac{\pi}{2\log t}+\frac{\pi(\sigma+Q_{0})^{2}}{4t\log^{2}t},\qquad a_{1}(\sigma,Q_{0},t)=\frac{\sigma+Q_{0}}{t}.$
+In addition, the right-hand side of (4) is decreasing in $\sigma$, provided that $k_{i}\geq 0$, and the following hold:
+$t\geq\left(\frac{k_{4}}{k_{1}(1+a_{1}(1+\delta_{r},Q_{0},t_{0}))^{k_{2}}}\right)^{1/k_{2}}.$ (5)
 Proof.
-The proof is analogous to that of [28, Lemma 2.1], except that we are more precise in estimating (9) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}$ and $1+a_{0}$.
-Note that different to the proof of [28, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$.
-This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing (by the condition (10)) in $\sigma$.
+The proof is analogous to that of [30, Lemma 2.1], through a direct application of Lemma 3 with $a=1/2$ and $b=1+\delta_{r}$, except that we are more precise in estimating (4) by retaining the dependence on $\sigma$ in the exponents of $1+a_{1}(\sigma,Q_{0},t)$ and $1+a_{0}(\sigma,Q_{0},t)$.
+Note that different to the proof of [30, Lemma 2.1], we use only two lines, $\sigma=1/2$ and $\sigma=1+\delta_{r}$, omitting the need for a bound at $\sigma=1$.
+This is because we intend to find the maximum bound over $\sigma\in[1/2,1+\delta_{r}]$, which we show occurs at $\sigma=1/2$, as our bound is decreasing in $\sigma$.
+The decreasing condition can be verified by checking when the derivative of the right-hand side of (4) is negative, and this is what gives rise to the condition in (5).
 Additionally, we found that further splitting the interval complicates (5) and raises the threshold of $t$ for the lemma to be valid.
@@
 provided that
-$t^{1/6}\log t\geq\frac{0.618(1+a_{0})(1+a_{1})^{1/6}}{\zeta(1+\delta_{r})}.$ (11)
+$t^{1/6}\log t\geq\frac{\zeta(1+\delta_{r})}{0.618(1+a_{1}(1+\delta_{r},1.31,t_{0}))^{1/6}},$ (6)
+where $\delta_{r}$, $a_{0}(\sigma,1.31,t)$, and $a_{1}(\sigma,1.31,t)$ are defined as in Lemma 4.
 Proof.
-Recall that Hiary, Patel, and Yang [15] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$.
-We now refer to Lemma 4, taking the maximum of (9) at $\sigma=1/2$. It has been shown in [30, p. 15] that $Q_{0}=1.31$ is permissible in (8) together with
-$(k_{1},k_{2},k_{3},Q_{0})=\left(0.618,\tfrac{1}{6},1,1.31\right),$
+Recall that Hiary, Patel, and Yang [16] have $|\zeta(1/2+it)|\leq 0.618t^{1/6}\log t$, for $t\geq 3$, while a quick calculation on Mathematica shows that $|\zeta(1/2+it)|\leq 1.461$ for $|t|<3$ (see also [14]).
+Meanwhile, one has by $|\zeta(s)|\leq\zeta(\sigma)$ that for all $t$,
+$|\delta_{r}+it||\zeta(1+\delta_{r}+it)|\leq\frac{\zeta(1+\delta_{r})}{\log t}|\delta_{r}+it|\log t.$
+Therefore, it can be verified that both bounds in (3) are met.
+We now refer to Lemma 4, taking the maximum of (4) at $\sigma=1/2$. It has been shown in [31, p. 15] that $Q_{0}=1.31$ is permissible in (3) together with
+$(k_{1},k_{2},k_{3},k_{4},Q_{0})=\left(0.618,\tfrac{1}{6},1,\tfrac{\zeta(1+\delta_{r})}{\log t},1.31\right),$
 hence we are done. ∎
@@
 Proof.
-We use same argument in the first paragraph of [29, p.
-2], which modifies Ford’s proof in [10].
-The only difference is that we now apply the argument to [2, §5], which itself is an improvement on Ford’s result.
+We use same argument in the first paragraph of [28, p.
+260], which modifies Ford’s proof in [11, Lemma 7.3].
+The only difference is that we now apply the argument with parameters $C$ and $D$ found in the proof of Theorem $1.1$ of [2, §5], which itself is an improvement on Ford’s result.
 This proves the lemma.
@@
 $|\zeta(\sigma+it)|\leq C_{3}(\log t)^{2/3},$
-with
-$C_{3}=58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right).$ (12)
+with $a_{0}$ defined as in Lemma 4, and
+$C_{3}=58.096\sqrt{1+9t_{0}^{-2}}\left(1+a_{0}(2,1,t_{0})\right)^{2/3}.$ (7)
 Proof.
-The proof is similar to that of [14, Lemma 7]. First, to show that the result holds for $1\leq\sigma\leq 2$, we use the Phragmén–Lindelöf Principle, i.e., Lemma 3, on the holomorphic function
+The proof is similar to that of [15, Lemma 7]. First, to show that the result holds for $1\leq\sigma\leq 2$, we use the Phragmén–Lindelöf Principle, i.e., Lemma 3, on the holomorphic function
 $f(s)=(s-1)\zeta(s).$
@@
 $|f(1+it)|=|t|\,|\zeta(1+it)|\leq 58.096\,|2+it|(\log|2+it|)^{2/3}.$
-This inequality is verified numerically for $|t|<3$, and is a consequence of Lemma 5 for $|t|\geq 3$.
-Also, on the line $s=2+it$,
+This inequality is verified numerically via Mathematica for $|t|<3$, and is a consequence of Lemma 5 for $|t|\geq 3$.
+Also, on the line $s=2+it$, the trivial bound $|\zeta(s)|\leq\zeta(\sigma)$ gives
 $|f(2+it)|\leq|1+it|\,\zeta(2)<58.096|3+it|(\log|3+it|)^{2/3},$
 for all real $t$. Now apply Lemma 3 with the following parameter values: $a=1$, $b=2$, $Q=1$, $\alpha_{1}=\beta_{1}=1$, $\alpha_{2}=\beta_{2}=2/3$, and $A=B=58.096$. We obtain
-$|f(s)|\leq 58.096|1+s|(\log|1+s|)^{2/3},$ (13)
+$|f(s)|\leq 58.096|1+s||\log(1+s)|^{2/3},$ (8)
 for $1\leq\sigma\leq 2$. Moreover, for the same range of $\sigma$, we have the following estimates:
-$\begin{split}\frac{|1+s|}{|s-1|}&\leq\frac{\sqrt{9+t^{2}}}{t}\leq\sqrt{1+9t_{0}^{-2}},\\ \log|1+s|&\leq\log\sqrt{9+t^{2}}\leq\log t+\log\sqrt{1+9t_{0}^{-2}}.\end{split}$
+$\begin{split}\frac{|1+s|}{|s-1|}&\leq\frac{\sqrt{9+t^{2}}}{t}\leq\sqrt{1+9t_{0}^{-2}},\\ |\log(1+s)|&\leq\left(1+a_{0}(2,1,t_{0})\right)\log t,\end{split}$
+where $a_{0}$ is defined as in Lemma 4. The second inequality above is derived using the same argument as that in [30, §2].
 Thus for $1\leq\sigma\leq 2$ and $t\geq t_{0}$, the inequality (8) implies that
-$|\zeta(\sigma+it)|$ $\leq 58.096\sqrt{1+9t_{0}^{-2}}\left(1+\frac{\log\sqrt{1+9t_{0}^{-2}}}{\log t_{0}}\right)\log^{2/3}t<C_{3}\log^{2/3}t,$
+$|\zeta(\sigma+it)|$ $\leq 58.096\sqrt{1+9t_{0}^{-2}}\left(1+a_{0}(2,1,t_{0})\right)^{2/3}\log^{2/3}t\leq C_{3}\log^{2/3}t,$
 as desired, where $C_{3}$ is defined in (7).
-Finally, we address the case where $\sigma\geq 2$. Here since $\sigma>1$, we apply the trivial bound in (49). Then for $\sigma\geq 2$, it is easy to verify that
+Finally, we address the case where $\sigma\geq 2$. Here since $\sigma>1$, we apply the trivial bound
+$|\zeta(s)|\leq\zeta(\sigma)\leq\frac{\sigma}{\sigma-1},$ (9)
+where $\sigma/(\sigma-1)$ is decreasing in $\sigma$. Then for $\sigma\geq 2$, it is easy to verify that
 $|\zeta(\sigma+it)|\leq 2\leq C_{3}(\log t)^{2/3}$
@@
 Lemma 7.
-[23, Lemma 4] In the region 1/$21/2<\sigma\leq 3/$ 2 and $0\leq t\leq 2\exp(e^{2})$, we have
+[24, Lemma 4] In the region $1/2<\sigma\leq 3/2$ and $0\leq t\leq 2\exp(e^{2})$, we have
 $\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\frac{4}{\sigma-(1/2)}.$
-In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones.
-We would like to point out that Lemma 8 and Corollary 9, which can both be viewed as an extension of Corollary 8 inside the critical strip, are not actually required for the purposes of this paper.
+In the remainder of this section, we present some estimates concerning the Riemann zeta function, including some for $\sigma>1$, which are better than the trivial ones in some finite range.
+We would like to point out that Lemmas 10 and 11, as well as Corollary 10, do not contribute to the main results of this paper, but are included as they are potentially useful estimates for anyone seeking to do explicit computations involving $\zeta(s)$.
+Similarly, Lemma 8 and Corollary 9, which can both be viewed as an extension of Corollary 8 inside the critical strip, are not actually required for the main purposes of this paper.
 However, we include them here as the proofs are similar to that of Corollary 8, and at the same time we provide a brief discussion on their usefulness.
-Lemma 8 is an improvement on [16, Proposition 5.3.2], which proves a similar result.
+Lemma 8 is an improvement on [17, Proposition 5.3.2], which proves a simlar result.
 Lemma 8.
@@
 $\zeta(s)=\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}+\frac{N^{1-s}}{s-1}+\frac{s}{12N^{s+1}}-\frac{s(s+1)}{2}\int_{N}^{\infty}\frac{\varphi^{*}(u)}{u^{s+2}}\text{d}u,$ (10)
-where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$.
+where $\varphi^{*}(u)$ is the periodic function obtained by extending the polynomial $\varphi(u)=u^{2}-u+1/6$ on the interval $[0,1]$ (for full details see e.g., [1, Equation (8)]).
 Note that (10) holds for $\sigma>0$ by analytic continuation (see [26, Theorem 3.5]).
@@
 $|N^{-s}|=N^{-\sigma}\leq\exp\left(-\Big{(}1-\frac{1}{W\log t}\Big{)}\log N\right)=N^{-1}\exp\left(\frac{\log N}{W\log t}\right).$ (11)
-This bound can be used throughout (14). The first term in (14) can also be bounded using the estimate in [11, Lemma 2.1], so for $N\geq 2$,
+This bound can be used throughout (10). The first term in (10) can also be bounded using the estimate in [12, Lemma 2.1], so for $N\geq 2$,
 $\left|\sum_{n=1}^{N-1}\frac{1}{n^{s}}+\frac{1}{2N^{s}}\right|$ $=\left|\sum_{n=1}^{N}\frac{1}{n^{s}}-\frac{1}{2N^{s}}\right|$
@@
 Let $s=\sigma+ikt$ with $1\leq\sigma\leq\sigma_{1}\leq 2$, $t\geq t_{0}\geq 3$, and a fixed $k\geq 1$. For $0<\eta/k<t_{0}$, we have
-$|\zeta(\sigma+ikt)|<\log kt+C(\sigma_{1},t_{0},k,\eta),$
+$|\zeta(\sigma+ikt)|<\log(kt)+C(\sigma_{1},t_{0},k,\eta),$
 where
@@
 Proof.
-The proof follows the argument of [3, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$. We omit the proof since it is almost identical to that of Lemma 8. ∎
+The proof follows the argument of [3, Proposition A.1], originally due to Backlund [1], and extends their result to $t<500$.
+Note that here we have instead the choice of $N=\lfloor kt/\eta\rfloor+1$.
+A slightly better estimate of the harmonic sum from [12, Lemma 2.1] was also used.
+We omit the proof since the process is almost identical to that of Lemma 8.
+∎
 To use Lemma 8, we want to choose a $t_{0}$ and optimise over $\eta$.
@@
    $t_{0}$    $B_{1}$    $\eta$
-      $3$       $2.1173$       $2/3$
+   33    $2.1173$    $2/3$
    $10^{3}$    $0.2254$    $0.41$
@@
 Ford’s bound, while asymptotically better, is only sharper than bounds of the form $\zeta(s)\ll\log t$ for very large $t$.
-For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 9 is an improvement to [10, Theorem 1].
+For $\sigma\in[1-1/(W_{0}\log t),1]$ and $3\leq t\leq 10^{10^{5}}$, Corollary 9 is an improvement to [11, Theorem 1].
 This makes it useful for results over smaller $t$.
@@
 $|\zeta(s)|\leq 3.13\log t,$
-whereas degrading a Richert-type bound like Ford’s [10, Theorem 1] would give
+whereas degrading a Richert-type bound like Ford’s [11, Theorem 1] would give
 $|\zeta(s)|\leq 102\log t,$
-and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [28, Lem.
+and a bound obtained via the Phragmén–Lindelöf principle like (2.4) in [30, Lem.
 2.1] using estimates on the $1/2$-line (see Hiary, Patel, and Yang [16]) and estimates on the $1$-line (see Backlund [1]) would yield
@@
 Next we have some bounds on $\zeta(\sigma)$ for real $\sigma>1$.
-These are useful in their own right, but we show they can be used to obtain good bounds on $-\zeta^{\prime}(\sigma)/\zeta(\sigma)$ in a similar region.
+These are useful in their own right, but we show they can be used to obtain good bounds on $\zeta^{\prime}(\sigma)/\zeta(\sigma)$ in a similar region.
 Lemma 9.
-[22, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
+[23, Lemma 5.4] Let $\gamma$ denote Euler’s constant. For $\sigma>1$ we have
 $|\zeta(\sigma+it)|\leq\zeta(\sigma)\leq\frac{e^{\gamma(\sigma-1)}}{\sigma-1}.$
 Lemma 10.
-Let $\gamma_{n}$ be the Laurent–Stieltjes constants (with $\gamma:=\gamma_{0}$). For integers $k$ such that $1\leq k\leq 1500$, and real $\sigma$ satisfying $31<\sigma\leq\sigma_{k}<$ 3, we have
+Let $\gamma_{n}$ be the Laurent–Stieltjes constants (with $\gamma:=\gamma_{0}$). For integers $k$ such that $1\leq k\leq 1500$, and real $\sigma$ satisfying $1<\sigma\leq\sigma_{k}<3$, we have
 $\zeta(\sigma)\geq\frac{1}{\sigma-1}+\gamma+\phi_{0}(\sigma,k),$
@@
 Proof.
-Recall from the Laurent series expansion of the zeta function
-$\zeta(\sigma)=\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}+\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}.$ (20)
-To prove the non-negativity of (19) we require the last two terms of (20) together to be positive. We have the bounds on $\gamma_{n}$ due to Lavrik [17]:
+To prove the first assertion of the lemma, we have from the Laurent series expansion (about $\sigma=1$) of the zeta function
+$\zeta(\sigma)$ $=\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}+\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}$ (16)
+$\geq\frac{1}{\sigma-1}+\gamma+\sum_{n=1}^{k}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}-\left|\sum_{n=k+1}^{\infty}\frac{(-1)^{n}(\sigma-1)^{n}\gamma_{n}}{n!}\right|.$
+We have the bounds on $\gamma_{n}$ due to Lavrik [18]:
 $|\gamma_{n}|\leq\frac{n!}{2^{n+1}},\quad\quad n=1,2,3,\ldots,$ (17)
@@
 by the geometric series with $|1-\sigma|<2$. Taking the lower bound from this, one can then verify that $\phi_{0}(\sigma,k)\geq 0$ for $1<s\leq\sigma_{k}$ by calculating the first $k$ terms. ∎
-It is likely that $\phi(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [22, Lemma 5.3].
+It is likely that $\phi_{0}(\sigma,k)$ is always positive for any $\sigma>1$, although proving that would require a different approach, perhaps one akin to [23, Lemma 5.3].
 Our method in Lemmas 10 and 11, while simpler, is restricted by bounds on $|\gamma_{n}|$ which loses some information about sign changes.
@@
 Let the hypotheses of Lemma 10 be fulfilled. Then for $\sigma$ real, we have
-$-\zeta(\sigma)\leq\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}},\qquad 1<\sigma<3,$
+$\zeta^{\prime}(\sigma)\leq\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}},\qquad 1<\sigma<3,$
 and
@@
 $=\frac{1}{(\sigma-1)^{2}}+\frac{1}{2(\sigma-1)}\sum_{n=1}^{\infty}n\left(\frac{\sigma-1}{2}\right)^{n}=\frac{1}{(\sigma-1)^{2}}+\frac{1}{(\sigma-3)^{2}}$
-for $|1-\sigma|<2$ by identification with the polylogarithm $\textup{Li}_{-1}(z)=z/(1-z)^{2}$. Dividing by the bound in the Lemma 10 proves the result. ∎
+for $|1-\sigma|<2$ by identification with the polylogarithm $\textup{Li}_{-1}(z)=z/(1-z)^{2}$. Dividing by the bound in Lemma 10 proves the result. ∎
 These bounds are good for $\sigma$ close to $1$. For instance, we have that
@@
 $\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\frac{-\zeta^{\prime}(\sigma)}{\zeta(\sigma)}\leq\frac{\phi_{2}(\sigma_{0})}{\sigma-1},$
-[…]
+where $\phi_{2}(\sigma_{0})=\max\left\{\phi_{1}(\sigma_{0},10),\,0.852\right\}$.
 Proof.
@@
 $\left|\frac{1}{\zeta(1+\kappa+it)}\right|\leq|\zeta(1+\kappa)|^{3/4}|\zeta(1+\kappa+2it)|^{1/4}.$ (21)
-The two factors on the right-hand side can be bounded with Lemmas 9 and 8, respectively, with the aid of the bound
+The two factors on the right-hand side can be bounded with Lemma 9 and Corollary 8, respectively, with the aid of the bound
 $\log 2t$ $\leq\left(1+\frac{\log 2}{\log t_{0}}\right)\log t.$
@@
 4.1 Assuming a partial Riemann hypothesis
-In this section, we assume $RH(T)$, and estimate $|\zeta^{\prime}(s)/\zeta(s)|$ in a right-half plane: $\sigma\geq 1/2+\epsilon$, $t_{0}\leq t\leq T$.
-We will first prove the result for $T_{0}$ and then make the substitution to $t$ at the end.
-The idea is to construct discs $|s-s_{0}|\leq 1/2$ which have origins slightly to the right of the $1$-line, centred at $s_{0}=\sigma_{0}+iT_{0}$, where
-$\sigma_{0}=1+\epsilon,\qquad e+1/2\leq t_{0}\leq T_{0}\leq T-1/2,$
-and $\epsilon\leq 1/2$ is a parameter to be determined later.
-We then extend their radii into the critical strip, so that Theorem 1 applied to $f(s):=-\zeta^{\prime}(s)/\zeta(s)$ gives us the desired estimates.
-Due to $RH(T)$, $\zeta(s)$ is analytic and non-zero in the discs.
-Choosing a fixed $R^{\prime}<\epsilon$, we have by Corollary 6 that for $1/2+\epsilon\leq\sigma\leq 3/2+\epsilon$,
-$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT_{0})\right|\leq M:=\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+iT_{0}+R^{\prime}e^{i\theta})||,$ (26)
-which we estimate, by Lemma 7 with $\delta_{r}=1/2+\epsilon$ and $T_{0}\geq t_{0}\geq e+R^{\prime}$. In the region $1/2\leq\sigma\leq 3/2+2\epsilon$, we have
-$|\zeta(s)|\leq A_{0}(T_{0}+R^{\prime})^{1/6}\log(T_{0}+R^{\prime})\leq A_{1}T_{0}^{1/6}\log T_{0},$ (27)
-where
-$A_{1}$ $=A_{0}\left(1+\frac{R^{\prime}}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{t_{0}}\right)}{\log t_{0}}\right),$ (28)
-with
-$A_{0}$ $=0.618\left(1+a_{1}\left(\frac{3}{2}+2\epsilon,1.31,t_{0}-R^{\prime}\right)\right)^{7/6}$ (29)
-×$\qquad\qquad\bm{\times}\left(1+a_{0}\left(\frac{3}{2}+2\epsilon,1.31,t_{0}-R^{\prime}\right)\right).$
-Combining (27), (29), and (28) with (26) thus gives for $1/2+\epsilon\leq\sigma\leq 3/2+\epsilon$,
-$M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{1}\right).$ (30)
-Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we can use the trivial bound for $\sigma>1$ [12, Lemma 70.1], which gives us
-$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{1}{\sigma_{0}-1}.$ (31)
-However, when bounding $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$ in an interval away from the $1$-line, that is to say, where $\sigma_{0}\leq\sigma\leq 1.5$ with $\sigma_{0}$ not tending to $1$, we can do better than the trivial bound by using Corollary 10.
-We have for $\sigma_{0}\leq\sigma\leq 1.5$,
-$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq-\frac{\zeta^{\prime}(\sigma_{0})}{\zeta(\sigma_{0})}<\frac{\phi_{2}(\sigma_{0})}{\sigma_{0}-1}.$ (32)
-Up to this point, the above with Theorem 1, where $M$ is as in (30), and $R=1/2$, will give a result that holds for
-$1+\epsilon-\frac{\alpha}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha}{2},$
-and for a result on any larger $\sigma$, we simply use the trivial bound in (31).
-Finally, putting everything together and making the substitution $T_{0}\mapsto t$, we prove our result.
+In this section, we assume $RH(T)$, and estimate $|\zeta^{\prime}(s)/\zeta(s)|$ in a right-half plane: $\sigma>1/2+\epsilon$, $t_{0}\leq t\leq T$.
 Theorem 2.
-Assume $RH(T)$. Let $\alpha$, $\epsilon$, and $R^{\prime}$ be parameters, such that $20<R^{\prime}<\epsilon\leq 1/$ 2 and $10<\alpha<$ 1. In the region
-$\sigma\geq 1+\epsilon-\frac{\alpha}{2}\quad\text{and}\quad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
+Assume $RH(T)$. Let $\alpha_{0}$ and $\epsilon$ be parameters, such that $0<\epsilon\leq 1/2$ and $0<\alpha_{0}<1$. In the region
+$\sigma\geq 1+\epsilon-\frac{\alpha_{0}}{2}\quad\text{and}\quad 3\leq t_{0}\leq t\leq T-\frac{1}{2},$
 we have
@@
 with
-$Q_{RH(T)}$ $=\max\Bigg{\{}\frac{2\alpha}{(1-\alpha)\pi R^{\prime}}\Bigg{(}\frac{1}{6}+B_{t_{0}}+\frac{\log A_{1}}{\log t_{0}}\Bigg{)}$ (33)
-$\qquad\qquad+\left(\frac{1+\alpha}{1-\alpha}\right)\frac{\phi_{2}(1+\epsilon)}{\epsilon\log t_{0}},\,\frac{1}{(\epsilon+(\alpha/2))\log t_{0}}\Bigg{\}},$
-where $A_{1}$ is defined in (28), $\phi_{2}(\sigma)$ is defined in Corollary 10, and
-$B_{t_{0}}=\begin{cases}\frac{1}{e}&\text{ if }t_{0}\leq e^{e},\\ \frac{\log\log t_{0}}{\log t_{0}}&\text{ if }t_{0}>e^{e}.\end{cases}$
+$Q_{RH(T)}$ $=\max\Bigg{\{}\frac{4}{(1-\alpha_{0})^{2}}\Bigg{(}\frac{1}{6}+2B_{t_{0}}+\frac{\log A_{3}}{\log t_{0}}\Bigg{)}\,,\,\frac{1}{(\epsilon+(\alpha_{0}/2))\log t_{0}}\Bigg{\}},$ (22)
+where $A_{3}$ is defined in (4.1), and
+$B_{t_{0}}=\begin{cases}\frac{1}{e}&\text{ if }t_{0}\leq e^{e},\\ \frac{\log\log t_{0}}{\log t_{0}}&\text{ if }t_{0}>e^{e},\end{cases}$ (23)
+provided that
+$t^{1/6}\log t\geq\frac{\zeta(\frac{3}{2}+\epsilon)}{0.618(1+a_{1}(\frac{3}{2}+\epsilon,1.31,t_{0}))^{1/6}}$ (24)
+is satisfied, with $a_{1}(\sigma,1.31,t)$ defined as in Lemma 4.
 Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
+Proof.
+We will first prove the result for $T_{0}$ and then make the substitution to $t$ at the end.
+The idea is to construct discs $|s-s_{0}|\leq 1/2$ which have origins slightly to the right of the $1$-line, centred at $s_{0}=\sigma_{0}+iT_{0}$, where
+$\sigma_{0}=1+\epsilon,\qquad 3\leq t_{0}\leq T_{0}\leq T-1/2,$
+and $\epsilon\leq 1/2$ is a parameter to be determined later. We then extend their radii into the critical strip, so that Corollary 6 applied with
+$f(s):=\zeta(s),\qquad s_{0}=\sigma_{0}+iT_{0},\qquad r=1/2,$
+gives us the desired estimates. Due to $RH(T)$, $\zeta(s)$ is analytic and non-zero in the discs.
+In order to do this, we first need to determine $A_{1}$.
+Let $V(\epsilon,\sigma_{1},t_{0},\eta)$ be defined as in Lemma 12, with $\sigma_{1}$ and $\eta$ as parameters.
+Then by Corollary 7 with $\delta_{r}=r+\epsilon$, and Lemma 12 with $\kappa=\epsilon$, we have for $1/2\leq\sigma\leq\sigma_{0}+r$ and each $T_{0}\in[T_{0}-r,T_{0}+r]$,
+$\left|\frac{\zeta(\sigma+iT_{0})}{\zeta(\sigma_{0}+iT_{0})}\right|\leq A_{1}:=A_{3}T_{0}^{1/6}(\log T_{0})^{2}$ (25)
+where
+$A_{3}:=0.618$ $(1+a_{0}(\sigma_{0}+r,1.31,t_{0}))(1+a_{1}(\sigma_{0}+r,1.31,t_{0}))^{7/6}$
+$\bm{\cdot}\left(1+\frac{r}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{r}{t_{0}}\right)}{\log t_{0}}\right)V(\epsilon,\sigma_{1},t_{0},\eta);$ (26)
+provided (6) is satisfied.
+Now choosing a fixed positive $\alpha_{0}<1$, we have by Corollary 6,
+$\left|\frac{\zeta^{\prime}(\sigma+iT_{0})}{\zeta(\sigma+iT_{0})}\right|$ $\leq\frac{4}{(1-\alpha_{0})^{2}}\left(\frac{1}{6}\log T_{0}+2\log\log T_{0}+\log A_{3}\right)$
+$\leq\frac{4}{(1-\alpha_{0})^{2}}\left(\frac{1}{6}+2B_{t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)\log T_{0},$ (27)
+which holds for
+$1+\epsilon-\frac{\alpha_{0}}{2}\leq\sigma\leq 1+\epsilon+\frac{\alpha_{0}}{2}.$
+Note that $B_{t_{0}}$ is defined in (23), and is utilised since $\log\log t/\log t$ attains its maximum at $t=e^{e}$, after which it is decreasing.
+To obtain a bound that is valid also for $\sigma\geq 1+\epsilon+\alpha_{0}/2$, we use the trivial bound for all $\sigma>1$ [13, Lemma 70.1]:
+$\left|\frac{\zeta^{\prime}(s)}{\zeta(s)}\right|\leq-\frac{\zeta^{\prime}(\sigma)}{\zeta(\sigma)}<\frac{1}{\sigma-1}$ (28)
+with $\sigma=1+\epsilon+\alpha_{0}/2$, and then take the maximum of (4.1) and (28).
+Finally, making the substitution from $T_{0}$ to $t$ proves the result. ∎
 Assuming $RH(H_{0})$ allows us to prove Corollary 3, which holds for $t\leq H$.
 Proof of Corollary 3.
-For $t\geq t_{0}$, in Theorem 2 set $\sigma_{0}=1+\epsilon-\frac{\alpha}{2}$.
-Then fix $t_{0}$ and optimise (33) over $\epsilon,\,R^{\prime},\,\alpha$, while ensuring that the restrictions set in Theorem 2 and (11) hold.
-In practice $R^{\prime}$ is taken to be arbitrarily smaller than $\epsilon$, differing only after the sixth decimal place.
-The final value of $t_{0}=5$ is taken to be the one where (11) holds for all $\sigma_{0}$ in Table 2.
-This gives the result.
+For $t\geq t_{0}$ and $\sigma\geq\sigma_{0}$, in Theorem 2 set $\sigma_{0}=1+\epsilon-\frac{\alpha_{0}}{2}$.
+Then fixing $t_{0}$ and $\sigma_{0}$, optimise (22) over $\epsilon$ (which also determines $\alpha_{0}$), and also over $\sigma_{1},\,\eta$ (due to the quantity $V$ in $A_{3}$), while ensuring that the restrictions set in Theorem 2 (and by extension Lemma 12 and Corollary 8), hold.
+The value $t_{0}$ is chosen by finding the smallest value of $t$ that satisfies (24).
 ∎
-$\sigma_{0}$ $Q_{RH(H_{0})}$ $\epsilon$ $\alpha$
-$0.51$ $38889$ $0.004948$ $0.989897$
-$0.55$ $1508$ $0.024334$ $0.948668$
-$0.6$ $361.726$ $0.047605$ $0.89521$
-$0.65$ $153.957$ $0.070599$ $0.841198$
-$0.7$ $82.739$ $0.092279$ $0.784558$
-$0.75$ $50.441$ $0.113238$ $0.726476$
-$0.8$ $33.249$ $0.129407$ $0.658814$
-$0.85$ $23.086$ $0.144234$ $0.588468$
-$0.9$ $16.601$ $0.160246$ $0.520492$
-$0.95$ $12.235$ $0.174136$ $0.448272$
-$1$ $9.161$ $0.176533$ $0.353067$
-Table 2: Values for $Q_{RH(H_{0})}$ and $\sigma\geq\sigma_{0}$ in Corollary 3, with corresponding $\epsilon,\,\alpha$ (all rounded to six decimal places). Each entry is valid for $5\leq t\leq H$.
+$\sigma_{0}$ $t_{0}$ $Q_{RH(H_{0})}$ $\epsilon$ $\sigma_{1}$ $\eta$
+$0.51$ $15$ $31447$ $0.000462$ $1.965754$ $2.608856$
+$0.55$ $14$ $1086$ $0.002766$ $1.901324$ $3.329589$
+$0.6$ $14$ $249.260$ $0.006041$ $1.795116$ $3.162609$
+$0.65$ $14$ $104.992$ $0.009573$ $1.846979$ $3.195987$
+$0.7$ $14$ $56.739$ $0.013287$ $1.797626$ $3.198866$
+$0.75$ $14$ $35.162$ $0.017144$ $1.155376$ $2.868261$
+$0.8$ $14$ $23.759$ $0.021126$ $1.392644$ $3.173843$
+$0.85$ $14$ $17.050$ $0.025202$ $1.763188$ $3.402204$
+$0.9$ $13$ $13.123$ $0.029430$ $1.227811$ $3.283187$
+$0.95$ $13$ $10.175$ $0.033682$ $1.227876$ $3.283802$
+11 $13$ $8.101$ $0.037999$ $1.889284$ $3.054339$
+Table 2: Values for $Q_{RH(H_{0})}$ and $\sigma\geq\sigma_{0}$ in Corollary 3, with corresponding $\epsilon,\,\sigma_{1},\,\eta$ (all rounded to six decimal places).
+Each entry is valid for $t_{0}\leq t\leq H$.
+We note that [5, Theorem 1.1] implies that
+$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 0.639\log t$
+for $10^{6}\leq t\leq 2.99997\cdot 10^{12}$. This result allows us to show that
+$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 40.440\log\log t$ (29)
+holds unconditionally for $e^{e}\leq t\leq 2.99997\cdot 10^{12}$.
+To prove this, recall from §2.1 that $RH$ has been verified up to $3.000175332800\cdot 10^{12}$.
+Now note that $\log t/\log\log t$ is increasing for $t\geq e^{e}$.
+Then repeating the process in the proof of Corollary 3, for $e^{e}\leq t\leq 10^{6}$, we obtain
+$\left|\frac{\zeta^{\prime}(1+it)}{\zeta(1+it)}\right|\leq 7.686\log t\leq 40.440\log\log t,$
+with rounded parameters $(\epsilon,\sigma_{1},\eta)=(0.037816,1.395842,3.177141)$. The second inequality above uses the fact that $t$ attains its maximum at $10^{6}$.
+We remark that (29) is an unconditional result due to the partial verification of $RH$.
+While this is nowhere near as sharp as the bounds in [5], it nevertheless holds for a wider range and might be useful for smaller values of $t$.
+If desired, one can extend the range for which the corollary is unconditional beyond this (albeit still in a finite range $t_{0}\leq t\leq T$), at the expense of a worse leading constant.
+This will prove a poor man’s version of [5, Theorem 1.1], but with a simpler method.
+As mentioned in §1, the range of validity is wider but the constants are worse.
+This extension requires a modification in the proof of Theorem 2, by fixing $r=\epsilon+1/W\log T$, utilising a classical zero-free region, and also using Lemma 8 in place of Corollary 7 when estimating $A_{1}$ in (25).
 4.2 Above the Riemann height
-Suppose $t_{0}\geq H$ and let there be a positive real number $d$.
-As before, we construct discs which have origins slightly to the right of the $1$-line, at $s_{0}=\sigma_{0}+iT_{0}$, where $T_{0}\geq t_{0}$,
-$\sigma_{0}=1+\delta_{T_{0}}=1+\frac{d}{\log T_{0}},$ (34)
-and extend their radii slightly into the critical strip, and then apply Theorem 1.
-This time however, we do not have the assumption of $RH(T)$ to help us with potential zeros.
-We will instead utilise the zero-free region mentioned in Section 2.1 and keep our discs within said region.
-This requires discs with radii decreasing to 00 at the rate of $1/\log T_{0}$, similar to the zero-free region.
-Let
-$R_{T_{0}}=2\delta_{T_{0}}=\frac{2d}{\log T_{0}}\qquad\text{and}\qquad r=\alpha R_{T_{0}},\quad 0<\alpha<1.$ (35)
-In addition, let $d$, and hence $R_{T_{0}}$, be chosen such that
-$\sigma_{0}-R_{T_{0}}=\sigma_{0}-2\delta_{T_{0}}\geq 1-\frac{1}{W_{0}\log(T_{0}+R_{T_{0}})},$
-which is fulfilled by
-$d\leq\frac{1}{W_{0}}\left(1+\frac{\log\big{(}1+\frac{R_{H}}{H}\big{)}}{\log H}\right)^{-1}.$ (36)
-This ensures that the disc $|s-s_{0}|\leq R_{T_{0}}$ lies entirely in the zero-free region, so $f(s)$ is analytic there.
-Next, we want to obtain a bound on $M$. This is done by choosing
-$R^{\prime}=\frac{1}{2}-\delta_{H}=\frac{1}{2}-\frac{d}{\log H}$
-in Lemma 2. Then as long as $d$ is chosen to satisfy (3), i.e.,
-$d\leq\frac{1}{W_{0}}\left(1+\frac{\log\big{(}1+\frac{R^{\prime}}{H}\big{)}}{\log H}\right)^{-1},$ (37)
-we have for $\sigma\geq 1-\delta_{H}$,
-$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT_{0})\right|\leq M:=\frac{4}{\pi R^{\prime}}\max_{0\leq\theta\leq 2\pi}|\log|\zeta(\sigma+iT_{0}+R^{\prime}e^{i\theta})||.$ (38)
-Estimating $M$ thus requires a uniform bound of $\zeta(s)$ in the region
-$\frac{1}{2}\leq 1-\delta_{T_{0}}-R^{\prime}\leq\sigma\leq 1+3\delta_{T_{0}}+R^{\prime}\leq 1+\frac{1}{2}+2\delta_{H}.$
-By Lemma 7 with $\delta_{r}=1/2+2\delta_{H}$, we have for $T_{0}\geq H$, and the region $1/2\leq\sigma\leq 3/2+2\delta_{H}$ that
-$|\zeta(s)|\leq A_{2}(T_{0}+R^{\prime})^{1/6}\log(T_{0}+R^{\prime})\leq A_{3}T_{0}^{1/6}\log T_{0},$ (39)
-where
-$A_{3}$ $=A_{2}\left(1+\frac{R^{\prime}}{H}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{H}\right)}{\log H}\right),$ (40)
-with
-$A_{2}$ $=0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{2d}{\log H},1.31,H-R^{\prime}\right)\right)^{7/6}$ (41)
-×$\qquad\qquad\bm{\times}\left(1+a_{0}\left(\frac{3}{2}+\frac{2d}{\log H},1.31,H-R^{\prime}\right)\right).$
-Combining (39), (41), and (40) with (38) thus gives for $1-\delta_{T_{0}}\leq\sigma\leq 1+3\delta_{T_{0}}$,
-$M\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}\log T_{0}+\log\log T_{0}+\log A_{3}\right).$ (42)
-Now to bound $|\zeta^{\prime}(s_{0})/\zeta(s_{0})|$, we use the trivial bound for $\sigma>1$, given in (31). This yields
-$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq\frac{1}{\sigma_{0}-1}\leq\frac{\log T_{0}}{d}.$ (43)
-Up to this point, Theorem 1 will give a result that holds for
-$1-\delta_{T_{0}}(2\alpha-1)\leq\sigma\leq 1+\delta_{T_{0}}(2\alpha+1),$ (44)
-and for a result on any larger $\sigma$, we simply use (31). At the same time, if we want a result that holds for $\sigma\geq 1-1/W\log T_{0}$, then by the left inequality of (44), we must have
-$W=\frac{1}{d(2\alpha-1)}\qquad\text{and}\qquad\alpha=\frac{1}{2}+\frac{1}{2dW}.$ (45)
-Finally, putting everything together and making the substitution $T_{0}\mapsto t$, we prove our result.
+$W$ $Q$ $\beta$ $\epsilon_{1}$ $\sigma_{1}$ $\eta$
+$5.559$ $928465$ $0.999980$ $0.012554$ $1.756408$ $1.050104$
+$5.56$ $170199$ $0.999890$ $0.012551$ $1.937703$ $2.228745$
+$5.5601$ $157356$ $0.999881$ $0.012563$ $1.393177$ $4.286069$
+$5.5602$ $146346$ $0.999872$ $0.012587$ $1.710791$ $1.367807$
+$5.5603$ $136750$ $0.999863$ $0.012582$ $1.776753$ $1.141731$
+$5.5604$ $128379$ $0.999854$ $0.012712$ $1.889117$ $2.700506$
+$5.5605$ $120884$ $0.999845$ $0.012614$ $1.711461$ $3.302948$
+$5.561$ $93736$ $0.999800$ $0.012690$ $1.339371$ $2.627981$
+$5.562$ $64686$ $0.999710$ $0.012842$ $1.299159$ $2.599102$
+$5.563$ $49364$ $0.999620$ $0.012699$ $1.427545$ $1.952944$
+$5.564$ $39919$ $0.999531$ $0.012747$ $1.557141$ $2.691489$
+$5.565$ $33530$ $0.999441$ $0.013347$ $1.715049$ $4.337899$
+$5.566$ $28887$ $0.999351$ $0.013295$ $1.234376$ $4.393614$
+$5.567$ $25383$ $0.999262$ $0.013634$ $1.973530$ $3.744506$
+$5.568$ $22625$ $0.999172$ $0.013225$ $1.814250$ $2.845626$
+$5.569$ $20414$ $0.999082$ $0.013180$ $1.641377$ $1.985693$
+$5.57$ $18592$ $0.998993$ $0.012868$ $1.322806$ $4.407476$
+$5.575$ $12870$ $0.998545$ $0.012963$ $1.920818$ $3.555767$
+$5.58$ $9863$ $0.998102$ $0.019214$ $1.690970$ $2.264745$
+$5.59$ $6698$ $0.997207$ $0.012557$ $1.929368$ $4.355523$
+$5.6$ $5082$ $0.996321$ $0.015401$ $1.421225$ $3.490106$
+$5.7$ $1501$ $0.987620$ $0.025973$ $1.590960$ $3.434482$
+$5.8$ $888.269$ $0.979209$ $0.018833$ $1.305491$ $4.203009$
+$5.9$ $635.099$ $0.971088$ $0.021934$ $1.977612$ $4.803605$
+66 $496.670$ $0.963240$ $0.027122$ $1.420679$ $3.567866$
+$6.1$ $409.353$ $0.955640$ $0.016196$ $1.713398$ $1.710689$
+$6.2$ $349.287$ $0.948293$ $0.019949$ $1.623774$ $3.117029$
+$6.3$ $305.581$ $0.941204$ $0.066228$ $1.701866$ $3.427829$
+$6.4$ $272.026$ $0.934286$ $0.023358$ $1.163604$ $3.125341$
+$6.5$ $246.170$ $0.927735$ $0.255900$ $1.454736$ $3.093587$
+$6.75$ $199.284$ $0.911787$ $0.057471$ $1.361313$ $4.269028$
+77 $168.924$ $0.897081$ $0.059017$ $1.777110$ $3.321819$
+88 $109.668$ $0.847438$ $0.038048$ $1.707618$ $2.623961$
+99 $84.858$ $0.808831$ $0.031281$ $1.249275$ $2.976333$
+$10$ $71.220$ $0.777942$ $0.016334$ $1.624690$ $4.127955$
+$11$ $62.611$ $0.752706$ $0.084730$ $1.177847$ $1.792189$
+$12$ $56.653$ $0.731621$ $0.020106$ $1.606738$ $2.688990$
+$13$ $52.306$ $0.713814$ $0.041793$ $1.671118$ $3.367414$
+Table 3: Values for $Q$ in Corollary 1, with corresponding $W$ and parameters $\beta$, $\epsilon_{1}$, $\sigma_{1}$, $\eta$ (all rounded to six decimal places). Each entry is valid for $t\geq 13$.
 Theorem 3.
-Let $d$ and $W>W_{0}$ be positive parameters, subject to (36), (37), and (45). Let also $10<\alpha<$ 1 be a parameter defined in (45), and depending on $d$ and $W$. In a region
-$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H$
+Let $d\leq 1/W_{0}$, $\epsilon_{1}\leq 1/2$, and $\beta<1$ be positive real parameters, subject to (35) and (38). Define $W>W_{0}$ by
+$W:=\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)^{-1},$ (30)
+where the function $a_{\epsilon_{1}}(t_{0})$ also depends on $\epsilon_{1}$ and is defined in (34).
+Then in a region
+$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq t_{0}\geq H$
 where $\zeta(\sigma+it)\neq 0$, we have
@@
 where
-$Q_{H}=\max\left\{\frac{\lambda}{6}+\frac{1+\alpha}{d(1-\alpha)}+\frac{\lambda(\log\log H+\log A_{3})}{\log H},\,\frac{1}{d(1+2\alpha)}\right\},$
-with $A_{3}$ defined in (40),
-$\lambda=\frac{8\alpha}{(1-\alpha)\pi R^{\prime}},\qquad\text{and}\qquad R^{\prime}=\frac{1}{2}-\frac{d}{\log H}.$
+$Q_{H}=\max\left\{\lambda_{1}\left(\frac{1}{6}+\frac{2\log\log t_{0}}{\log t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)+\lambda_{2},\,\frac{1}{(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d)}\right\},$
+with
+$\lambda_{1}=\frac{16\beta}{(1-\beta)}\left(1-\frac{8d}{2d+\log t_{0}}\right)^{-2}\qquad\text{and}\qquad\lambda_{2}=\frac{1+\beta}{d(1-\beta)},$ (31)
+provided that
+$t^{1/6}\log t\geq\frac{\zeta(\frac{3}{2}+\frac{2d}{\log t_{0}})}{0.618(1+a_{1}(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}))^{1/6}}$ (32)
+is satisfied.
+The quantity $A_{3}$ is defined in (4.2), and depends on $d$ and $t_{0}$, on $a_{0}$ and $a_{1}$, which are defined as in Lemma 4, and on $V(\kappa,\sigma_{1},t_{0},\eta)$ and its corresponding parameters, defined as in Lemma 12 and subject to (37).
+Proof.
+The proof is similar to that of Theorem 2.
+Suppose $t_{0}\geq H$ and let there be a positive real parameter $d\leq 1/W_{0}$ to be determined later.
+As before, we construct discs which have origins slightly to the right of the $1$-line, at $s_{0}=\sigma_{0}+iT_{0}$, where $T_{0}\geq t_{0}$,
+$\sigma_{0}=1+\frac{d}{\log T_{0}},\qquad r=\frac{1}{2}+\frac{d}{\log T_{0}},$ (33)
+extend their radii slightly into the critical strip, and then apply Lemma 2 with $f(s)=\zeta(s)$, $s=\sigma+iT_{0}$, and $A_{1}$ defined by (25).
+This time however, we do not have the assumption of $RH(T)$ to help us with potential zeros.
+We will instead utilise the zero-free region mentioned in §2.1 and keep our discs within said region.
+This requires discs with radii decreasing to 0 at the rate of $1/\log T_{0}$, similar to the zero-free region.
+Let $\epsilon_{1}$ be a fixed positive real parameter, and for the sake of brevity define the decreasing function (for $T_{0}>1$)
+$a_{\epsilon_{1}}(T_{0}):=1+\frac{\log\left(1+\frac{\epsilon_{1}}{\log T_{0}}\right)}{\log T_{0}}>1.$ (34)
+To ensure that the non-vanishing condition on $f(s)=\zeta(s)$ is met, notice that if we impose the condition $\alpha r\leq\epsilon_{1}\leq r$, then the disc $|s-s_{0}|\leq\alpha r$ with
+$\alpha r\leq\sigma_{0}-\left(1-\frac{d}{\log(T_{0}+\epsilon_{1})}\right),$
+is completely contained in the zero-free region as defined in §2.1. This also implies that the following choice of $\alpha$ and bounds are admissible:
+$\alpha:=\frac{2d\left(1+a_{\epsilon_{1}}(t_{0})^{-1}\right)}{\log T_{0}+2d}<\frac{4d}{2d+\log t_{0}}\qquad\text{and}\qquad\frac{2d}{\log t_{0}}\leq\epsilon_{1}\leq\frac{1}{2},$ (35)
+which in turn requires that $d<(\log t_{0})/6$ since we need $\alpha<1/2$ for Lemma 2. This is easily achieved since $d\leq 1/W_{0}<(\log t_{0})/6$ for all $t_{0}\geq 3$.
+Next, we determine $A_{1}$ and $A_{2}$.
+Suppose (6) is satisfied with $1+\delta_{r}=\sigma_{0}+r$ and define $A_{1}$ (which depends on $A_{3}$) by (25), but use (33) with the argument of Lemma 12 to replace (4.1) by
+$A_{3}:=0.618\left(1+a_{0}\left(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}\right)\right)\left(1+a_{1}\left(\frac{3}{2}+\frac{2d}{\log t_{0}},1.31,t_{0}\right)\right)^{7/6}$
+$\bm{\cdot}\left(1+\frac{1}{2t_{0}}+\frac{d}{t_{0}\log t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{1}{2t_{0}}+\frac{d}{t_{0}\log t_{0}}\right)}{\log t_{0}}\right)V\left(\frac{d}{\log t_{0}},\sigma_{1},t_{0},\eta\right),$ (36)
+where $\sigma_{1}$ and $\eta$ satisfies
+$1+\frac{d}{\log t_{0}}\leq\sigma_{1}\leq 2\qquad\text{and}\qquad 0<\eta<2t_{0}.$ (37)
+On the other hand, $A_{2}$ is determined by invoking (28), which gives us
+$\left|\frac{\zeta^{\prime}(s_{0})}{\zeta(s_{0})}\right|\leq A_{2}:=\frac{1}{d}\log T_{0}.$
+Now, for any results from Lemma 2 to be meaningful for us, we require the left most point of $|s-s_{0}|\leq\alpha\beta r$ to lie within the zero-free region and in the critical strip.
+In other words, $\sigma_{0}-\alpha\beta r\leq 1$, which implies
+$\beta\geq\frac{2d}{\alpha(2d+\log T_{0})}\geq\left(1+\frac{1}{a_{\epsilon_{1}}(t_{0})}\right)^{-1}\geq\frac{1}{2}.$ (38)
+In particular, since
+$\sigma_{0}-\alpha\beta r=1-\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)\frac{1}{\log T_{0}},$
+to obtain a result that holds for
+$1-\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)\frac{1}{\log T_{0}}\leq\sigma\leq 1+\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d\right)\frac{1}{\log T_{0}},$
+we choose $W>W_{0}$ such that
+$W:=\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})-d\right)^{-1},$
+while to cover the remaining range of $\sigma\geq 1+\left(d\beta(1+a_{\epsilon_{1}}(t_{0})^{-1})+d\right)/\log T_{0}$, we again use (28), and take the maximum of the two results.
+Finally, putting everything together into Lemma 2 and making the substitution $T_{0}\mapsto t$, we prove our result. ∎
 4.3 Combining both cases $t\leq H$ and $t\geq H$
 Proof of Corollary 1.
-For the case of $t\geq H$, with a fixed choice of $W$ as defined in (45), it remains to optimise $Q_{H}$ in Theorem 3 over $\alpha$ and $d$, subject to the conditions mentioned in said theorem, while also checking that (11) holds.
-We found that the optimal value for $d$ in each case was approximately $d=0.179898\ldots$, and any differences in the final result caused by varying $d$ further occurred only after the sixth decimal place.
-We note also that the optimal $d$ is one that is as close to $1/W_{0}$ as possible while still satisfying (36) and (37).
-On the other hand, a simple calculation shows that
-$0.821-\frac{1}{W_{0}\log t}>$ 0.82
-for $t\geq t_{0}=e$. Now, for $\sigma\geq 0.8$ and $5\leq t\leq H$, Corollary 3 and Table 2 asserts that
-$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=33.249\log t,$
+For the case of $t\geq H$, with a fixed choice of $W$ as defined in (30), it remains to optimise $Q_{H}$ in Theorem 3 over $d$, $\beta$, $\epsilon_{1}$, $\sigma_{1}$, and $\eta$, subject to the conditions mentioned in said theorem, while also checking that (32) holds, which is done in a similar way as in the proof of Corollary 3.
+We found that the optimal value for $d$ in all cases is $d=1/W_{0}$, and (32) holds for all $t\geq t_{0}\geq H$.
+All computed parameters are presented in Table 3, rounded to six decimal places.
+On the other hand, a simple calculation shows that for $t\geq t_{0}=13$,
+$1-\frac{1}{W_{0}\log t}>0.929.$
+Now, for $\sigma\geq 0.9$ and $13\leq t\leq H$, Corollary 3 and Table 2 asserts that
+$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=13.123\log t,$
 which is smaller than any value of optimised $Q_{H}$ found. Thus this same value of $Q_{H}$ holds for all $t\geq 13$. ∎
@@
 Proof of Corollary 2.
-First note that Corollary 3 and Table 2 has for $\sigma\geq 1$ and $5\leq t\leq H$,
-$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=9.161\log t.$
-Then for $t\geq H$, the proof is identical to that of Theorem 3 and Corollary 1, except we can restrict the size of $\alpha$ even further, since we now require $\sigma_{0}-\alpha R_{T_{0}}\geq 1$, which is equivalent to $\alpha\leq 1/2$.
-In order for the result to hold for all $\sigma\geq 1$, we choose $\alpha=1/2$, so in this case,
-$Q_{H}=\max\left\{\frac{8}{6\pi R^{\prime}}+\frac{3}{d}+\frac{8(\log\log H+\log A_{3})}{\pi R^{\prime}\log H},\,\frac{1}{2d}\right\}.$
-As before, we then optimise over $d$, while ensuring that we fulfil the conditions in Theorem 3 and (11). The optimal value found was $d=0.179898$. ∎
-$W$ $Q$ $\alpha$ $W$ $Q$ $\alpha$
-$5.559$ $450041$ $0.999972$ $5.59$ $4458$ $0.997200$
-$5.56$ $106223$ $0.999882$ $5.6$ $3383$ $0.996312$
-$5.5601$ $98674$ $0.999873$ $5.7$ $1002$ $0.987604$
-$5.5602$ $92137$ $0.999864$ $5.8$ $594.087$ $0.979198$
-$5.5603$ $86414$ $0.999855$ $5.9$ $425.318$ $0.971076$
-$5.5604$ $81354$ $0.999846$ $6$ $333.034$ $0.963224$
-$5.5605$ $76859$ $0.999837$ $6.1$ $274.847$ $0.955630$
-$5.561$ $60223$ $0.999792$ $6.2$ $234.806$ $0.948282$
-$5.562$ $42026$ $0.999703$ $6.3$ $205.568$ $0.941166$
-$5.563$ $32278$ $0.999613$ $6.4$ $183.281$ $0.934273$
-$5.564$ $26201$ $0.999523$ $6.5$ $165.729$ $0.927592$
-$5.565$ $22051$ $0.999433$ $6.75$ $134.741$ $0.911755$
-$5.566$ $19036$ $0.999343$ $7$ $114.502$ $0.897049$
-$5.567$ $16748$ $0.999254$ $8$ $74.999$ $0.847418$
-$5.568$ $14951$ $0.999164$ $9$ $58.454$ $0.808816$
-$5.569$ $13502$ $0.999074$ $10$ $49.359$ $0.777935$
-$5.57$ $12310$ $0.998985$ $11$ $43.607$ $0.752668$
-$5.575$ $8541$ $0.998537$ $12$ $39.642$ $0.731612$
-$5.58$ $6542$ $0.998091$ $13$ $36.742$ $0.713796$
-Table 3: Values for $Q$ in Corollary 1, with corresponding $W$ and $\alpha$ (rounded to six decimal places). Each entry is valid for $t\geq 5$.
+First note that Corollary 3 and Table 2 has for $\sigma\geq 1$ and $13\leq t\leq H$,
+$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{RH(H_{0})}\log t=8.101\log t.$ (39)
+Then for $t\geq H$, the proof is identical to that of Theorem 3 and Corollary 1, except we can restrict the size of $\alpha\beta r$ even further, since we can now set $\sigma_{0}-\alpha\beta r=1$, which is equivalent to
+$\beta=(1+a_{\epsilon_{1}}(t_{0})^{-1})^{-1}.$ (40)
+So in this case, for $\sigma\geq 1$ and $t\geq t_{0}\geq H$, we have that
+$\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq\max\left\{\lambda_{1}\left(\frac{1}{6}+\frac{2\log\log t_{0}}{\log t_{0}}+\frac{\log A_{3}}{\log t_{0}}\right)+\lambda_{2},\,\frac{1}{2d}\right\}\,\log t,$ (41)
+where $\lambda_{1}$ and $\lambda_{2}$ are defined as in (31), but with the added restriction of (40).
+Similar to the proof of Corollary 1, we then optimise (41) over $d$, $\epsilon_{1}$, $\sigma_{1}$, and $\eta$, while ensuring that we fulfil the appropriate conditions in Theorem 3 and (32).
+The optimal value of $d$ found was again $d=1/W_{0}$.
+We then compare the result with (39) and take the maximum to obtain a bound that holds for all $t\geq 13$.
+∎
 5 Estimates for $|1/\zeta(s)|$
@@
 for some $\delta_{1}>0$.
-The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [3, Proposition A.2], [19]), while in the past the above integral has been estimated by bounding the real part of the logarithmic derivative by a uniform bound $|\zeta^{\prime}(s)/\zeta(s)|$.
-Here some savings can be made by splitting up the interval $[\sigma,1+\delta_{1}]$ and using the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval (see [7, Lemma 10]).
-However, we show that by using directly a bound on the real part instead, through Lemma 2, we are able to do much better.
-Furthermore, through a trick utilising the trigonometric polynomial, we are able to get an slight asymptotic improvement on the bounds, at the cost of a slightly worse constant.
-As before, the proofs will be split into two cases, $t\leq H$ and $t\geq H$.
-5.1 Assuming a partial Riemann hypothesis
+The estimation of the first term on the right-hand side can be improved using the trigonometric polynomial (see [3, Proposition A.2], [20]), while the above integral is estimated by bounding the real part of the logarithmic derivative by a uniform bound on its absolute value $|\zeta^{\prime}(s)/\zeta(s)|$.
+Here some savings can be made by using the method of [7, Lemma 10]), which splits the interval $\sigma,1+\delta_{1}]$ and uses the sharpest bounds for $|\zeta^{\prime}(s)/\zeta(s)|$ in the appropriate range, rather than one uniform bound for the entire interval.
+Furthermore, through an argument utilising the trigonometric polynomial, we are able to get a slight asymptotic improvement on the bounds, at the cost of a worse constant.
 Theorem 4.
-Assume $RH(T)$.
-Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log t_{0}\leq\sigma_{1}\leq 2$.
-Suppose also $R^{\prime}=1/2-d/\log t_{0}$.
-Then in the region
-$\sigma\geq 1-\frac{1}{W\log t}\qquad\text{and}\qquad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
+Let $t_{0}\geq 13$ and $d_{1}>0$ be a constant. Let $\{W_{j}:1\leq j\leq J\}$ be a sequence of increasing real numbers where $(W_{j},Q_{1,j})$ is a pair for which $t\geq t_{0}$ and
+$\sigma\geq 1-\frac{1}{W_{j}\log t}\quad\text{implies}\quad\left|\frac{\zeta^{\prime}(\sigma+it)}{\zeta(\sigma+it)}\right|\leq Q_{1,j}\log t.$
+Then for
+$\sigma\geq 1-\frac{1}{W_{1}\log t},$
 we have
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{RH(T)}(d,d_{1},\sigma_{1},t_{0},\eta)\log t,$
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{0}(d_{1},\sigma_{1},t_{0},\eta)\log t,$
 where
-$Y_{RH(T)}$ $(d,d_{1},\sigma_{1},t_{0},\eta)=\max\Bigg{\{}\frac{1}{d_{1}}+\frac{1}{\log t_{0}},$
-$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\Bigg{\}},$
-with
-$A_{4}$ $=\left(1+\frac{R^{\prime}}{t_{0}}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{t_{0}}\right)}{\log t_{0}}\right)\left(1+a_{0}\left(\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},1.31,t_{0}-R^{\prime}\right)\right)$
-×$\qquad\qquad\bm{\times}0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},1.31,t_{0}-R^{\prime}\right)\right)^{7/6},$
-$B_{t_{0}}$ defined in Theorem 2, and $V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta),\,\eta$ defined in Lemma 12.
-Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
+$Y_{0}$ $(d_{1},\sigma_{1},t_{0},\eta)=\max\Biggl{\{}\frac{1}{d_{1}}+\frac{1}{\log t_{0}},$ (43)
+$V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta)\cdot\exp\Bigg{(}\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}\Bigg{)}\Biggr{\}},$
+with any $1+d_{1}/\log t_{0}\leq\sigma_{1}\leq 2$, $\gamma$ denoting Euler’s constant, and $V$ defined as in Lemma 12.
 Proof.
-Let $\delta_{t}:=d/\log t$ and $\delta_{1,t}:=d_{1}/\log t$, and first suppose $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
-Since by $RH(T)$ there are no zeros for $\sigma\geq 1/2$, we do not have to check the condition in (3).
-The same argument in Section 4.1 with Lemma 2 leading to (30), requires a uniform bound on $|\zeta(s)|$ in the region
-$\frac{1}{2}\leq 1-\delta_{t}-R^{\prime}\leq\sigma\leq 1+\delta_{1,t}+R^{\prime}\leq\frac{3}{2}+\frac{d_{1}-d}{\log t_{0}},$
-and in return gives us
-$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\log t$
-uniformly for $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. Thus we have from (46) that
-$\log\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\log\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|+\int_{1-\delta_{t}}^{1+\delta_{1,t}}\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(x+it)\right|\text{d}x,$ (47)
-where the integral is bounded from above by
-$\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right).$
-To bound the first term on the right-hand side of (47), we use Lemma 12 with $\kappa=\delta_{1,t}=d_{1}/\log t$ to see that
-$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|\leq V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$ (48)
-Hence (47) becomes
-$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$
-for $t_{0}\leq t\leq T-1/2$ and $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
-Now we turn to the case $\sigma\geq 1+\delta_{1,t}$. Here we use the trivial bound
-$|\zeta(s)|\leq\zeta(\sigma)\leq\frac{\sigma}{\sigma-1},\qquad(\sigma>1),$ (49)
-where $\sigma/(\sigma-1)$ is decreasing in $\sigma$. Together with the trigonometric inequality from (25), we see that
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\zeta(1+\delta_{1,t})\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t.$ (50)
-We reiterate that although there are sharper bounds available to use here, this is sufficient for our purposes as the case where $\sigma\leq 1+\delta_{1,t}$ is dominant when computing constants.
-Note that the results proved hold also if one assumes $RH$, and allow us to remove the upper bound restriction on $t$, as explained at the end of the proof of Theorem 2. ∎
-5.2 Above the Riemann height
-Now we consider $t\geq H$ and prove the following.
-Theorem 5.
-Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log H\leq\sigma_{1}\leq 2$.
-Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37).
-Then in the region
-$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
-we have
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{H}(d,d_{1},\sigma_{1},\eta)\log t,$
-where
-$Y_{H}$ $(d,d_{1},\sigma_{1},\eta)=\max\Bigg{\{}\frac{1}{d_{1}}+\frac{1}{\log H},$
-$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}V\left(\frac{d_{1}}{\log H},\sigma_{1},H,\eta\right)\Bigg{\}},$
-with
-$A_{5}$ $=\left(1+\frac{R^{\prime}}{H}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{H}\right)}{\log H}\right)\left(1+a_{0}\left(\frac{3}{2}+\frac{d_{1}-d}{\log H},1.31,H-R^{\prime}\right)\right)$
-×$\qquad\qquad\bm{\times}0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{d_{1}-d}{\log H},1.31,H-R^{\prime}\right)\right)^{7/6},$
-and $V(d_{1}/\log H,\sigma_{1},H,\eta),\,\eta$ are defined in Lemma 12. Admissible values of $Y_{H}$ with corresponding $W$ are presented in Table 4.
-Proof.
-The proof is analogous to that of Theorem 4.
-Let $\delta_{t}:=d/\log t$ and $\delta_{1,t}:=d_{1}/\log t$, and first suppose $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
-As we do not assume $RH(T)$, we have to check that (3) holds, and we do this by our hypothesis of $d$ satisfying (37).
-The same argument in Section 4.2 with Lemma 2 leading to (42), requires a uniform bound on $|\zeta(s)|$ in the region
-$\frac{1}{2}\leq 1-\delta_{t}-R^{\prime}\leq\sigma\leq 1+\delta_{1,t}+R^{\prime}\leq\frac{3}{2}+\frac{d_{1}-d}{\log H},$
-and in return gives us
-$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\log t$
-uniformly for $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
-The rest of the proof then follows that of Theorem 4, and it remains to optimise $Y_{H}$ for a fixed choice of $W=1/d$, over $d_{1}$, $\sigma_{1}$ and $\eta$, subject to the conditions mentioned in the statement of the theorem, while also checking that (11) holds.
-∎
-$W$ $Y_{H}$ $d_{1}$ $\sigma_{1}$ $\eta$
-$5.5587$ $2.307$ $1.059544$ $1.412209$ $3.464102$
-$6$ $2.284$ $1.060503$ $1.608124$ $3.775135$
-$7$ $2.244$ $1.062065$ $1.604395$ $3.772048$
-$8$ $2.214$ $1.063547$ $1.895304$ $3.621918$
-$9$ $2.191$ $1.064570$ $1.399356$ $3.464227$
-$10$ $2.173$ $1.065554$ $1.135564$ $3.65928$
-$11$ $2.158$ $1.065999$ $1.43905$ $3.446264$
-$12$ $2.146$ $1.066605$ $1.345521$ $3.769427$
-$13$ $2.136$ $1.067103$ $1.345557$ $3.77319$
-Table 4: Values for $Y_{H}$ in Theorem 5, with corresponding $W$, and parameters (rounded to six decimal places) $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq H$.
-5.3 Combining both cases $t\leq H$ and $t\geq H$
+Let $\delta_{1}=\delta_{1}(t):=d_{1}/\log t$.
+We will split the proof into two cases: $\sigma\leq 1+\delta_{1}$ and $\sigma>1+\delta_{1}$.
+First see that in the range $1-1/W_{1}\log t\leq\sigma\leq 1+\delta_{1}$ we have
+$\log\left|\frac{1}{\zeta(\sigma+it)}\right|$ $=-\textup{Re}\log\zeta\left(1+\delta_{1}+it\right)+\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x.$ (44)
+Writing $\Delta=1/\log t$, we can split the integral and rewrite it as
+$\left(\int_{1-\tfrac{\Delta}{W_{1}}}^{1-\tfrac{\Delta}{W_{2}}}+\cdots+\int_{1-\tfrac{\Delta}{W_{j}}}^{1-\tfrac{\Delta}{W_{j+1}}}+\cdots+\int_{1-\tfrac{\Delta}{W_{J}}}^{1}+\int_{1}^{1+\delta_{1}}\right)\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x.$ (45)
+By Corollary 2 and our assumptions we have
+$\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x$ $\leq\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}.$
+We therefore have, from (44),
+$\log\left|\frac{1}{\zeta(\sigma+it)}\right|<$ $-\log\left|\zeta\left(1+\frac{d_{1}}{\log t}+it\right)\right|$ (46)
+$+\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}.$
+To estimate the first term on the right-hand side of (46) we apply Lemma 12, keeping in mind our restrictions on $\sigma_{1}$. This gives us
+$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta)\log t.$ (47)
+We combine this with (46) and exponentiate to obtain $Y_{0}$ for $\sigma\leq 1+\delta_{1}$.
+For the case $\sigma>1+\delta_{1}$ we have from the trivial bound $|1/\zeta(s)|\leq\zeta(\sigma)$ with (9) that
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t,$ (48)
+since $\sigma/(\sigma-1)$ is decreasing in $\sigma$.
+Taking the maximum from both cases completes the proof. ∎
+Corollary 4 follows from Theorem 4 by computing $Y_{0}$ for specific choices of $t_{0}$ and $W_{j}$, and optimizing over $d_{1}$, $\sigma_{1}$, and $\eta$.
+These computed values of $Y_{0}$ are labelled $Y$, and are presented in Table 4, with its corresponding parameters (rounded to six decimal places).
+The process is the same as that of [7, Corollary 3], but for completeness sake, we repeat it here.
 Proof of Corollary 4.
-For the first assertion of the corollary, we begin by assuming $RH(H_{0})$ and dealing with the case $t_{0}\leq t\leq H$.
-To obtain $Y(W,5)$, fix $t_{0}$ and $W=1/d$ in Theorem 4 and optimise $Y_{RH(H_{0})}$ over $d_{1}$, $\sigma_{1}$, and $\eta$.
-At the same time ensure that the restrictions set in Theorem 4 and Corollary 8 hold, and raise $t_{0}$ if (11) is not satisfied.
-We find that the smallest value of $t_{0}$ such that (11) is fulfilled for all entries in Tables 6 and 5 is $t_{0}=5$.
-Then we compare the resulting values with the corresponding $Y_{H}$ (for $t\geq H$) in Theorem 5 and Table 4, taking the larger of the two.
-[…]
-For the second assertion, i.e., when $\sigma=1$, we set $d=0$ and repeat the above process using Theorems 4 (under $RH(H_{0})$) and 5 (unconditionally).
-This allows us to obtain a result for $\sigma\geq 1$.
-We have
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 2.018\log t\quad\text{for}\quad t\geq H,$ (51)
-[…]
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 2.751\log t\quad\text{for}\quad 2\exp(e^{2})\leq t\leq H,$ (52)
-[…]
-Finally, combine (51) and (52) with the following two bounds. First, we utilise Lemma 7, which was proven using numerical methods. Taking $\sigma=1$ in said lemma, we derive
-$\left|\frac{1}{\zeta(1+it)}\right|\leq\frac{8}{\log(500)}\log t\leq 1.288\log t\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}).$
-Second, a computation from the proof of [3, Proposition A.2], implemented in interval arithmetic, states
-$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.079\log t\quad\text{for}\quad 2\leq t\leq 500.$
-This completes the proof. ∎
-When $\sigma$ is fixed (in this case $\sigma=1$), the limiting factor for the constant in the second assertion of Corollary 4 comes from the case $2\exp(e^{2})\leq t\leq H$.
-In particular, the lower threshold of $2\exp(e^{2})$ is what determines the constant $2.751$.
-Using numerical methods, it is possible to push this threshold higher and hence improve (52).
-To illustrate this idea, following the proof of Corollary 4, we can prove unconditionally that for $t\geq 10^{6}$,
-$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.391\log t,$ (53)
-which is an improvement over what one gets from [5, Theorem 1.1].
-We make the remark that we include our computed values of $\sigma_{1}$ and $\eta$ in Tables 4, 5, and 6, for completeness sake.
-In practice, as we limit the overall constant to three decimal places, $d_{1}$ is the main parameter we optimise over.
-For $\sigma_{1}$ and $\eta$, any sufficiently large values will do, and further optimisations here will only improve the final constant from the fourth decimal place onwards.
-$t_{0}$ $Y(5.5587,t_{0})$ $d_{1}$ $\sigma_{1}$ $\eta$
-$10$ $8.101$ $0.307606$ $1.330692$ $3.068541$
-$10^{2}$ $4.339$ $0.542825$ $1.851563$ $3.417699$
-$10^{3}$ $3.632$ $0.648814$ $1.282661$ $3.465621$
-$10^{4}$ $3.264$ $0.725026$ $1.748467$ $3.463629$
-$10^{5}$ $3.021$ $0.787049$ $1.208003$ $3.464054$
-$10^{7}$ $2.711$ $0.885011$ $1.101137$ $3.464109$
-$10^{9}$ $2.518$ $0.960512$ $1.272668$ $3.445896$
-$10^{11}$ $2.383$ $1.021396$ $1.655004$ $3.464097$
-Table 5: Values for $Y(5.5587,t_{0})$ in Corollary 4, with corresponding $t_{0}$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$.
-Each entry is valid for $t\geq t_{0}$ and $W=5.5587$.
-$W$ $Y(W,5)$ $d_{1}$ $\sigma_{1}$ $\eta$
-$5.5587$ $18.060$ $0.170031$ $1.105647$ $2.784753$
-$6$ $16.722$ $0.173040$ $1.107519$ $2.781151$
-$7$ $14.654$ $0.178370$ $1.110828$ $2.783844$
-$8$ $13.339$ $0.182334$ $1.113292$ $2.786533$
-$9$ $12.433$ $0.185403$ $1.115198$ $2.782321$
-$10$ $11.771$ $0.187839$ $1.116712$ $2.782458$
-$11$ $11.268$ $0.189804$ $1.117933$ $2.786745$
-$12$ $10.872$ $0.191448$ $1.118954$ $2.786457$
-$13$ $10.553$ $0.192838$ $1.119819$ $2.786611$
-Table 6: Values for $Y(W,5)$ in Corollary 4, with corresponding $W$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$.
-Each entry is valid for $t\geq t_{0}=5$.
+After choosing $t_{0}=13$, we aim to minimise $Y_{0}$ by optimising over $d_{1}>0$, $\eta$, and $\sigma_{1}$, while making sure that the conditions on these parameters are fulfilled.
+This optimisation process first requires fixing the desired $W_{1}$, which then tells us the number of $W_{j}$ we are summing over, since from Table 3 we always have $(W_{J},Q_{1,J})=(13,52.306)$.
+Computing more values of $(W_{J},Q_{1,J})$ will give a better result, but the improvements eventually become negligible, as $Y_{0}$ is largely determined by the initial few $W_{j}$.
+For our computations, we only used values for $W_{j}=W$ from Table 3 (Corollary 1), thus we require $t_{0}=13$.
+Note that one can adjust Theorem 4 to compute $Y$ for the case $\sigma\geq 1$. In that case, we will modify (43) by setting
+$\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}=0$
+before optimising. We obtain $Y=30.812$ for $\sigma\geq 1$ and $t\geq 13$, with $(d_{1},\sigma_{1},\eta)=(0.032871,1.662479,3.216997)$. This proves the second assertion of the corollary.
+To prove the third assertion, we first compute $Y=29.388$ for $\sigma\geq 1$ and $t\geq 2\exp(e^{2})$, with $(d_{1},\sigma_{1},\eta)=(0.034172,1.286253,3.718165)$. Then see that from Lemma 7,
+$\left|\frac{1}{\zeta(1+it)}\right|\leq\frac{8}{\log(500)}\log t\leq 1.288\log t\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}),$
+while a computation from the proof of [3, Proposition A.2], implemented in interval arithmetic, states that
+$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.079\log t\quad\text{for}\quad 2\leq t\leq 500,$ (49)
+so we are done. ∎
+$W$ $Y$ $d_{1}$ $\sigma_{1}$ $\eta$
+$5.56$ $7.59\cdot 10^{31}$ $0.030648$ $1.151072$ $3.150608$
+66 $2.46\cdot 10^{9}$ $0.030648$ $1.150464$ $3.150009$
+77 $1.83\cdot 10^{6}$ $0.030647$ $1.151694$ $3.149136$
+88 $89357$ $0.030647$ $1.148849$ $3.149156$
+99 $19482$ $0.030647$ $1.150203$ $3.150832$
+$10$ $7589$ $0.030648$ $1.149868$ $3.150556$
+$11$ $3972$ $0.030650$ $1.150918$ $3.151137$
+$12$ $2472$ $0.030646$ $1.148656$ $3.150041$
+$13$ $1719$ $0.030647$ $1.149567$ $3.150198$
+Table 4: Values for $Y$ in Corollary 4, with corresponding $W$, and parameters (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq 13$.
 5.1 An asymptotic improvement
-In this section we prove an asymptotic improvement over the bounds given earlier in Section 5, at the expense of a worse constant.
-For the sake of clarity, we first outline the general idea here, before proving the next two theorems.
-Assume we have the bound
+In this section we prove an asymptotic improvement over the bounds given earlier in §5, at the expense of a worse constant.
+For the sake of clarity, we first outline the general idea here, before proving the next theorem.
+Suppose we have the bound
 $\max_{1-\delta\leq\sigma\leq 1+\delta_{1}}\left|{\textup{Re}}\left(\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right)\right|\leq C\log t,$
@@
 $\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq|\zeta(1+\delta_{1})|^{3/4}|\zeta(1+\delta_{1}+2it)|^{1/4},$ (51)
-and at the expense of a worse constant, we can estimate the second factor in (55) differently from the previous sections.
+and at the expense of a worse constant, we can estimate the second factor in (51) differently from Theorem 4.
 We do this by using Lemma 6, which states that $|\zeta(s)|\leq C_{3}(\log t)^{2/3}$ for $\sigma\geq 1$.
-Along with Lemma 9, this leads to
+Along with the second inequality in Lemma 9 and our choice of $\delta_{1}$, this leads to
 $\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq(C_{4}\log t)^{3/4}(C_{3}(\log t)^{2/3})^{1/4},$
@@
 Theorem 5.
-Assume $RH(T)$.
-Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log t_{0}\leq 1$.
-Suppose also $R^{\prime}=1/2-d/\log t_{0}$.
-Then in the region
-$\sigma\geq 1-\frac{1}{W\log t}\qquad\text{and}\qquad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
+Let the hypotheses of Theorem 4 be satisfied. Then for $t\geq t_{0}\geq 13$ and
+$\sigma\geq 1-\frac{1}{W_{1}\log t},$
 we have
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{RH(T)}(d,d_{1},t_{0})(\log t)^{11/12},$
+$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{0}(d_{1},t_{0})(\log t)^{11/12},$
 where
-$Y^{\prime}_{RH(T)}$ $(d,d_{1},t_{0})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)^{3/4},$
-$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log t_{0}}}{d_{1}}\right)^{3/4}\Bigg{\}},$
-with $A_{4}$ defined in Theorem 4, $B_{t_{0}}$ defined in Theorem 2, and $C_{3}$ defined in Lemma 6.
-Furthermore, by assuming $RH$ instead of $RH(T)$, the upper bound restriction on $t$ can be removed.
+$Y^{\prime}_{0}$ $(d_{1},t_{0})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)^{3/4},$
+$\exp\Bigg{(}\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log t_{0}}}{d_{1}}\right)^{3/4}\Bigg{\}},$
+with $d_{1}$, $J$, $Q_{1,j}$, $W_{j}$ defined as in Theorem 4, and $C_{3}$ defined as in Lemma 6.
 Proof.
-The proof is identical to that of Theorem 4, except that for $\sigma\leq 1+\delta_{1,t}$, by Lemmas 6 and 9, we can replace (48) with
-$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|$ $\leq|\zeta(1+\delta_{1,t})|^{3/4}|\zeta(1+\delta_{1,t}+2it)|^{1/4}$
-$\leq\left(\frac{e^{\gamma(d_{1}/\log t_{0})}}{d_{1}}\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}.$
-Meanwhile, for $\sigma\geq 1+\delta_{1,t}$, the same argument with (24) and (49) leads to
+The proof is identical to that of Theorem 4, except that in the case $\sigma\leq 1+\delta_{1}$, we replace (47) with
+$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|$ $\leq|\zeta(1+\delta_{1})|^{3/4}|\zeta(1+\delta_{1}+2it)|^{1/4}$
+$\leq\left(\frac{e^{\gamma(d_{1}/\log t_{0})}}{d_{1}}\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}$
+by using (51) with Lemmas 6 and 9.
+Meanwhile, for $\sigma>1+\delta_{1}$, the same argument with (20) and (9) leads to
 $\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\left(\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}$
 in place of (48). ∎
-Theorem 7.
-Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log H\leq 1$.
-Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37).
-Then in the region
-$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
-we have
-$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{H}(d,d_{1})(\log t)^{11/12},$
-where
-$Y^{\prime}_{H}$ $(d,d_{1})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log H}\right)^{3/4},$
-$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log H}}{d_{1}}\right)^{3/4}\Bigg{\}},$
-with $A_{5}$ defined in Theorem 5, and $C_{3}$ defined in Lemma 6.
-Proof.
-The proof is similar to that of Theorem 6, except we do not assume $RH(T)$. Then the process follows the proof of Theorem 5, making similar changes as with the earlier case $t\leq H$ in Theorem 6. ∎
-Combining Theorems 6 and 7, we prove Corollary 5, a result which holds for all $t\geq t_{0}\geq 5$.
 Proof of Corollary 5.
-The proof is analogous to that of Corollary 4, where the final result is obtained by optimising over the appropriate parameters.
-For the result on $\sigma=1$, the limiting region is again $2\exp(e^{2})\leq t\leq H$, where we have
-$\left|\frac{1}{\zeta(1+it)}\right|\leq 7.475(\log t)^{11/12},$
-with $d_{1}=0.72042$. ∎
-$t_{0}$ $Y^{\prime}(5.5587,t_{0})$ $d_{1}$ $W$ $Y^{\prime}(W,5)$ $d_{1}$
-$10$ $21.123$ $0.307588$ $5.5587$ $47.054$ $0.170052$
-$10^{2}$ $11.643$ $0.542830$ $6$ $43.568$ $0.173052$
-$10^{3}$ $9.842$ $0.648792$ $7$ $38.179$ $0.178401$
-$10^{4}$ $8.885$ $0.725028$ $8$ $34.753$ $0.182368$
-$10^{5}$ $8.247$ $0.787048$ $9$ $32.391$ $0.185428$
-$10^{7}$ $7.425$ $0.885011$ $10$ $30.668$ $0.187861$
-$10^{9}$ $6.907$ $0.960531$ $11$ $29.356$ $0.18984$
-$10^{11}$ $6.545$ $1.021397$ $12$ $28.325$ $0.191481$
-$H$ $6.340$ $1.059544$ $13$ $27.494$ $0.192568$
-Table 7: Values for $Y^{\prime}(W,t_{0})$ in Corollary 5, with corresponding $W$ or $t_{0}$, and parameter $d_{1}$ for the case $t_{0}\leq t\leq H$ (rounded to six decimal places).
-Each entry is valid for $t\geq t_{0}$.
+The proof, and process of optimisation here, even for the case $\sigma\geq 1$, are exactly the same as that of Corollary 4.
+These computed values of $Y^{\prime}_{0}$ for $t\geq 13$, are labelled $Y^{\prime}$, and are presented in Table 5.
+For all these cases, we found an optimal parameter of $d_{1}=0.030648$ (rounded to six decimal places).
+For the case $\sigma\geq 1$, we computed $Y^{\prime}=87.725$ for $t\geq 13$, also with optimal $d_{1}=0.030648$ (rounded).
+Finally, for the case $\sigma=1$, we computed $Y^{\prime}=81.900$ for $\sigma\geq 1$ and $t\geq 2\exp(e^{2})$, with $d_{1}=0.030793$ (rounded).
+Then by a similar calculation as in the proof of Corollary 4, we have by Lemma 7,
+$\left|\frac{1}{\zeta(1+it)}\right|\leq 1.5(\log t)^{11/12}\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}).$
+On the other hand, (49) implies
+$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.421(\log t)^{11/12}\quad\text{for}\quad 2\leq t\leq 500,$
+where this estimate is obtained by noting that $t$ attains its maximum at $500$. ∎
+$W$ $Y^{\prime}$
+$5.56$ $2.17\cdot 10^{32}$
+66 $6.99\cdot 10^{9}$
+77 $5.21\cdot 10^{6}$
+88 $2.55\cdot 10^{5}$
+99 $55579$
+$10$ $21649$
+$11$ $11331$
+$12$ $7051$
+$13$ $4904$
+Table 5: Values for $Y^{\prime}$ in Corollary 5, with corresponding $W$. Each entry is valid for $t\geq 13$.
 6 Concluding Remarks
@@
 Minor improvements can be made by a more careful treatment of the sum over zeros and the integral arising from the application of Lemma 2.
-Other possibilities are: an improvement on the Borel–Carathéodory by specialising it more to this problem, proving a uniform (in $\sigma$) bound for $\arg\zeta(s)$, or using stronger estimates for $\zeta(s)$ under $RH$ or $RH(T)$.
+For Theorem 4, a better smoothing argument could be used in place of (45).
+Other possibilities for improvement are: a refinement on the Borel–Carathéodory by specialising it more to this problem, using stronger estimates for $\zeta(s)$ under $RH$ or $RH(T)$, or obtaining a good bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ for use in (42).
 It is evident from the results that the closer $s$ is to the boundary of the zero-free region, and hence a potential zero, the larger the constants.
-However, the large constants can also be attributed in part to the factor $(R+r)/(R-r)$ in Theorem 1, which increases rapidly as $r$ approaches $R$.
-This contributes significantly to the overall estimate, even if $|f(s_{0})|$ is small, and seems to suggest that Theorem 1 in its current form might not be optimal for this problem.
-The main advantage of our method using Lemma 1 is that it eliminates the need for Theorem 1 in estimating $|1/\zeta(s)|$, and also reduces the number of applications of the theorem when estimating $|\zeta^{\prime}(s)/\zeta(s)|$.
-It might be possible to remove this entirely, by directly bounding the real and imaginary part of $|\zeta^{\prime}(s)/\zeta(s)|$, as one can easily prove an analogue of Lemma 1 which relates $\textup{Im}(f^{\prime}/f)$ to $\arg f$.
-To use such a result in the case of $f(s)=\zeta(s)$, besides dealing with the corresponding sum over zeros, a uniform bound (over a range of $\sigma$) on $S_{\sigma}(t):=(1/\pi)\arg\zeta(\sigma+it)$ is needed.
+However, the large constants can also be attributed in part to the factor (R+r)/(Rr)(R+r)/(R-r) in Theorem 1, which increases rapidly as $r$ approaches $R$.
+This contributes significantly to the overall estimate, even if $|f(s_{0})|$ is small, and seems to suggest that Theorem 1 in its current form might not be optimal for bounding $|\zeta^{\prime}(s)/\zeta(s)|$.
+An advantage of obtaining a bound on ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))$ is to reduce the usage of Theorem 1, when bounding $|1/\zeta(s)|$.
+This is because we would be able to apply such a bound directly to (42), whereas we currently use ${\textup{Re}}(\zeta^{\prime}(s)/\zeta(s))\leq|\zeta^{\prime}(s)/\zeta(s)|$.
+This alternative could be approached via Lemma 2, provided one were able to deal with the resulting sum over zeros well.
 To get unconditional results we have (due to knowledge of the Riemann height) assumed a partial Riemann hypothesis to prove some of our results.
-As an example, for $W=5.5587$, the condition (37) in Theorem 5 is only fulfilled for $t\geq 3.22\cdot 10^{9}$.
-The conditions on $d$ (which depends on $t$) are what prevents us from taking $W$ too small, and this is where the Riemann height and $RH(T)$ helps us.
-In fact we only need a ‘weak’ version of $RH$, for instance in (26), that up to height $T$, no zeros lie to the right of $1/2-R^{\prime}$.
 The downside to this is that we have not used better available bounds that require the full strength of $RH$.
@@
 As an illustration, Corollary 5, though asymptotically superior, beats Corollary 4 only when $t$ is large.
-In the case of $\sigma=1$, this happens for $t\geq\exp(161975)$.
+In the case of $\sigma=1$, this happens for $t\geq\exp(2.2\cdot 10^{5})$.
 This is in part due to the relatively large constants in Corollary 5, and a direct way of lowering these constants is an improvement in the bound $|\zeta(1+it)|$, which correspondingly lowers $C_{3}$ in Lemma 6.
@@
 In general, when studying zero-free regions for $\zeta(s)$, increasing the degree of the non-negative trigonometric polynomial improves the results one gets.
-Surprisingly, the same cannot be said in our setting (with the exception of results in Section 5.4), and the classical polynomial we used is actually optimal.
-Take for instance Lemma 12, which is used in Theorems 4 and 5.
+Surprisingly, the same cannot be said in our setting, and the classical polynomial we used is actually optimal.
+Take for instance Lemma 12, which is used in Theorem 4.
 Here the overall order of the bound is $O(\log t)$, and in this scenario the article [20] has an in-depth discussion on why the classical polynomial is best.
-On the other hand, in Theorems 6 and 7, the question is whether higher degree polynomials lead to a reduction in the exponent $11/12$, even if overall constants get worse.
+On the other hand, in Theorem 5, the question is whether higher degree polynomials lead to a reduction in the exponent $11/12$, even if overall constants get worse.
 This is an interesting question to consider, since in this case the answer is less obvious, although preliminary calculations suggest that the classical polynomial is still the optimal one.
 Acknowledgements
-We thank Tim Trudgian for his guidance throughout the writing of this paper. We are also very grateful to Bryce Kerr, Olivier Ramaré, and Aleksander Simonič for the helpful discussions.
+We thank Tim Trudgian for his guidance throughout the writing of this paper.
+We also thank the referee for the care taken in going through our article.
+Finally, we are very grateful to Andrew Fiori, Bryce Kerr, Olivier Ramaré, Aleksander Simonič, for the helpful discussions, and especially to Harald Helfgott for pointing out an error in an earlier version of this paper.